### v1

# Stationary solutions to stochastic 3D Euler equations in Hölder space

Abstract.
We establish the existence of infinitely many global and stationary solutions in $C(\mathbb{R};C^{\vartheta})$ space for some $\vartheta>0$ to the three dimensional Euler equations driven by an additive noise. The result is based on a new stochastic version of the convex integration method, incorporating the stochastic convex integration method developed in [HZZ22b] and pathwise estimates to derive uniform moment estimates independent of time.
Key words and phrases:
stochastic Euler equations, stationary solutions, Hölder space, convex integration.
2010 Mathematics Subject Classification:
60H15; 35R60; 35Q30
Research supported by National Key R&D Program of China (No. 2022YFA1006300) and the NSFC (No. 12271030). The financial support by the DFG through the CRC 1283 ”Taming uncertainty and profiting from randomness and low regularity in analysis, stochastics and their applications” is greatly acknowledged.
1. Introduction
In this paper, we are considered with stochastic Euler equations governing the time evolution of the velocity $u$ of an inviscid fluid on the three dimensional torus $\mathbb{T}^{3}=\mathbb{R}^{3}/(2\pi\mathbb{Z})^{3}$. The system reads as
${\mathord{{\rm d}}}u+\mathord{{\rm div}}(u\otimes u)\,{\mathord{{\rm d}}}t+\nabla P\,{\mathord{{\rm d}}}t$ $={\mathord{{\rm d}}}B,$ (1.1)
$\mathord{{\rm div}}u$ $=0,$
where $P$ stands for the corresponding pressure. The right hand side represents a random external force acting on the fluid, which is given by a $GG^{*}$-Wiener process $B$ defined on some probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $GG^{*}$ being trace class operator. In early works, stochastic Euler equations have been widely studied in both two and three dimensions, see e.g.[BFM16], [BP01], [CFH19], [GHV14], where local-in-time solutions for (1.1) have been constructed. Our goal is to investigate the existence of global and stationary solutions with Hölder continuity to the above system (1.1). To achieve this, we employ the method of convex integration, a technique that has gained prominence in fluid dynamics over the past decades. This method was initially introduced following [DLS10, DLS12, DLS13] by De Lellis and Székelyhidi Jr. In these works, the method of convex integration was used in order to prove non-uniqueness of weak solutions with Hölder continuity to Euler equations. Subsequently, this method has led to several breakthroughs in the fluid dynamics, including proof of Onsager’s conjecture for the incompressible Euler equations [Ise18, BDLSV19]. We refer to the excellent review article by Buckmaster and Vicol [BV19] for a gentle introduction and further references. In recent years, convex integration has already been applied in a stochastic setting, namely, to the isentropic Euler system by Breit, Feireisl and Hofmanová [BFH20] and to the full Euler system by Chiodaroli, Feireisl and Flandoli [CFF19] with linear multiplicative noise. Futhermore, ill/well-posedness of dissipative martingale solutions to stochastic 3D Euler equations was studied by Hofmanová, Zhu and Zhu [HZZ22a], along with the existence and non-uniqueness of strong Markov solutions. Later, Hölder continuous, global-in-time probabilistically strong solutions to 3D Euler equations perturbed by Stratonovich transport noise was constructed in [HLP22]. In the mostly recent work [HZZ22b], existence and non-uniqueness of stationary solutions in $L^{2}$ to the stochastic 3D Euler equations was established, using a new stochastic convex integration method. However, the regularity of the global and stationary solutions to the stochastic 3D Euler equations with additive noise, as constructed in the previous work, cannot achieve Hölder continuity. Inspired by these previous works and the progress in the deterministic case, it is natural to investigate the existence and non-uniqueness of global and stationary Hölder continuous solutions to the Euler equations in the stochastic setting (1.1). Another motivation for our research is linked to the Onsager’s conjecture. It is constructed in [Ise18, BDLSV19] that solutions in $C^{\frac{1}{3}-}$ do not conserve energy to the deterministic Euler equations. It is natural to ask whether the regularity of the global, nonunique solutions in the stochastic setting can stay $C^{\frac{1}{3}-}$ a.s.. To summarize the above discussion, it is natural to investigate the validity of the following claims.
  1. (i)
    Existence and non-uniqueness of global $\vartheta-$Hölder continuous solutions to the stochastic Euler equations (1.1), for some $\vartheta>0$.
  2. (ii)
    Existence and non-uniqueness of stationary solutions to the stochastic Euler equations (1.1).
1.1. Main results
Our aim is to provide some answers to the above problems (i) and (ii) in the physically relevant context of the stochastic Euler equations on $\mathbb{T}^{3}$ driven by additive stochastic noise. As the main tool within this study, we apply a stochastic version of the convex integration method. Unlike the previous works using convex integration for the Navier–Stokes and Euler equations in the stochastic setting [HZZ19, HZZ22a, HZZ23a, HZZ23b], we are now inspired by [CDZ22] and [HZZ22b] where they overcome the limitation originating from stopping times, previously used to control the noise terms in the iteration. More precisely, we no longer work with stopping times and instead we include expectations in the inductive iteration and provide uniform moment estimates. However, the construction of Hölder continuous Euler flows in the deteministic setting, as demonstrated in [BV19], requires to introduce transport equations, and pathwise positive lower bounds for the solutions are used to improve regularity. As a result, we also introduce some pathwise estimates during iteration for convex integration. To this end, we use cutoff method to control the growth of noise pathwisely. Compared to only using stochastic convex integration, the combination of pathwise estimates with it can provide better Hölder regularity (see Remark 1.3 below). This approach also allows us not to control all the moments bounds during the iteration. We hope this cutoff method can be applied to enhance regularity of the solutions in the future work. As we are interested in the long time behavior of solutions and particularly in the construction of statistically stationary solutions, we work with entire solutions solving the equations for all times $t\in{\mathbb{R}}$. Accordingly, the norms in the convex integration scheme must be chosen appropriately in order to provide the desired global-in-time estimates. This is achieved through bounds of the form
$\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{t\leqslant s\leqslant t+1}\|u(s)\|_{C^{\vartheta}}^{p}\right]<\infty,$
with $p\geqslant 1$ and small $\vartheta>0$. Such bounds provide uniform moment estimates locally in $C(\mathbb{R};C^{\vartheta})$ and guarantee the convergence of the corresponding ergodic averages. This leads to the existence of stationary solutions. Within this study, we focus on analytically weak solutions which satisfy the equations in the following sense.
Definition 1.1.
We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is an analytically weak solution to the Euler system (1.1) provided
  1. (1)
    $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ is a stochastic basis with a complete right-continuous filtration;
  2. (2)
    $B$ is an $\mathbb{R}^{3}$-valued, spatial mean and divergence free, two-sided $GG^{*}$-Wiener process with respect to the filtration $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$;
  3. (3)
    the velocity $u\in C({\mathbb{R}}\times\mathbb{T}^{3})$ $\mathbf{P}$-a.s. and is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted;
  4. (4)
    for every $-\infty<s\leqslant t<\infty$ it holds $\mathbf{P}$-a.s.
    $\langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r\ =\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle$
    for all $\psi\in C^{\infty}(\mathbb{T}^{3})$, $\mathord{{\rm div}}\psi=0$.
With this definition at hand, our first main result reads as follows, tackling the problem (i) above.
Theorem 1.2.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exists an analytically weak solution to (1.1) in the sense of Definition 1.1 which belongs to $C(\mathbb{R},C^{\vartheta})$, for some $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}\right)$.
Remark 1.3.
(i) In Section 3, we prove this result by combining stochastic convex integration method with pathwise estimates to derive uniform moment estimates independent of time, and obtain Hölder continuous solutions, where $\vartheta$ can be chosen less than ${\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}$. While we also prove this result in a separable article (see [Lü24]), which applies stochastic convex integration method to the the construction of Hölder continuous Euler flows in [DLS13] to derive uniform moment estimates independent of time. The value of $\vartheta$ in this case decreases due to the higher moments required, and we may take $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{73{\cdot}62^{3}},\frac{1}{7{\cdot}62^{3}}\}}$. (ii) It is natural to ask whether we could construct the solutions with better regulartiy to (1.1) by combining the gluing method in [BDLSV19] with our cutoff method. This will be studied in our future work.
The second result is the existence and non-uniqueness of stationary solutions in Hölder space. In the sequel, we understand stationarity in the sense of shift invariance of laws of solutions on the space of trajectories, see also [BFHM19, BFH20e, FFH21, HZZ22b]. More precisely, we define the joint trajectory space for the solution and the driving Brownian motion as
${\mathcal{T}}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$
for some $\kappa>0$ and let $S_{t}$, $t\in{\mathbb{R}}$, be shifts on trajectories given by
$S_{t}(u,B)(\cdot)=(u(\cdot+t),B(\cdot+t)-B(t)),\quad t\in{\mathbb{R}},\quad(u,B)\in{\mathcal{T}}.$
We note that the shift in the second component acts differently in order to guarantee that for a Brownian motion $B$ the shift $S_{t}B$ is again a Brownian motion.
Definition 1.4.
We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is a stationary solution to the stochastic Euler equations (1.1) provided it satisfies (1.1) in the sense of Definition 1.1 and its law is shift invariant, that is,
$\mathcal{L}[S_{t}(u,B)]=\mathcal{L}[u,B]\qquad\text{ for all }\quad t\in{\mathbb{R}}.$
With these definitions at hand, we are able to state our second main result, tackling the problem (ii) above. The result then reads as follows and is proved in Section 5.
Theorem 1.5.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exist infinitely many stationary solutions to the stochastic Euler equations (1.1). Moreover, the solutions belong to $C({\mathbb{R}},C^{\vartheta})$ a.s. for some $\vartheta>0$, satisfying
$\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{t\leqslant s\leqslant t+1}\|u(s)\|_{C^{\vartheta}}\right]<\infty.$
Organization of the paper
In Section 2, we collect the basic notations used throughout the paper. Section 3 and Section 4 are the core of our proofs: the stochastic convex integration is developed and employed to construct entire non-unique analytically weak solutions. In Section 3, we present the main iteration Proposition 3.2 to prove our main Theorem 3.3. Section 4 is devoted to the construction of stochastic convex integration with pathwise estimates, namely, proof of iteration Proposition 3.2. Then the results in Section 3 are used in Section 5 together with a Krylov–Bogoliubov’s argument to obtain existence of non-unique stationary solutions to the stochastic Euler equations. In Appendix A, we recall the construction of Beltrami waves from [BV19] . In Appendix B, we give estimates on amplitude functions used in the convex integration construction. Finally, the estimates for transport equations are completed in Appendix C.
2. Notations
2.1. Function spaces
Throughout the paper, we employ the notation $a\lesssim b$ if there exists a constant $c>0$ such that $a\leqslant cb$. We let ${\mathbb{N}}_{0}:={\mathbb{N}}\cup\{0\}$. Given a Banach space $E$ with a norm $\|\cdot\|_{E}$ and $t\in{\mathbb{R}}$, we write $C_{t}E=C([t,t+1];E)$ for the space of continuous functions from $[t,t+1]$ to $E$, equipped with the supremum norm $\|f\|_{C_{t}E}=\sup_{s\in[t,t+1]}\|f(s)\|_{E}$. For $\alpha\in(0,1)$ we define $C^{\alpha}_{t}E$ as the space of $\alpha$-Hölder continuous functions from $[t,t+1]$ to $E$, endowed with the norm $\|f\|_{C^{\alpha}_{t}E}=\sup_{s,r\in[t,t+1],s\neq r}\frac{\|f(s)-f(r)\|_{E}}{|r-s|^{\alpha}}+\sup_{s\in[t,t+1]}\|f(s)\|_{E}$. For $\alpha\in(0,1)$, we also write $C^{\alpha}$ for functions in the space of $\alpha$-Hölder continuous functions from $\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$ endowed with the norm $\|f\|_{C^{\alpha}}=\sup_{x,y\in\mathbb{T}^{3},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in\mathbb{T}^{3}}|f(x)|.$ We use $L^{p}$ to denote the set of standard $L^{p}$-integrable functions from $\mathbb{T}^{3}$ to $\mathbb{R}^{3}$. For $s>0$, $p>1$ we set $W^{s,p}:=\{f\in L^{p};\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}<\infty\}$ with the norm $\|f\|_{W^{s,p}}=\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}$. Set $L^{2}_{\sigma}=\{f\in L^{2};\int_{\mathbb{T}^{3}}f\,{\mathord{{\rm d}}}x=0,\mathord{{\rm div}}f=0\}$. For $s>0$, we define $H^{s}:=W^{s,2}\cap L^{2}_{\sigma}$. For $s<0$ we define $H^{s}$ to be the dual space of $H^{-s}$. For $t\in\mathbb{R}$ and a domain $D\subset\mathbb{R}$ we denote by $C^{N}_{t,x}$ and $C^{N}_{D,x}$, respectively, the space of $C^{N}$-functions on $[t,t+1]\times\mathbb{T}^{3}$ and on $D\times\mathbb{T}^{3}$, respectively, $N\in\mathbb{N}_{0}$. The spaces are equipped with the norms
$\|f\|_{C^{N}_{t,x}}=\sum_{\begin{subarray}{c}0\leqslant n+|\alpha|\leqslant N\\ n\in\mathbb{N}_{0},\alpha\in\mathbb{N}^{3}_{0}\end{subarray}}\|\partial_{t}^{n}D^{\alpha}f\|_{L^{\infty}_{[t,t+1]}L^{\infty}},\qquad\|f\|_{C^{N}_{D,x}}=\sum_{\begin{subarray}{c}0\leqslant n+|\alpha|\leqslant N\\ n\in\mathbb{N}_{0},\alpha\in\mathbb{N}^{3}_{0}\end{subarray}}\sup_{t\in D}\|\partial_{t}^{n}D^{\alpha}f\|_{L^{\infty}}.$
For a Polish space $H$ we denote by $\mathcal{B}(H)$ the $\sigma$-algebra of Borel sets in $H$. We also use $\mathring{\otimes}$ to denote the trace-free part of the tensor product. We recall the inverse divergence operator $\mathcal{R}$ from [DLS13, Section 4.5], which acts on vector fields $v$ with $\int_{\mathbb{T}^{3}}v{\mathord{{\rm d}}}x=0$ as
$(\mathcal{R}v)^{kl}=(\partial_{k}\Delta^{-1}v^{l}+\partial_{l}\Delta^{-1}v^{k})-\frac{1}{2}(\delta_{kl}+\partial_{k}\partial_{l}\Delta^{-1})\mathord{{\rm div}}\Delta^{-1}v,$
for $k,l\in\{1,2,3\}$. Then $\mathcal{R}v(x)$ is a symmetric trace-free matrix for each $x\in\mathbb{T}^{3}$, and $\mathcal{R}$ is a right inverse of the div operator, i.e. $\mathord{{\rm div}}(\mathcal{R}v)=v$. By [CL22, Theorem B.3] we have for $1\leqslant p\leqslant\infty$
$\|\mathcal{R}f\|_{L^{p}(\mathbb{T}^{3})}\lesssim\|f\|_{L^{p}(\mathbb{T}^{3})}.$ (2.1)
2.2. Probabilistic elements
For a given probability measure $\mathbf{P}$ we denote by $\mathbf{E}$ the expectation under $\mathbf{P}$. Regarding the driving noise, we assume that $B$ is $\mathbb{R}^{3}$-valued two-sided $GG^{*}$-Wiener process with zero spatial mean and zero divergence, defined on some probability space $(\Omega,\mathcal{F},\mathbf{P})$ and $G$ is a Hilbert–Schmidt operator from $U$ to $L_{\sigma}^{2}$ for some Hilbert space $U$. For $p\in[1,\infty)$, $\delta\in(0,\frac{1}{2})$ and some $\kappa>0$, we denote
$|\!|\!|u|\!|\!|^{p}_{C^{0},p}:=\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{s\in[t,t+1]}\|u(s)\|^{p}_{C^{0}}\right],\quad|\!|\!|u|\!|\!|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa},p}:=\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|u\|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa}}\right].$
These norms define function spaces of random variables on $\Omega$ taking values in $C({\mathbb{R}}\times\mathbb{T}^{3})$ and $C^{\frac{1}{2}-\delta}({\mathbb{R}},C^{\kappa})$, respectively, with bounds in $L^{p}(\Omega;{C^{0}(I\times\mathbb{T}^{3}))}$ and $L^{p}(\Omega;{C^{\frac{1}{2}-\delta}(I,C^{\kappa})})$ for bounded interval $I\subset{\mathbb{R}}$. Furthermore, the bounds only depend on the length of the interval $I$, not on its location within ${\mathbb{R}}$. In the sequel, we simply say that $u$ has a uniform moment of order $p$ locally in $C^{0}({\mathbb{R}}\times\mathbb{T}^{3})$ provided $|\!|\!|u|\!|\!|^{p}_{C^{0},p}<\infty$. Similarly, we define the corresponding norms with $C^{0}$ replaced by $C^{\kappa}$, $C^{1}$, $L^{1}$ and $H^{\frac{3}{2}+\kappa}$.
3. Stochastic convex integration with pathwise estimates and results
In [CDZ22] and [HZZ22b], they presented a stochastic convex integration method constructing directly solutions on the whole time line $\mathbb{R}$ by introducing expectations to the iterative estimates in convex integration. In this section, we integrate their methods with pathwise estimates in the context of the stochastic Euler equations. Our goal is to construct a Hölder continuous Euler flow, mirroring the deterministic case as presented in [BV19], where the transport equations (see (4.7) below) are used to improve regularity. This method necessitates pathwise positive lower bounds for the solutions of the transport equations. As a result, moment bounds during the inductive iteration become insufficient. Hence, we introduce additional pathwise estimates in convex integration schemes. More precisely, our idea employs the cutoff method to control the growth of the noise, enabling the introduction of pathwise estimates during the inductive iteration. This is combined with moment bounds to derive desired uniform estimates. Compared to only using moment bounds in the inductive iteration in [CDZ22] and [HZZ22b], the advantage of this approach is that it only requires lower moments of the solutions, and avoids the requirements for higher moments as in [HZZ22b]. This can simplify the computations and enhance the regularity of the solutions. At the first step, we decompose stochastic Euler system (1.1) into two parts, one is linear and involves the noise, whereas the second one is a random PDE. More precisely, we consider the stochastic linear equation
${\mathord{{\rm d}}}z+z{\mathord{{\rm d}}}t$ $={\mathord{{\rm d}}}B,$ (3.1)
$\mathord{{\rm div}}z$ $=0,$
where $B$ is $\mathbb{R}^{3}$-valued two-sided trace-class $GG^{*}$-Wiener process with spatial zero mean and divergence free (see e.g. [LR15, Section 2.1]), and $v$ solves the nonlinear equation
$\partial_{t}v-z+\mathord{{\rm div}}((v+z)\otimes(v+z))+\nabla P$ $=0,$ (3.2)
$\mathord{{\rm div}}v$ $=0.$
Here, $z$ is divergence free by the assumptions on the noise $B$ and by $P$ we denote the pressure term associated to $v$. We include the linear zero order term $z$ on the left hand side of (3.1) in order to obtain a stabilization of the equation needed for the necessary global in time estimates. It will be seen in the course of the construction that the corresponding counter term in (3.2) will not cause any difficulties. Using the factorization method, it is standard to derive regularity of the stochastic convolution $z$ on a given stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ canonical filtration given in [LR15, Section 2.1]. In particular, the following result follows from [DPZ92, Theorem 5.16] together with the Kolmogorov continuity criterion.
Proposition 3.1.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\kappa}GG^{*})<\infty$ for some $\kappa>0$. Then for any $\delta\in(0,\frac{1}{2})$, $p\geqslant 2$
$\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}{C^{\kappa}}}\right]\lesssim\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}H^{\frac{3}{2}+\kappa}}\right]\leqslant(p-1)^{p/2}L^{p},$ (3.3)
where $L\geqslant 1$ depends on $\mathrm{Tr}((-\Delta)^{3/2+\kappa}GG^{*})$, $\delta$ and is independent of $p$.
Proof.
We recall that the unique stationary solution to (3.1) has the explicit form $z(t)=\int_{-\infty}^{t}e^{s-t}{\mathord{{\rm d}}}B_{s}$. By [LR15, Section 2.1], the Wiener process $B$ can be written as $B=\sum_{k\in{\mathbb{N}}}\sqrt{c_{k}}\beta_{k}e_{k}$ for an orthonormal basis $\{e_{k}\}_{k\in{\mathbb{N}}}$ of $L^{2}_{\sigma}$ consisting of eigenvectors of $GG^{*}$ with corresponding eigenvalues $c_{k}$, and the coefficients satisfy $\sum_{k\in{\mathbb{N}}}c_{k}<\infty$, where $\{\beta_{k}\}_{k\in{\mathbb{N}}}$ are a sequence of mutually independent standard two-sided real-valued Brownian motions. Then it holds for $t\geqslant s$,
$\mathbf{E}\|z(t)-z(s)\|^{2}_{H^{\frac{3}{2}+\kappa}}$
$=\mathbf{E}\left\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}\int_{s}^{t}e^{r-t}{\mathord{{\rm d}}}B_{r}\right\|_{L^{2}}^{2}+\mathbf{E}\left\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}\int_{-\infty}^{s}(e^{r-t}-e^{r-s}){\mathord{{\rm d}}}B_{r}\right\|^{2}_{L^{2}}$
$=\sum_{k=1}^{\infty}c_{k}\int_{s}^{t}\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}e_{k}\|_{L^{2}}^{2}e^{2(r-t)}{\mathord{{\rm d}}}r$
$\qquad+\sum_{k=1}^{\infty}c_{k}\int_{-\infty}^{s}\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}e_{k}\|_{L^{2}}^{2}(e^{r-t}-e^{r-s})^{2}{\mathord{{\rm d}}}r$
$\leqslant\sum_{k=1}^{\infty}c_{k}\left(\|e_{k}\|_{L^{2}}^{2}+\|(-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}e_{k}\|_{L^{2}}^{2}\right)\left(\int_{s}^{t}e^{2(r-t)}{\mathord{{\rm d}}}r+\int_{-\infty}^{s}(e^{r-t}-e^{r-s})^{2}{\mathord{{\rm d}}}r\right)$
$=(\mathrm{Tr}(GG^{*})+\mathrm{Tr}((-\Delta)^{3/2+\kappa}GG^{*}))\left(\int_{s}^{t}e^{2(r-t)}{\mathord{{\rm d}}}r+\int_{-\infty}^{s}(e^{r-t}-e^{r-s})^{2}{\mathord{{\rm d}}}r\right)$
$\leqslant(\mathrm{Tr}(GG^{*})+\mathrm{Tr}((-\Delta)^{3/2+\kappa}GG^{*}))(t-s).$
Using Gaussianity we have
$\mathbf{E}\|z(t)-z(s)\|^{p}_{H^{\frac{3}{2}+\kappa}}\leqslant(p-1)^{\frac{p}{2}}\Big{(}\mathbf{E}\|z(t)-z(s)\|_{H^{\frac{3}{2}+\kappa}}^{2}\Big{)}^{\frac{p}{2}}.$
The result then follows from Kolmogorov’s continuity criterion and fundamental Sobolev embedding $H^{\frac{3}{2}+\kappa}\hookrightarrow C^{\kappa}$, i.e. for any $\delta\in(0,\frac{1}{2})$
$\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}{C^{\kappa}}}\right]\lesssim\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}H^{\frac{3}{2}+\kappa}}\right]\leqslant(p-1)^{p/2}L^{p},$
where $L\geqslant 1$ depends on $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\kappa}GG^{*})$, $\delta$ and is independent of $p$. ∎
3.1. Stochastic convex integration set-up and iterative proposition
Let us now explain how the convex integration iteration is set up. We consider an increasing sequence $\{\lambda_{q}\}_{q\in\mathbb{N}_{0}}$ which diverges to $\infty$, and a bounded sequence $\{\delta_{q}\}_{q\in\mathbb{N}}$ which decreases to 00. We choose $a,b\in\mathbb{N}$ and $\beta\in(0,1)$ and let
$\lambda_{q}=a^{b^{q}},\quad\delta_{1}=3rL^{2},\quad\delta_{q}=\frac{1}{2}\lambda_{2}^{2\beta}\lambda_{q}^{-2\beta},\ q\geqslant 2.$
Here $\beta$ will be chosen sufficiently small, $a$ will be chosen sufficiently large, and more details on the choice of these parameters will be given below in the course of the construction. The iteration is indexed by a parameter $q\in\mathbb{N}_{0}$. At each step $q$, a pair $(v_{q},\mathring{R}_{q})$ is constructed to solve the following system
$\partial_{t}v_{q}-z_{q}+\mathord{{\rm div}}((v_{q}+z_{q})\otimes(v_{q}+z_{q}))+\nabla p_{q}$ $=\mathord{{\rm div}}\mathring{R}_{q},$ (3.4)
$\mathord{{\rm div}}v_{q}$ $=0.$
In the above we define
$z_{q}(t,x)=\chi_{q}\left(\|\tilde{z}_{q}(t)\|_{C^{1}_{x}}\right)\tilde{z}_{q}(t,x),$ (3.5)
where $\chi_{q}$ is a non-increasing smooth function such that
$\chi_{q}(x)=\begin{cases}1,&x\in[0,\lambda_{q+2}^{\gamma}],\\ 0,&x\in(\lambda_{q+3}^{\frac{1}{4}\gamma},\infty),\end{cases}$
with derivative bounded by $1$, which requires
$\frac{1}{\lambda_{q+3}^{\gamma/4}-\lambda_{q+2}^{\gamma}}\leqslant 1.$ (3.6)
And $\tilde{z}_{q}=\mathbb{P}_{\leqslant f(q)}z$ with $f(q)=\lambda_{q+1}^{\gamma/8}$. Here $\mathbb{P}_{\leqslant f(q)}$ is the Fourier multiplier operator, which projects a function onto its Fourier frequencies $\leqslant f(q)$ in absolute value, and $\gamma$ is another small parameter which will be fixed in Section 4.1. $\mathring{R}_{q}$ on the right hand of (3.4) is a $3\times 3$ matrix which is trace-free and we put the trace part into the pressure. Keeping (3.3) in mind, by the Sobolev embedding $\|f\|_{L^{\infty}}\lesssim\|f\|_{H^{\frac{3}{2}+\kappa}}$ for $\kappa>0$, we observe that for any $p\geqslant 2$ and $\delta\in(0,\frac{1}{2})$
$|\!|\!|z_{q}|\!|\!|_{C^{0},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{H^{\frac{3}{2}+\kappa},p}\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\kappa},p}\lesssim(p-1)^{\frac{1}{2}}L,$ (3.7)
$|\!|\!|z_{q}|\!|\!|_{C^{1},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{1},p}\lesssim f(q)|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\kappa},p}\lesssim(p-1)^{\frac{1}{2}}L\lambda_{q+1}^{\frac{\gamma}{8}},$
$|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{1}_{x},2p}|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2p}\lesssim\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}H^{\frac{3}{2}+\kappa},2p}^{2}\lesssim(2p-1)L^{2}\lambda_{q+1}^{\frac{\gamma}{8}}.$
The above estimates (3.7) will be frequently used in the sequel. Under the above assumptions, our main iteration reads as follows, and the proof of this result is presented in Section 4 below.
Proposition 3.2.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$ and let $r>1$ be fixed. Given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$, where $L$ is from Proposition 3.2 and $c_{R}>0$ is a small universal constant (see Section 4.3). There exists a choice of parameters $a,b,\beta$ and $\gamma$ such that the following holds true: Let $(v_{q},\mathring{R}_{q})$ for some $q\in\mathbb{N}_{0}$ be an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted solution to (3.4) satisfying
$\|v_{q}\|_{C_{t,x}^{0}}$ $\leqslant\lambda_{q+2}^{\gamma},$ (3.8a)
$|\!|\!|v_{q}|\!|\!|_{C^{0},2r}$ $\leqslant\lambda_{q}^{\beta},$ (3.8b)
$\|v_{q}\|_{C_{t,x}^{1}}$ $\leqslant\lambda^{\frac{3}{2}}_{q}\delta_{q}^{\frac{1}{2}},$ (3.8c)
$\|\mathring{R}_{q}\|_{C^{0}_{t,x}}$ $\leqslant\lambda^{\gamma}_{q+3},$ (3.8d)
$|\!|\!|\mathring{R}_{q}|\!|\!|_{C^{0},r}$ $\leqslant\delta_{q+1},$ (3.8e)
$|\!|\!|\mathring{R}_{q}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{48}c_{R}\delta_{q+2}\underline{e},$ (3.8f)
Moreover, for any $t\in{\mathbb{R}}$
$\frac{3}{4}\delta_{q+1}e(t)\leqslant e(t)-\mathbf{E}\|(v_{q}+z_{q})(t)\|_{L^{2}}^{2}\leqslant\frac{5}{4}\delta_{q+1}e(t),\quad$ $q\geqslant 1,$ (3.9)
$\mathbf{E}\|v_{0}(t)+z_{0}(t)\|^{2}_{L^{2}}\leqslant\frac{1}{2}e(t),\quad$ $q=0.$
Then there exists an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted process $(v_{q+1},\mathring{R}_{q+1})$ which solves (3.4), obeys (3.8a)-(3.8f) and (3.9) at the level $q+1$ and satisfies
$|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\bar{M}\delta_{q+1}^{\frac{1}{2}},$ (3.10)
where $\bar{M}$ is a universal constant which will be fixed throughout the iteration.
We intend to start the iteration from $v_{0}\equiv 0$ on ${\mathbb{R}}$. In this case, we have $\mathring{R}_{0}=z_{0}\mathring{\otimes}z_{0}-\mathcal{R}z_{0}$. Recalling the definition of $z_{0}$ in (3.5), we obtain
$\|\mathring{R}_{0}\|_{C^{0}_{t,x}}\leqslant\|z_{0}\|^{2}_{C^{0}_{t,x}}+\|z_{0}\|_{C^{0}_{t,x}}\leqslant\lambda_{3}^{\frac{1}{2}\gamma}+\lambda_{3}^{\frac{1}{4}\gamma}\leqslant\lambda_{3}^{\gamma},$
where we choose $a$ sufficiently large to absorb the constant in the last inequality. Taking expectation and using (3.7), we obtain
$|\!|\!|\mathring{R}_{0}|\!|\!|_{C^{0},r}\leqslant|\!|\!|z_{0}|\!|\!|^{2}_{C^{0},2r}+|\!|\!|z_{0}|\!|\!|_{C^{0},r}\leqslant|\!|\!|z|\!|\!|^{2}_{H^{\frac{3}{2}+\sigma},2r}+|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},r}\leqslant 2rL^{2}+rL\leqslant\delta_{1},$
$|\!|\!|\mathring{R}_{0}|\!|\!|_{L^{1},1}\leqslant(2\pi)^{3}|\!|\!|\mathring{R}_{0}|\!|\!|_{C^{0},1}\leqslant 3{\cdot}(2\pi)^{3}L^{2}\leqslant\frac{1}{96}c_{R}\underline{e}\leqslant\frac{1}{48}c_{R}\delta_{2}e(t),$
where we used $6{\cdot}48{\cdot}(2\pi)^{3}c_{R}^{-1}L^{2}\leqslant\underline{e}$, $\delta_{1}=3rL^{2}$ and $\delta_{2}=\frac{1}{2}$. Hence, (3.8d), (3.8e) and (3.8f) are satisfied on the level $q=0$. Similarly, (3.9) also holds at the level $q=0$, since we have
$\mathbf{E}\|z_{0}(t)\|^{2}_{L^{2}}\leqslant(2\pi)^{3}\mathbf{E}\|z_{0}(t)\|^{2}_{C^{0}}\leqslant(2\pi)^{3}\mathbf{E}\|z(t)\|^{2}_{H^{\frac{3}{2}+\sigma}}\leqslant(2\pi)^{3}L^{2}\leqslant\frac{1}{2}e(t).$
We will see in the following context, as long as the $q=0$ step in (3.9) holds ture, $\zeta_{0}(t)$ in (4.11) below is well-defined, thus the iteration can proceed.
3.2. Proof of main result
Since the first iteration is established, we deduce the following result by using Proposition 3.2.
Theorem 3.3.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$. Let $r>1$ be fixed and given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$. There exists an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted process $u$ which belongs to $C({\mathbb{R}};C^{\vartheta})$ $\mathbf{P}$-a.s. for some $\vartheta>0$ and is an analytically weak solution to (1.1) in the sense of Definition 1.1. Moreover, the solution satisfies
$|\!|\!|u|\!|\!|_{C^{\vartheta},2r}<\infty,$ (3.11)
and for all $t\in{\mathbb{R}}$
$\mathbf{E}\|u(t)\|_{L^{2}}^{2}=e(t).$ (3.12)
There are infinitely many such solutions by choosing different energies $e$.
Proof.
We may thus start the iteration Proposition 3.2 with the pair $(v_{0},\mathring{R}_{0})$ and obtain a sequence of solutions $(v_{q},\mathring{R}_{q})$. By (3.8c), (3.10) and interpolation we deduce for any $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}b^{5}},\frac{1}{21{\cdot}b^{4}}\}}\right)\in(0,\frac{2}{3}\beta)$, the following series is summable
$\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\vartheta},2r}$ $\leqslant\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{1-\vartheta}_{C^{0},2r}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{\vartheta}_{C^{1},2r}$
$\lesssim\sum_{q\geqslant 0}\delta_{q+1}^{\frac{1-\vartheta}{2}}\lambda_{q+1}^{\frac{3}{2}\vartheta}\delta_{q+1}^{\frac{1}{2}\vartheta}\leqslant\sqrt{3r}La^{\frac{3}{2}b\vartheta}+\lambda_{2}^{\beta}\sum_{q\geqslant 1}\lambda_{q+1}^{\frac{3}{2}\vartheta-\beta}<\infty.$
Thus, we may define a limiting function $v=\lim_{q\to\infty}v_{q}$ which lies in $L^{2r}(\Omega;C({\mathbb{R}},C^{\vartheta}))$. Since $v_{q}$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted for every $q\in\mathbb{N}_{0}$, the limit $v$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted as well. Combining (3.7) with (4.33) in the proof of Proposition 3.2 and using interpolation again, we deduce for the same $\vartheta$ as above and any $p\geqslant 1$
$\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{\vartheta},p}$ $\leqslant\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{1-\vartheta}_{C^{0},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{\vartheta}_{C^{1},p}$
$\lesssim p\sum_{q\geqslant 0}(\lambda_{q+1}^{-\frac{\gamma}{8}\sigma}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})^{1-\vartheta}\lambda_{q+2}^{\frac{\gamma}{8}\vartheta}$
$\lesssim p\sum_{q\geqslant 1}\left(\lambda_{q+1}^{\frac{\gamma}{8}(b\vartheta+\sigma\vartheta-\sigma)}+\lambda_{q+1}^{\gamma(\frac{9}{8}b\vartheta+\frac{1}{8}-b)}\right)<\infty,$
since $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{b+\sigma},\frac{8b-1}{9b}\}}$. Then we obtain $\lim_{q\to\infty}z_{q}=z$ in $L^{p}(\Omega;C({\mathbb{R}},C^{\vartheta}))$ for any $p\geqslant 1$. Furthermore, it follows from (3.8e) that $\lim_{q\to\infty}\mathring{R}_{q}=0$ in $L^{1}(\Omega;C({\mathbb{R}},C^{0}))$. Thus $v$ is an analytically weak solution to (3.2). Hence letting $u=v+z$ we obtain an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted analytically weak solution to (1.1). Moreover, the estimate (3.11) for $u$ holds. Finally, (3.12) follows from (3.9), and choosing different energy functions, we can obtain infinitely many such solutions. This concludes the proof of Theorem 3.3. ∎
4. Proof of Proposition 3.2
The proof proceeds in several main steps which are similar as in many convex integration schemes. First of all, we start the construction by fixing the parameters in Section 4.1 and proceed with a mollification step in Section 4.2. Section 4.3 is devoted to introduce the new iteration $v_{q+1}$. This is the main part of the construction which differs in general convex integration schemes since we aim to construct global Hölder continuous flows. Here, we combine the construction methods of [BV19] and [HZZ22b] to obtain a new perturbation $w_{q+1}$. More precisely, we construct new amplitude functions $a_{(\xi)}$ similarly to [HZZ22b], meanwhile, we incorporate the solutions for transport equations into Beltrami waves to obtain an acceptable transport error similar to [BV19]. As pathwise positive lower bounds of the solutions to the transport equations are required, we use cutoff method to control the growth of noise and add pathwise estimtes of the velocity field $v_{q}$ in inductive iteration. Section 4.4 contains the inductive estimates of $v_{q+1}$, especially the pathwise and moment estimates. In Section 4.5, we show how to construct the new stress $\mathring{R}_{q+1}$. We establish the inductive pathwise and moment estimates on $\mathring{R}_{q+1}$ in Section 4.6. Finally, in Section 4.7 we show how the energy is controlled.
4.1. Choice of parameters
In the sequel, additional parameters will be indispensable and their value has to be carefully chosen in order to respect all the compatibility conditions appearing in the estimations below. First, for a fixed integer $b\geqslant 7$, we let $\beta,\gamma\in(0,1)$ be small parameters satisfying
$4b^{3}\gamma+4b^{3}\beta<1,\qquad 3b\beta<\gamma.$
In the sequel, we also use the following bound
$2b^{2}\beta+\beta<\frac{\gamma\sigma}{8},$
which can be obtained by choosing $\gamma=\frac{2}{9b^{3}}$ and choosing $\beta$ small such that $\beta<{\mathord{{\rm min}}}{\{\frac{\sigma}{78{\cdot}b^{5}},\frac{2}{27{\cdot}b^{4}}\}}$. The last free parameter $a\in 2^{36b^{2}\mathbb{N}}$ is chosen sufficiently large such that $a^{\frac{b\gamma}{8}}\in\mathbb{N}$ and $2<a^{(\frac{b}{4}-1)b^{2}\gamma}$. The choice of $a$ ensures $f(q)\in\mathbb{N}$ and (3.6) holds. The above requirements for parameters will be used in the following sections. In the sequel, we increase $a$ in order to absorb various implicit and universal constants.
4.2. Mollification
We replace $v_{q}$ by a mollified velocity field $v_{\ell}$, we choose a small parameter $\ell\in(0,1)$ as $\ell=\lambda_{q}^{-2}$. Let $\{\phi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers on $\mathbb{R}^{3}$, and $\{\varphi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers with support in $(0,1)$. The one-sided mollifiers here is used in order to preserve adaptedness. We define a mollification of $v_{q}$ and $\mathring{R}_{q}$ in space and time by
$v_{\ell}=(v_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad\mathring{R}_{\ell}=(\mathring{R}_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad z_{\ell}=({z_{q}}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$
where $\phi_{\ell}=\frac{1}{\ell^{3}}\phi(\frac{\cdot}{\ell})$ and $\varphi_{\ell}=\frac{1}{\ell}\varphi(\frac{\cdot}{\ell})$. By definition it is easy to know that $z_{\ell}$ is adapted. Then using (3.4) we obtain that $(v_{\ell},\mathring{R}_{\ell})$ satisfies
$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{com}}),$ (4.1)
$\mathord{{\rm div}}v_{\ell}$ $=0,$
where
$R_{\textrm{com}}=(v_{\ell}+z_{\ell})\mathring{\otimes}(v_{\ell}+z_{\ell})-(((v_{q}+z_{q})\mathring{\otimes}(v_{q}+z_{q}))*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$ (4.2)
$p_{\ell}=(p_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell}-\frac{1}{3}\big{(}|v_{\ell}+z_{\ell}|^{2}-(|v_{q}+z_{q}|^{2}*_{x}\phi_{\ell})*_{t}\varphi_{\ell}).$
Using the definition of $z_{q}$ (3.5) and the pathwise inductive assumption (3.8a), we can obtain the following pathwise estimate for $R_{\textrm{com}}$
$\|R_{\textrm{com}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q}+z_{q}\|^{2}_{C^{0}_{[t-1,t+1],x}}\leqslant 4(\lambda_{q+2}^{2\gamma}+\lambda_{q+3}^{\frac{\gamma}{2}})\leqslant\frac{1}{10}\lambda_{q+4}^{\gamma},$ (4.3)
where we choose $a$ sufficiently large to absorb the constant in the last inequality. Using a mollification estimate and inductive assumption (3.8a) and (3.8c), for any $t\in\mathbb{R}$ and $\delta\in(0,\frac{1}{12})$, we have
$\|R_{\textrm{com}}(t)\|_{L^{\infty}}$
$\lesssim\left(\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}+\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C_{x}^{1}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{\frac{1}{2}-\delta}_{[t-1,t]}C_{x}^{0}}\right)\left(\|z_{q}\|_{C^{0}_{[t-1,t],x}}+\|v_{q}\|_{C^{0}_{[t-1,t],x}}\right).$
Taking expectation and using (3.7), (3.8a) and (3.8b), we obtain
$|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{C^{0},r}$ $\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell|\!|\!|z_{q}|\!|\!|_{C^{1},2r}+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2r}\right)\left(|\!|\!|z_{q}|\!|\!|_{C^{0},2r}+|\!|\!|v_{q}|\!|\!|_{C^{0},2r}\right)$ (4.4)
$\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell\lambda_{q+1}^{\frac{\gamma}{8}}2rL+\ell^{\frac{1}{2}-\delta}\lambda_{q+1}^{\frac{\gamma}{8}}2pL^{2}\right)\left(\lambda_{q}^{\beta}+2rL\right)$
$\lesssim\lambda_{q}^{-\frac{1}{2}+\beta}+\lambda_{q}^{-\frac{3}{4}+\beta}\lambda_{q+1}^{\frac{\gamma}{8}}$
$\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
where we used $b\gamma+4b^{3}\beta<1$ to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{-\frac{3}{4}+\beta+\frac{1}{8}b\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $a$ was chosen sufficiently large to absorb the constant. By (4.4), we also obtain
$\qquad|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$ (4.5)
Similarly, by inductive assumption (3.8c) and the same requirements for parameters as above, we also obtain for any $t\in\mathbb{R}$
$\|(v_{q}-v_{\ell})(t)\|_{L^{\infty}}\lesssim\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}\leqslant\ell\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (4.6)
4.3. Perturbation of the velocity
As usual in convex integration schemes, in order to guarantee smoothness throughout the construction, the new velocity field $v_{q+1}$ will be constructed as a perturbation of $v_{\ell}$:
$v_{q+1}=:v_{\ell}+w_{q+1},$
where $w_{q+1}:=w_{q+1}^{(p)}+w_{q+1}^{(c)}$ is constructed following [BV19]. We proceed construction with introducing transport equations and time cutoffs.
4.3.1. Flow maps and cutoffs
In order to have an acceptable transport error in the estimates of $\mathring{R}_{q+1}$, the perturbation $w_{q+1}$ needs to be transported by the flow of the vector field $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$. Similar to [BV19], a natural way to achieve this, is to replace the linear phase $\xi\cdot x$ in the definition of the Beltrami wave $W_{\xi,\lambda}$ by the nonlinear phase $\xi\cdot\Phi(x,t)$, where $\Phi$ is transported by the aforementioned vector field. A slight difference from [BV19] is that we extend the definition of $\Phi$ to the whole line w.r.t time variable. For any integer $k\in\mathbb{Z}$, we subdivide $[k,k+1]$ into time intervals of size $\ell$ and solve transport equations on these intervals. For $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$ we define the adapted map $\Phi_{k,j}:\Omega\times{{\mathbb{R}}^{3}}\times[k+(j-1)\ell,k+(j+1)\ell]\to{{\mathbb{R}}^{3}}$ as the ${\mathbb{T}}^{3}$ periodic solution of
$(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\Phi_{k,j}$ $=0,$ (4.7)
$\Phi_{k,j}(k+(j-1)\ell,x)$ $=x.$
We put further details on the $C_{x}^{n}$-estimates of $\Phi_{k,j}$ in Appendix C. Specifically, the following two estimates which follows from (C.5) and (C.6)
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\textrm{Id}\|_{C_{x}^{0}}\leqslant\lambda^{-\frac{1}{2}}_{q}\ll 1,$ (4.8a)
$\frac{1}{2}\leqslant$ $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\leqslant 2.$ (4.8b)
As indicated by the proof in Appendix C, the validity of (4.8a) requires that both $v_{q}$ and $z_{q}$ have pointwise bounds, and this is why we need to cut off $\tilde{z}_{q}$ and introduce (3.8c) in the inductive iteration, instead of solely considering the moment bounds in the inductive iteration as in [HZZ22b]. The above estimates play a crucial role in the cancellation of the estimates for the new Reynolds stress $\mathring{R}_{q+1}$. In particular, (4.8b) ensures the validity of stationary phase Lemma 4.2 below in Section 4.6, which requires the positive lower bounds of $\nabla\Phi_{k,j}$. We could easily extend $\Phi_{k,j}$ from $\Omega\times\mathbb{R}^{3}\times[k+(j-1)\ell,k+(j+1)\ell]$ to $\Omega\times\mathbb{R}^{3}\times[k,k+1]$ in an adapted way, which is still denoted as $\Phi_{k,j}$. The extension is not unique but in the following we only use the value of $\Phi_{k,j}$ on $[k+(j-1)\ell,k+(j+1)\ell]$. We also let $\eta$ be a non-negative function supported in $(-1,1)$, which is 1 on $(-\frac{1}{4},\frac{1}{4})$ and such that the square of the shifted bump functions
$\eta_{j}(t)=\eta(\ell^{-1}t-j)$
form a partition of unity
$\sum_{j}\eta^{2}_{j}(t)=1,$
[…]
$\eta_{k,j}(t)=\eta(\ell^{-1}(t-k)-j)$
form a partition of unity
$\sum_{j}\eta^{2}_{k,j}(t)=1,$ (4.9)
for all $t\in[k,k+1]$.
4.3.2. Construction of $w_{q+1}$
Let us now proceed with the construction of the perturbation $w_{q+1}$. To this end, the building blocks of the perturbation are the Beltrami waves presented in [BV19, section 5.4], which we recall in Appendix A. Since we use moment bounds of $\mathring{R}_{q}$ to the iterative estimates in Proposition 3.2, we have to adjust the amplitude functions $a_{(\xi)}$ such that we can apply Lemma A.2 in our setting. More precisely, we first define $\rho$ as follows
$\rho(t,x):=\sqrt{\ell^{2}+|\mathring{R}_{\ell}(t,x)|^{2}}+c_{*}\zeta_{\ell}(t),$ (4.10)
$\zeta_{q}(t):=\frac{1}{3{\cdot}(2\pi)^{3}}\Big{[}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q}(t)+{z_{q}}(t)\|_{L^{2}}^{2}\Big{]},$ (4.11)
and
$\zeta_{\ell}:=\zeta_{q}*_{t}\varphi_{\ell},$
where $c_{*}>0$ is a constant introduced in Lemma A.2 and we take the constant $c_{R}$ in (3.8f) satisfying $c_{R}<c_{*}$. Now, we define the amplitude functions
$a_{(\xi)}(t,x)=a_{q+1,j,\xi}(t,x)=c_{*}^{-\frac{1}{2}}\rho^{\frac{1}{2}}\eta_{k,j}(t)\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right),\ t\in[k,k+1],$ (4.12)
where $\gamma_{\xi}^{(j)}$ is introduced in Appendix A. By the definition of $\rho$ we know
$\left\|\textrm{Id}-\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}(t,x)}{\rho}\right)\right\|_{C^{0}}\leqslant c_{*}.$
[…]
$c_{*}^{-1}\rho\textrm{Id}-\mathring{R}_{\ell}=\frac{1}{2}\sum_{j}\sum_{\xi\in\Lambda_{j}}a^{2}_{(\xi)}(\textrm{Id}-\xi\otimes\xi)$ (4.13)
[…]
$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (4.14)
and
$\|a_{(\xi)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma},\qquad\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (4.15)
With these preparations in hand, we define the principal part $w_{q+1}^{(p)}$ of the perturbation $w_{q+1}$ as in [BV19], more precisely, for $t\in[k,k+1]$, we define
$w_{(\xi)}(t,x):=a_{q+1,j,\xi}(t,x)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)},$
and
$w^{(p)}_{q+1}(t,x):$ $=\sum_{j}\sum_{\xi\in\Lambda_{j}}a_{q+1,j,\xi}(t,x)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)}$ (4.16)
$=\sum_{j}\sum_{\xi\in\Lambda_{j}}c_{*}^{-\frac{1}{2}}\rho^{\frac{1}{2}}\eta_{k,j}(t)\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)},$
where $B_{\xi}$ is suitable vector defined in Appendix A. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(p)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. Next, we define the incompressibility correction $w_{q+1}^{(c)}$. We aim to add a corrector to $w_{q+1}^{(p)}$ such that the resulting function is perfect $\mathrm{curl}$, making it thus divergence free. To this end, it is useful to introduce the following scalar $phase$ function
$\phi_{(\xi)}(t,x):=e^{i\lambda_{q+1}\xi\cdot(\Phi_{k,j}(t,x)-x)},\quad t\in[k,k+1].$ (4.17)
Using the same notation as in Appendix A, we define
$W_{(\xi)}=B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}.$
Since from Appendix A $\mathrm{curl}$ $W_{(\xi)}=\lambda_{q+1}W_{(\xi)}$, and $\phi_{(\xi)}$ and $a_{(\xi)}$ are scalar functions, we have
$a_{(\xi)}\phi_{(\xi)}W_{(\xi)}=\frac{1}{\lambda_{q+1}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right)-\frac{1}{\lambda_{q+1}}\nabla\left(a_{(\xi)}\phi_{(\xi)}\right)\times W_{(\xi)}.$
We therefore define
$w_{(\xi)}^{(c)}(t,x):$ $=\frac{1}{\lambda_{q+1}}\nabla(a_{(\xi)}\phi_{(\xi)})\times B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}$ (4.18)
$=\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}(t,x)-\textrm{Id})\xi\right)\times W_{(\xi)}(\Phi_{k,j}(t,x)),\ t\in[k.k+1].$
The incompressibility correction $w^{(c)}_{q+1}$ is defined as
$w^{(c)}_{q+1}(t,x)=\sum_{j}\sum_{\xi\in\Lambda_{j}}w_{(\xi)}^{(c)}(t,x)$
for all $t\in[k,k+1]$, $k\in\mathbb{Z}$. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(c)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted, and the new perturbation is defined as
$w_{q+1}=w^{(p)}_{q+1}+w^{(c)}_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j}\sum_{\xi\in\Lambda_{j}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right),$ (4.19)
and so clearly $w_{q+1}$ is mean zero, divergence free and $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. We define the velocity field at level $q+1$ as
$v_{q+1}:=v_{\ell}+w_{q+1}.$
Thus, it is also $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted.
4.4. Inductive estimates for $v_{q+1}$
In this section, we verify the inductive estimates (3.8a), (3.8b), (3.8c) and (3.10). By the definition of $w^{(p)}_{q+1}$ (4.16), it follows from (3.8d), (4.10) and (A.5) that
$\|w^{(p)}_{q+1}\|_{C^{0}_{t,x}}\lesssim 2|\Lambda_{j}|c_{*}^{-\frac{1}{2}}M(\|\mathring{R}_{q}\|^{\frac{1}{2}}_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}})+2|\Lambda_{j}|M\bar{e}\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\frac{3}{4}\gamma},$ (4.20)
where $M$ is the universal constant in (A.5). Taking expectation on the both side of (4.20), we then use (3.8e) to obtain
$|\!|\!|w^{(p)}_{q+1}|\!|\!|_{C^{0},2r}\lesssim 2|\Lambda_{j}|c_{*}^{-\frac{1}{2}}M(|\!|\!|\mathring{R}_{q}|\!|\!|_{C^{0},r}^{\frac{1}{2}}+\ell^{\frac{1}{2}})+2|\Lambda_{j}|M\bar{e}\delta_{q+1}^{\frac{1}{2}}\leqslant\frac{1}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$ (4.21)
where $\bar{M}$ is a universal constant satisfying $9c_{*}^{-\frac{1}{2}}M|\Lambda_{j}|+8M\bar{e}|\Lambda_{j}|<\bar{M}$ and we choose $a$ sufficiently large to absorb the constant in the above two inequalities. Applying (4.8b), (4.15) and (C.8), we also obtain
$\|w^{(p)}_{q+1}\|_{C^{1}_{t,x}}$ $\lesssim\sup_{j}\sum_{\xi\in\Lambda_{j}}\|a_{(\xi)}\|_{C^{1}_{t,x}}+\lambda_{q+1}\|a_{(\xi)}\|_{C^{0}_{t,x}}\left(\|\partial_{t}\Phi_{k,j}\|_{C^{0}}+\|\nabla\Phi_{k,j}\|_{C^{0}}\right)$ (4.22)
$\lesssim\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}+\lambda_{q+3}^{2\gamma}\lambda_{q+1}\leqslant\frac{1}{4}\lambda_{q+1}^{\frac{3}{2}}\delta_{q+1}^{\frac{1}{2}},$
where we used $b\geqslant 7$ and $12b^{3}\gamma+b\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma+4}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}b-b\beta{\tiny}}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant. We proceed to estimate $w^{(c)}_{q+1}$, using a standard mollification estimate, (4.8a) and (4.15) implies
$\|w^{(c)}_{q+1}\|_{C^{0}_{t,x}}$ $\lesssim\sup_{j}\sum_{\xi\in\Lambda_{j}}\left(\frac{\|\nabla a_{(\xi)}\|_{C^{0}_{t,x}}}{\lambda_{q+1}}+\|a_{(\xi)}\|_{C^{0}_{t,x}}\cdot\|\nabla\Phi_{k,j}-\textrm{Id}\|_{C^{0}}\right)$ (4.23)
$\lesssim\frac{\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{\gamma}\lambda_{q}^{-\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}},$
where we used $b\geqslant 7$ and $12b^{3}\gamma+2b\beta<$ 1 to have $\lambda_{q}^{b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\beta}$ and $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-b\beta}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant. Taking expectation we obtain
$|\!|\!|w_{q+1}^{(c)}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (4.24)
Using (4.8a), (4.8b), (4.14), (4.15), (4.23) and estimates (C.7)–(C.9) for $\Phi_{k,j}$, we obtain
$\|w^{(c)}_{q+1}\|_{C_{t,x}^{1}}\lesssim$ $\lambda_{q+1}\|w^{(c)}_{q+1}\|_{C^{0}_{t,x}}\sup_{j}(\|\partial_{t}\Phi_{k,j}\|_{C^{0}}+\|\nabla\Phi_{k,j}\|_{C^{0}})$ (4.25)
$+\sup_{j}\sum_{\xi\in\Lambda_{j}}\bigg{[}\frac{\|\nabla a_{(\xi)}\|_{C^{1}_{t,x}}}{\lambda_{q+1}}+\|a_{(\xi)}\|_{C^{1}_{t,x}}\|\nabla\Phi_{k,j}-\textrm{Id}\|_{C^{0}}$
$\qquad\qquad\quad+\|a_{(\xi)}\|_{C^{0}_{t,x}}(\|\nabla^{2}\Phi_{k,j}\|_{C^{0}}+\|\partial_{t}\nabla\Phi_{k,j}\|_{C^{0}})\bigg{]}$
$\lesssim$ $\lambda_{q+1}\lambda_{q+3}^{\gamma}\delta_{q+1}^{\frac{1}{2}}+\frac{\lambda_{q}^{9}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}}+\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}\lambda_{q}^{-\frac{1}{2}}+\lambda_{q+3}^{\gamma}(\ell^{-\frac{3}{4}}+\lambda_{q}^{\frac{3}{2}}\lambda_{q+3}^{\gamma})$
$\leqslant$ $\lambda_{q+1}\lambda_{q+3}^{\gamma}\delta_{q+1}^{\frac{1}{2}}+\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}+\frac{\lambda_{q}^{9}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}}\leqslant\frac{1}{4}\lambda_{q+1}^{\frac{3}{2}}\delta_{q+1}^{\frac{1}{2}},$
where we used $b\geqslant 7$ and $12b\gamma^{3}+b\beta<$ 1 to have $\lambda_{q}^{b^{3}\gamma}<\lambda_{q}^{\frac{b}{2}}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{9+3b^{3}\gamma}<\lambda_{q}^{\frac{5}{2}b-b\beta}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant.
By (4.21) and (4.24), we have
$|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\frac{3}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$
[…]
$|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$
Combining the above two estimates, we obtain
$|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{0},2r}\leqslant|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}+|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\bar{M}\delta_{q+1}^{\frac{1}{2}}.$
Hence, (3.10) holds at the level $q+1$. Moreover, it follows from (3.8a), (4.20) and (4.23) that
$\|v_{q+1}\|_{C^{0}_{t,x}}\leqslant\|v_{\ell}\|_{C^{0}_{t,x}}+\|w_{q+1}^{(p)}\|_{C^{0}_{t,x}}+\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+2}^{\gamma}+\lambda_{q+3}^{\frac{3}{4}\gamma}+\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\gamma}.$ (4.26)
By (3.8c), (4.22) and (4.25) we obtain
$\|v_{q+1}\|_{C^{1}_{t,x}}\leqslant\|v_{\ell}\|_{C^{1}_{t,x}}+\|w_{q+1}^{(p)}\|_{C^{1}_{t,x}}+\|w_{q+1}^{(c)}\|_{C^{1}_{t,x}}\leqslant\lambda^{\frac{3}{2}}_{q}\delta^{\frac{1}{2}}_{q}+\frac{1}{2}\lambda^{\frac{3}{2}}_{q+1}\delta^{\frac{1}{2}}_{q+1}\leqslant\lambda^{\frac{3}{2}}_{q+1}\delta^{\frac{1}{2}}_{q+1}.$
By (3.8b), (4.21) and (4.24) we obtain
$|\!|\!|v_{q+1}|\!|\!|_{C^{0},2r}\leqslant|\!|\!|v_{q}|\!|\!|_{C^{0},2r}+|\!|\!|w^{(p)}_{q+1}|\!|\!|_{C^{0},2r}+|\!|\!|w^{(c)}_{q+1}|\!|\!|_{C^{0},2r}\leqslant\lambda_{q}^{\beta}+\bar{M}\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+1}^{\beta}.$ (4.27)
Therefore, we have verified that (3.8a), (3.8b) and (3.8c) hold at the level $q+1$.
4.5. Definition of the Reynolds stress $\mathring{R}_{q+1}$
[…]
$\mathord{{\rm div}}\mathring{R}_{q+1}-\nabla p_{q+1}$ (4.28)
$=\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}}_{\mathord{{\rm div}}(R_{\textrm{transport}})}+\underbrace{\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}}$
$+\underbrace{w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{Nash}})}+\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}+\mathord{{\rm div}}(w_{q+1}^{(c)}\otimes w_{q+1}+w_{q+1}^{(p)}\otimes w_{q+1}^{(c)})}_{\mathord{{\rm div}}(R_{\textrm{corrector}})+\nabla p_{\textrm{corrector}}}$
$+\underbrace{\mathord{{\rm div}}(v_{q+1}\otimes z_{q+1}-v_{q+1}\otimes z_{\ell}+z_{q+1}\otimes v_{q+1}-z_{\ell}\otimes v_{q+1}+z_{q+1}\otimes z_{q+1}-z_{\ell}\otimes z_{\ell})-z_{q+1}+z_{\ell}}_{\mathord{{\rm div}}(R_{\textrm{commutator1}})+\nabla p_{\textrm{commutator1}}}$
$+\mathord{{\rm div}}(R_{\textrm{commutator}})-\nabla p_{\ell}.$
[…]
$R_{\textrm{transport}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}\right),$
$R_{\textrm{Nash}}$ $:=\mathcal{R}(w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})),$
$R_{\textrm{corrector}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right)+\left(w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}\right),$
$R_{\textrm{commutator1}}$ $:=v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})+(z_{q+1}-z_{\ell})\mathring{\otimes}v_{q+1}+z_{q+1}\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}-\mathcal{R}(z_{q+1}-z_{\ell}),$
$\nabla p_{\textrm{corrector}}$ $:=\frac{1}{3}(2w_{q+1}^{(c)}\cdot w_{q+1}^{(p)}+|w_{q+1}^{(c)}|^{2}),$
$\nabla p_{\textrm{commutator1}}$ $:=\frac{1}{3}(v_{q+1}\cdot z_{q+1}-v_{q+1}\cdot z_{\ell}+z_{q+1}\cdot v_{q+1}-z_{\ell}\cdot v_{q+1}+z_{q+1}\cdot z_{q+1}-z_{\ell}\cdot z_{\ell}).$
In order to define the remaining oscillation error from the second line in (4.28), first note that for $j,j^{\prime}$ such that $|j-j^{\prime}|\geqslant 2$, we have $\eta_{k,j}(t)\eta_{k,j^{\prime}}(t)=0$ Second, we have $\Lambda_{j}\cap\Lambda_{j^{\prime}}=\emptyset$ for $|j-j^{\prime}|=1$, and by Lemma A.1 we have $\mathord{{\rm div}}(W_{\xi}\otimes W_{\xi^{\prime}}+W_{\xi^{\prime}}\otimes W_{\xi})=\nabla(W_{\xi}\cdot W_{\xi^{\prime}})$. Therefore, we may use (4.13) and Lemma A.2 to obtain
$\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})$ $=\mathord{{\rm div}}\left(\sum_{j,j^{{}^{\prime}},\xi,\xi^{{}^{\prime}}}w_{(\xi)}\otimes w_{(\xi^{{}^{\prime}})}+\mathring{R}_{\ell}\right)$ (4.29)
$=\mathord{{\rm div}}\left(\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell}\right)^{2}B_{\xi}\otimes B_{-\xi}+\mathring{R}_{\ell}\right)$
$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
$=\mathord{{\rm div}}\left(c^{-1}_{*}\rho(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell})+\mathring{R}_{\ell}\right)$
$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
=$\nabla\left(c_{*}^{-1}\rho\right)+\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})})\nabla\left(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\right)$
$\qquad\qquad\qquad+\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\nabla\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)$
$=\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}.$
Note that compared to [BV19], a slight difference in this term is the appearance of $\mathord{{\rm div}}(\rho\mathrm{Id})$, which can be put into the pressure term $p_{\textrm{oscillation}}$. Thus, we have denoted
$p_{\textrm{oscillation}}:=\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)+c_{*}^{-1}\rho,$
and
$R_{\textrm{oscillation}}:=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}\left(\left(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})}-\frac{W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}}{2}\mathrm{Id}\right)\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})})\right).$
Finally we define the Reynolds stress at the level $q+1$ by
$\mathring{R}_{q+1}=R_{\textrm{transport}}+R_{\textrm{oscillation}}+R_{\textrm{Nash}}+R_{\textrm{corrector}}+R_{\textrm{commutator}}+R_{\textrm{commutator1}},$ (4.30)
and
$p_{q+1}=p_{\ell}-p_{\textrm{oscillation}}-p_{\textrm{corrector}}-p_{\textrm{commutator1}}.$
4.6. Inductive estimates for $\mathring{R}_{q+1}$
In this section, we shall verify (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$. We first proceed with estimates of $R_{\mathrm{commutator1}}$ in subsection 4.6.1. Then we recall a useful sationary phase lemma in subsection 4.6.2. Finally, the remianing errors are controlled in subsection 4.6.34.6.6
4.6.1. Estimate on $R_{\mathrm{commutator1}}$
First, by the definition of $z_{q}$ we have
$\|z_{q}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+3}^{\frac{\gamma}{4}}.$ (4.31)
Then we use the pathwise estimates (4.26) and (4.31) to obtain the following pathwise bounds
$\|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}$ $\leqslant\|v_{q+1}\|_{C^{0}_{t,x}}\|z_{q+1}-z_{\ell}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+3}^{\gamma}\lambda_{q+4}^{\frac{\gamma}{4}}+\lambda_{q+3}^{\frac{5}{4}\gamma}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
$\|z_{q+1}\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}\|_{C^{0}_{t,x}}$ $\leqslant\|z_{q+1}\|^{2}_{C^{0}_{t,x}}+\|z_{\ell}\|^{2}_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\frac{\gamma}{2}}+\lambda_{q+3}^{\frac{\gamma}{2}}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
$\|\mathcal{R}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}$ $\leqslant\|z_{q+1}-z_{\ell}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\frac{\gamma}{4}}+\lambda_{q+3}^{\frac{\gamma}{4}}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
where we chose $a$ sufficiently large to absorb the constant in the above. Summing over the estimates above, the following pathwise bound holds true for any $t\in\mathbb{R}$
$\|R_{\textrm{commutator1}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}+\|z_{q+1}$$\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}\|_{C^{0}_{t,x}}$ (4.32)
$+\|\mathcal{R}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}\leqslant\frac{4}{5}\lambda_{q+4}^{\gamma}.$
Next we estimate the $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of $R_{\textrm{com1}}$. Before proceeding, we give an estimate of $z_{q+1}-z_{q}$.
Lemma 4.1.
For any $p\geqslant 1$, we have
$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},p}\lesssim(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})p.$ (4.33)
In particular, for the fixed $r>1$ in Proposition 3.2, we have
$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})r,$ (4.34)
where the implicit constants are independent of $q$.
Proof.
By the definition of $z_{q}$, we deduce for any $p\geqslant 1$
$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},p}$ $=\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|z_{q+1}(s)-z_{q}(s)\|_{C_{x}^{0}}^{p}\right)^{\frac{1}{p}}$
$\leqslant\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q+1}(s)-\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{p}\cdot\chi_{q+1}(\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}})^{p}\right)^{\frac{1}{p}}$
$+\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{p}\cdot\left|\chi_{q+1}(\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}})-\chi_{q}(\|\tilde{z}_{q}(s)\|_{C_{x}^{1}})\right|^{p}\right)^{\frac{1}{p}}$
[…]
$\tilde{z}_{q+1}(t,x)-\tilde{z}_{q}(t,x)=\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}e^{ik\cdot x}\hat{z}(t,k),$
where $\hat{z}$ is the Fourier transform of $z$, and $k\in\mathbb{Z}^{3}$. Then using Hölder’s inequality, we deduce for any $t\in{\mathbb{R}}$ and previous $\sigma>0$,
$\|\tilde{z}_{q+1}(t)-\tilde{z}_{q}(t)\|_{L^{\infty}}$ $\leqslant\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}|\hat{z}(t,k)|$
$=\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}(1+|k|^{2})^{\frac{1}{2}(\frac{3}{2}+\sigma)}|\hat{z}(t,k)|(1+|k|^{2})^{-\frac{1}{2}(\frac{3}{2}+\sigma)}$
$\leqslant\left(\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}(1+|k|^{2})^{\frac{3}{2}+\sigma}|\hat{z}(t,k)|^{2}\right)^{\frac{1}{2}}\left(\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}(1+|k|^{2})^{-(\frac{3}{2}+\sigma)}\right)^{\frac{1}{2}}$
$\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}\|z(t)\|_{H^{\frac{3}{2}+\sigma}}.$
Taking expectation and using (3.7) we obtain the bound
[…]
$\lesssim$ $\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|z(s)\|_{H^{\frac{3}{2}+\sigma}}^{p}\right)^{\frac{1}{p}}\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2p}\leqslant\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}\sqrt{2p}L.$
[…]
$|\chi_{q+1}$ $(\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}})-\chi_{q}(\|\tilde{z}_{q}(s)\|_{C_{x}^{1}})|$
$\leqslant$ $\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}\leqslant\lambda_{q+2}^{\gamma},\ \|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}}>\lambda_{q+3}^{\gamma}\}}+\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}>\lambda_{q+2}^{\gamma}\}}$
$\leqslant$ $\mathbf{1}_{\{\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}}>\lambda_{q+3}^{\gamma}\}}+\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}>\lambda_{q+2}^{\gamma}\}}.$
Using Hölder’s inequality, we obtain
[…]
$+\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{p}\cdot\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}>\lambda_{q+2}^{\gamma}\}}\right)^{\frac{1}{p}}$
$\leqslant\sup_{t\in\mathbb{R}}\mathbf{E}$ $\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{2p}\right)^{\frac{1}{2p}}\mathbf{P}\left(\|\tilde{z}_{q+1}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+3}^{\gamma}\right)^{\frac{1}{2p}}$
$+\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{2p}\right)^{\frac{1}{2p}}\mathbf{P}\left(\|\tilde{z}_{q}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+2}^{\gamma}\right)^{\frac{1}{2p}}.$
By Chebyshev’s inequality and (3.7), we have the following bound
$\sup_{t\in\mathbb{R}}\mathbf{P}\left(\|\tilde{z}_{q}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+2}^{\gamma}\right)^{\frac{1}{2p}}$ $\leqslant\lambda_{q+2}^{-\gamma}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|\tilde{z}_{q}\|^{2p}_{C_{[t,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2p}}$
$\lesssim\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2p}\leqslant\sqrt{2p}\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}L.$
Similarly as above, we have
$\sup_{t\in\mathbb{R}}\mathbf{P}\left(\|\tilde{z}_{q+1}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+3}^{\gamma}\right)^{\frac{1}{2p}}$ $\leqslant\lambda_{q+3}^{-\gamma}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|\tilde{z}_{q+1}\|^{2p}_{C_{[t,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2p}}$
$\lesssim\lambda_{q+3}^{-\gamma}\lambda_{q+2}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2p}\leqslant\sqrt{2p}\lambda_{q+3}^{-\gamma}\lambda_{q+2}^{\frac{\gamma}{8}}L.$
With the above estimates in hand, using (3.7) again, we obtain
[…]
Finally, combining (4.35) with (4.36) implies
[…]
In particular, taking $p=2r$ impies (4.33). Hence, the proof of Lemma 4.1 is completed. ∎
With Lemma 4.1 in hand, we proceed to estimate $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$, using estimates (4.27), (4.34) and Hölder’s inequality, we have
$|\!|\!|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}\leqslant|\!|\!|v_{q+1}|\!|\!|_{C^{0},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\beta}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}).$
Combining (3.3), (3.7) with (4.34) and using Hölder’s inequality we obtain
$|\!|\!|z_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|z_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}$
$\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2r}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}.$
Using a standard mollification estimate, we obtain for any $\delta\in(0,\frac{1}{12})$
$\|z_{\ell}(t)-z_{q}(t)\|_{L^{\infty}}\lesssim\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C^{1}_{x}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{1/2-\delta}_{[t-1,t]}C^{0}_{x}}.$
Taking expectation on the both side and using (3.7), we deduce
$|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}$ $\lesssim\ell\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z_{q}\|^{2r}_{C_{[t-1,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2r}}+\ell^{\frac{1}{2}-\delta}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z_{q}\|^{2r}_{C_{[t-1,t+1]}^{1/2-\delta}C_{x}^{0}}\right)^{\frac{1}{2r}}$ (4.37)
$\lesssim\ell^{\frac{5}{12}}\lambda_{q+1}^{\frac{\gamma}{8}}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z\|^{2r}_{C_{[t-1,t+1]}^{1/2-\delta}H^{\frac{3}{2}+\sigma}}\right)^{\frac{1}{2r}}\lesssim\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}\sqrt{2r}L.$
Hence, using (4.27), (4.37) and Hölder’s inequality, we deduce
$|\!|\!|v_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}\leqslant|\!|\!|v_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\beta+\gamma}\lambda_{q}^{-\frac{3}{4}}.$
Similarly, using (3.7) and (4.37) again, we deduce
$|\!|\!|z_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|z_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}$
$\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}.$
Summarizing over the estimates above, we obtain
$|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|v_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},r}$ (4.38)
$+|\!|\!|z_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},r}$
$\lesssim\lambda_{q+1}^{\beta+\gamma}\lambda_{q}^{-\frac{3}{4}}+\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}+\lambda_{q+1}^{\beta}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})$
$\leqslant\frac{1}{48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|\mathring{R}_{\textrm{commutator1}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e},$
where we used $2b^{2}\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta-\frac{b}{8}\gamma\sigma}<\lambda_{q}^{-2b^{3}\beta}$. We also used $3b\beta<\gamma$, $14b^{3}\beta+2b\gamma<$ 1 to have $\lambda_{q}^{b\beta+b\gamma-b^{2}\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{b\gamma+b\beta-\frac{3}{4}}<\lambda_{q}^{-2b^{3}\beta}$ in the above two inequalities and we chose $a$ sufficiently large to absorb the constant.
4.6.2. Stationary phase Lemma
Before we estimate the remaining terms in (4.30) separately to show that the stress $\mathring{R}_{q+1}$ obeys the bound (3.8d), (3.8e) and (3.8f) at level $q+1$, we first introduce the following lemma which makes rigorous the fact that $\mathcal{R}$ obeys the same elliptic regularity estimates as $|\nabla|^{-1}$. We recall the following stationary phase lemma (see for example [BV19, Lemma 5.7] and [DS17, Lemma 2.2]), adapted to our setting.
Lemma 4.2.
Given $\xi\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$, let $\lambda\xi\in\mathbb{Z}^{3}$ and $\alpha\in(0,1)$. Assume that $a\in C^{m,\alpha}(\mathbb{T}^{3})$ and $\Phi\in C^{m,\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$ are smooth functions such that the phase function $\Phi$ obeys
$C^{-1}\leqslant|\nabla\Phi|\leqslant C$
on $\mathbb{T}^{3}$, for some constant $C\geqslant 1$. Then, with the inverse divergence operator $\mathcal{R}$ defined in Section 2.1 we have for any $m\in\mathbb{N}$
$\left|\int_{\mathbb{T}^{3}}a(x)e^{i\lambda\xi\cdot\Phi(x)}{\mathord{{\rm d}}}x\right|\lesssim\frac{\|a\|_{C^{m}}+\|a\|_{C^{0}}\|\nabla\Phi\|_{C^{m}}}{\lambda^{m}},$
$\left\|\mathcal{R}\left(a(x)e^{i\lambda\xi\cdot\Phi(x)}\right)\right\|_{C^{\alpha}}\lesssim\frac{\|a\|_{C^{0}}}{\lambda^{1-\alpha}}+\frac{\|a\|_{C^{m,\alpha}}+\|a\|_{C^{0}}\|\nabla\Phi\|_{C^{m,\alpha}}}{{\lambda^{m-\alpha}}},$
where the implicit constant depends on $C,\alpha$ and m (in particular, not on the frequency $\lambda$).
The above lemma is used when estimating the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of the remaining terms in (4.30). Indeed, for a fixed $t\in[k+(j-1)\ell,k+(j+1)\ell]$, in view of the bound (4.8b), we have $\frac{1}{2}\leqslant\|\Phi_{k,j}(\cdot,t)\|_{C^{0}_{x}}\leqslant 2$ on $\mathbb{T}^{3}$. Thus, it follows from Lemma 4.2 that if $a$ is a smooth periodic function, then we have for any $n\in\mathbb{N}$
$\left\|\mathcal{R}(aW_{(\xi)}\circ\Phi_{k,j})\right\|_{C^{0}_{t}C^{\alpha}_{x}}\lesssim\frac{\|a\|_{C^{0}_{t,x}}}{\lambda_{q+1}^{1-\alpha}}+\frac{\|a\|_{C^{0}_{t}C^{n,\alpha}_{x}}+\|a\|_{C^{0}_{t,x}}\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C^{n,\alpha}_{x}}}{{\lambda_{q+1}^{n-\alpha}}}.$ (4.39)
In addition, the same proof that was used in [DS17] to prove Lemma 4.2 gives other useful estimates. Let $\xi\in\Lambda_{j}$ and $\xi^{\prime}\in\Lambda_{j^{\prime}}$ for $|j-j^{\prime}|\leqslant 1$ be such that $\xi+\xi^{\prime}\neq 0$, we have that $|\xi+\xi^{\prime}|\geqslant\bar{c}>0$, for some universal constant $\bar{c}\in(0,1)$. Then for a smooth periodic function $a$, we have the following estimates (see [BV19, (5.37)]), for $n\in\mathbb{N}$
$\left|\int_{\mathbb{T}^{3}}a(x)e^{i\lambda(\xi\cdot\Phi_{k,j}+\xi^{\prime}\cdot\Phi_{k,j^{\prime}})}{\mathord{{\rm d}}}x\right|\lesssim\frac{\|a\|_{C^{0}_{t}C^{n}_{x}}+\|a\|_{C^{0}_{t,x}}(\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C_{x}^{n}}+\|\nabla\Phi_{k,j^{\prime}}\|_{C^{0}_{t}C_{x}^{n}})}{\lambda^{n}},$ (4.40)
and
$\|\mathcal{R}(a(W_{(\xi)}\circ\Phi_{k,j}\otimes$ $W_{(\xi^{\prime})}\circ\Phi_{k,j^{\prime}}))\|_{C^{0}_{t}C_{x}^{\alpha}}$ (4.41)
$\lesssim\frac{\|a\|_{C^{0}_{t,x}}}{\lambda^{1-\alpha}}+\frac{\|a\|_{C^{0}_{t}C_{x}^{n,\alpha}}+\|a\|_{C^{0}_{t,x}}(\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C_{x}^{n,\alpha}}+\|\nabla\Phi_{k,j^{\prime}}\|_{C^{0}_{t}C_{x}^{n,\alpha}})}{{\lambda^{n-\alpha}}}.$
The implicit constant depends only on $\alpha$ and $n$. In the following subsection, we aim to use the above discussion to control the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$ norms, respectively, of the transport error, oscillation error, Nash error and corrector error. We first fix an integer $b\geqslant 7$, then choose $\beta$, $\alpha$ sufficiently small such that $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 and choose $m\in\mathbb{N}$ such that $m>\frac{2b+5}{b-5}$, at last we choose $a$ sufficiently large to absorb the constant.
4.6.3. Estimate on the transport error.
Inspecting the definition of $w_{q+1}^{(p)}$ as (4.16), we notice that the material derivative $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$ cannot lead to higher frequency term, since the term $W_{(\xi)}\circ\Phi_{k,j}$ is perfectly transported by $v_{\ell}+z_{\ell}$. Therefore, for $t\in[k,k+1]$, $k\in\mathbb{Z}$ we have
$(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}=\sum_{j,\xi}(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}W_{(\xi)}\circ\Phi_{k,j}.$
Returning to the $C^{N}_{t,x}$-norm estimate of $a_{(\xi)}$ in (4.14) and (4.15) and the definition of $z_{q}$ (3.5), we use chain rule and (3.8a) to obtain that for any $n\in\mathbb{N}_{0}$
$\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}\|_{C^{n}_{t,x}}$ $\lesssim\|a_{(\xi)}\|_{C^{n+1}_{t,x}}+\sum_{k=0}^{n}\|v_{\ell}+z_{\ell}\|_{C^{k}_{t,x}}\cdot\|\nabla a_{(\xi)}\|_{C^{n-k}_{t,x}}$ (4.42)
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+2)\gamma}+\sum_{k=0}^{n}\ell^{-k}\lambda_{q+3}^{\gamma}\ell^{-2(n-k+1)-1}\lambda_{q+3}^{(n-k+2)\gamma}$
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+3)\gamma}.$
[…]
$\|R_{\textrm{transport}}\|_{C^{0}_{t,x}}$ $\lesssim\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-4-m}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
$\lesssim\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)$
$\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $13b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
4.6.4. Estimate on the oscillation error.
Based on previous computations (4.29) in Section 4.5, we can rewrite oscillation error as follows, for $t\in[k,k+1]$, $k\in\mathbb{Z}$
$R_{\textrm{oscillation}}=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}$ $\bigg{[}\left(W_{(\xi)}\circ\Phi_{k,j}\otimes W_{(\xi^{{}^{\prime}})}\circ\Phi_{k,j^{{}^{\prime}}}-\frac{1}{2}(W_{(\xi)}\circ\Phi_{k,j}\cdot W_{(\xi^{{}^{\prime}})}\circ\Phi_{k,j^{{}^{\prime}}})\textrm{Id}\right)$
×$\times\left(\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\right)\bigg{]}.$
Then using (4.14), (4.15) and chain rule, we have that for $n\in\mathbb{N}_{0}$
$\|\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})\|_{C^{n}_{t,x}}$ $\lesssim\sum_{k=0}^{n+1}\|a_{(\xi)}\|_{C^{k}_{t,x}}\|a_{(\xi)}\|_{C^{n-k+1}_{t,x}}$ (4.43)
$\lesssim\sum_{k=0}^{n+1}\ell^{-2k-\frac{1}{2}}\lambda_{q+3}^{(k+1)\gamma}\ell^{-2(n-k+1)-\frac{1}{2}}\lambda_{q+3}^{(n-k+2)\gamma}$
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+3)\gamma}.$
Using (4.8a) and (4.15), we find
$\lambda_{q+1}\|a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\|_{C^{0}_{t,x}}\lesssim\lambda_{q+1}\lambda_{q}^{-\frac{1}{2}}\lambda_{q+3}^{2\gamma}.$ (4.44)
Using (4.8a), (4.14), (4.15), (4.43), (C.7) and chain rule, we obtain for $n\geqslant 1$
$\lambda_{q+1}\|a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\|_{C^{0}_{t}C^{n}_{x}}$ (4.45)
$\lesssim$ $\lambda_{q+1}\sum_{l=0}^{n}\|a_{(\xi)}a_{(\xi^{\prime})}\|_{C^{0}_{t}C^{l}_{x}}\|(\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}}\|_{C^{0}_{t}C^{n-l}_{x}}$
$\lesssim$ $\lambda_{q+1}\sum_{l=0}^{n}\ell^{-2l-1}\lambda_{q+3}^{(l+2)\gamma}\ell^{-(n-l)}\lesssim\lambda_{q+1}\ell^{-2n-1}\lambda_{q+3}^{(n+2)\gamma},$
where we used $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) in the second inequality, namely $\|\nabla\Phi_{k,j}(t)\|_{C^{n}_{x}}\leqslant\ell^{-n}$ for $n\geqslant 1$. Combining (4.43), (4.44) with (4.45) and using $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) again, applying (4.41) with $n=m$ and $a$ replaced by $\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})$ implies
$\|R_{\textrm{oscillation}}\|_{C^{0}_{t,x}}\lesssim$ $\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
$\qquad+\frac{\lambda_{q+1}\lambda_{q}^{-\frac{1}{2}}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-3}\lambda_{q+3}^{(m+3)\gamma}+\ell^{-m-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m-1-\alpha}}$
$\lesssim$ $\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-2}}\right)+\frac{\lambda_{q+3}^{2\gamma}}{\lambda_{q}^{\frac{1}{2}}\lambda_{q+1}^{-\alpha}}$
$\leqslant$ $\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\alpha-2b^{3}\beta}$, and used $5m+2b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-2)b}$, and we chose $a$ sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
4.6.5. Estimate on the Nash error.
Similar to transport error, we write
$w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})=w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell})+w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell}).$ (4.46)
For the first term on the right hand side, we have for $t\in[k,k+1]$, $k\in\mathbb{Z}$
$w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell})=\sum_{j,\xi}a_{(\xi)}W_{(\xi)}\circ\Phi_{k,j}\cdot\nabla(v_{\ell}+z_{\ell}).$
Similar to (4.42), using (3.8a), (4.14), (4.15), (4.31) and chain rule, we have for $n\in\mathbb{N}_{0}$
$\|a_{(\xi)}\nabla(v_{\ell}+z_{\ell})\|_{C^{n}_{t,x}}\lesssim\ell^{-2n-2}\lambda_{q+3}^{(n+2)\gamma}.$ (4.47)
[…]
$\|\mathcal{R}(w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}\lesssim\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-4}\lambda_{q+3}^{(m+3)\gamma}+\ell^{-m-3}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m-\alpha}}.$ (4.48)
The second term on the right hand side of (4.46) is
$w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell})=\sum_{j,\xi}\left(\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id})\xi\right)\times W_{(\xi)}\circ\Phi_{k,j}\right)\cdot\nabla(v_{\ell}+z_{\ell}).$
Similar as before, we use (3.8a), (4.14), (4.15) and (4.31) to have
$\left\|\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}\cdot\nabla(v_{\ell}+z_{\ell})\right\|_{C^{n}_{t,x}}$ $\lesssim\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\|\nabla a_{(\xi)}\|_{C^{k}_{t,x}}\|\nabla(v_{\ell}+z_{\ell})\|_{C^{n-k}_{t,x}}$
$\lesssim\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\ell^{-2k-3}\lambda_{q+3}^{(k+2)\gamma}\ell^{-(n-k+1)}\lambda_{q+3}^{\gamma}$
$\lesssim\frac{\ell^{-2n-4}}{\lambda_{q+1}}\lambda_{q+3}^{(n+3)\gamma},$
and we use (4.47) and (C.7) to have
$\|a_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id})\cdot\nabla(v_{\ell}+z_{\ell})\|_{C^{0}_{t}C^{n}_{x}}$ $\lesssim\sum_{l=0}^{n}\|a_{(\xi)}\nabla(v_{\ell}+z_{\ell})\|_{C^{0}_{t}C^{l}_{x}}\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C_{x}^{n-l}}$
$\lesssim\sum_{l=0}^{n}\ell^{-2l-2}\lambda_{q+3}^{(l+2)\gamma}\ell^{-(n-l)}$
$\lesssim\ell^{-2n-2}\lambda_{q+3}^{(n+2)\gamma}.$
[…]
$\|\mathcal{R}(w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}$ $\lesssim\frac{\ell^{-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2m-6}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-5}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m+1-\alpha}}$ (4.49)
$\qquad+\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-4}\lambda_{q+3}^{(m+3)\gamma}+\ell^{-m-3}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m-\alpha}}.$
Combining (4.48) with (4.49) we obatin
$\|R_{\textrm{Nash}}\|_{C^{0}_{t,x}}$ $\lesssim\|\mathcal{R}(w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}+\|\mathcal{R}(w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}$
$\lesssim\left(\frac{\ell^{-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}\right)\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $13b^{3}\gamma+3b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{4+2b^{3}\gamma+b\alpha}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+3b^{3}\gamma+b\alpha}<\lambda_{q}^{2b-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation, we obtain
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
4.6.6. Estimate on the corrector error.
The corrector error has two pieces, the transport derivative of $w^{(c)}_{q+1}$ by the flow of $v_{\ell}+z_{\ell}$, and the residual contribution from the nonlinear term, which are easier to estimate due to (4.20) and (4.23). Thus we obtain
$\|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\leqslant\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}^{2}\leqslant\frac{\lambda_{q}^{8}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-1}\leqslant\frac{1}{3{\cdot}6{\cdot}96}c_{R}\delta_{q+3},$
and
$\|w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma}\left(\frac{\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{\gamma}\lambda_{q}^{-\frac{1}{2}}\right)\leqslant\frac{1}{3{\cdot}6{\cdot}96}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-2b^{3}\beta}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{3}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Combining the above estimates and taking expectation, we obtain
$|\!|\!|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}|\!|\!|_{C^{0},r}$ $\leqslant\frac{1}{12}\delta_{q+2},$ (4.50)
$|\!|\!|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{6{\cdot}96}c_{R}\delta_{q+3}\underline{e}.$
The remaining term in corrector error is
$\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right).$
We use the definition of $w_{q+1}^{(c)}$ to have
$(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}=\sum_{j,\xi}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id})\xi\right)\right)\times W_{(\xi)}\circ\Phi_{k,j}.$
Using (3.8a), (4.14) and (4.15), we obtain
$\left\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}\right\|_{C^{n}_{t,x}}$ $\lesssim\frac{1}{\lambda_{q+1}}\|a_{(\xi)}\|_{C^{n+2}_{t,x}}+\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\|a_{(\xi)}\|_{C^{k+2}_{t,x}}\|v_{\ell}+z_{\ell}\|_{C^{n-k}_{t,x}}$
$\lesssim\frac{1}{\lambda_{q+1}}\ell^{-2n-5}\lambda_{q+3}^{(n+3)\gamma}+\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\ell^{-2k-5}\lambda_{q+3}^{(k+3)\gamma}\ell^{-(n-k)}\lambda_{q+3}^{\gamma}$
$\lesssim\frac{\ell^{-2n-5}}{\lambda_{q+1}}\lambda_{q+3}^{(n+4)\gamma}.$
Using (4.42) and (C.7), we obtain
$\left\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)(a_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id}))\right\|_{C^{0}_{t}C^{n}_{x}}$ $\lesssim\sum_{l=0}^{n}\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}\|_{C^{0}_{t}C^{l}_{x}}\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C^{n-l}_{x}}$
$\lesssim\sum_{l=0}^{n}\ell^{-2l-3}\lambda_{q+3}^{(l+3)\gamma}\ell^{-(n-l)}$
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+3)\gamma}.$
[…]
$\left\|\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right)\right\|_{C^{0}_{t,x}}\lesssim$ $\frac{\ell^{-5}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2m-7}\lambda_{q+3}^{(m+5)\gamma}+\ell^{-m-6}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{m+1-\alpha}}$
$+\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
$\lesssim$ $\left(\frac{\ell^{-5}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}\right)\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)$
$\leqslant$ $\frac{1}{6{\cdot}96}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $14b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{10+4b^{3}\gamma}<\lambda_{q}^{2b-\alpha b-2b^{3}\beta}$ and $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-\alpha b-2b^{3}\beta}$, and used $5m+5+b<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Taking expetation and combining (4.50), we obatin
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6{\cdot}48}c_{R}\delta_{q+3}\underline{e}.$
Summarizing over all the estimates we have established above and by (4.3), (4.4) and (4.32), we obtain
$\|\mathring{R}_{q+1}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\gamma},\quad|\!|\!|\mathring{R}_{q+1}|\!|\!|_{C^{0},r}\leqslant\delta_{q+2},\quad|\!|\!|\mathring{R}_{q+1}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e}.$ (4.51)
Hence, (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$.
4.7. Inductive estimates for the energy
To conclude the proof of Proposition 3.2, we shall verify (3.9) holds at level $q+1$. To this end, we define
$\delta E(t):=\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}.$
Proposition 4.3.
It holds for $t\in{\mathbb{R}}$
$\delta E(t)\leqslant$ $\frac{1}{4}\delta_{q+2}e(t).$ (4.52)
Proof.
By the definition of $\zeta_{q}$ in Section 4.3, we find
[…]
which shall be estimated. Let us begin with the bound of the first term on the right hand side of (4.53). Similar as the computation of oscillation error (4.29) in Section 4.5, by (4.13) we have for any $t\in[k,k+1]$, $k\in\mathbb{Z}$
$w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}$
$=\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)^{2}B_{\xi}\otimes B_{-\xi}+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}$
$=\sum_{j}c^{-1}_{*}\rho\eta_{k,j}^{2}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}$
$=c^{-1}_{*}\rho\textrm{Id}-\mathring{R}_{\ell}+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}.$
Taking the trace on both sides of the equation and using the fact that $\mathring{R}_{\ell}$ is traceless, we deduce for $t\in[k,k+1]$, $k\in\mathbb{Z}$
$|w_{q+1}^{(p)}|^{2}-3\zeta_{q}=3c_{*}^{-1}\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}+3(\zeta_{\ell}-\zeta_{q})+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}).$
Hence, we have
$\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(p)}|^{2}-3\zeta_{q}{\mathord{{\rm d}}}x\right|\leqslant$ $3c_{*}^{-1}(2\pi)^{3}\ell+3c_{*}^{-1}\mathbf{E}\|\mathring{R}_{\ell}\|_{L^{1}}+3\cdot(2\pi)^{3}|\zeta_{\ell}-\zeta_{q}|$ (4.54)
$+$ $\mathbf{E}\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}){\mathord{{\rm d}}}x\right|.$
In the following, we estimate each term separately, by the choice of parameters, we find
$3c_{*}^{-1}(2\pi)^{3}\ell=3c_{*}^{-1}(2\pi)^{3}\lambda_{q}^{-2}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
which requires $b^{2}\beta<1$ and $a$ sufficiently large to absorb the constant. Using (3.8f) on $\mathring{R}_{q}$, $\operatorname{supp}\varphi_{\ell}\subset[0,\ell]$ and $c_{R}<c_{*}$ we deduce for $t\in{\mathbb{R}}$
$3c_{*}^{-1}\mathbf{E}\|\mathring{R}_{\ell}(t)\|_{L^{1}}\leqslant\frac{1}{16}\delta_{q+2}e(t).$
For the third term in (4.54), we use (3.3), (3.7), (3.8b) and (3.8c) to obtain
$3\cdot(2\pi)^{3}|\zeta_{\ell}-\zeta_{q}|$ $\lesssim\ell\|e^{\prime}\|_{C^{0}_{[t-1,t]}}+\ell\mathbf{E}\|v_{q}\|_{C^{1}_{[t-1,t],x}}(\|v_{q}\|_{C_{[t-1,t],x}^{0}}+{\|z_{q}\|_{C_{[t-1,t],x}^{0}}})$
$\qquad\qquad+\ell^{1/2-\delta}\mathbf{E}\|{z_{q}}\|_{C_{[t-1,t]}^{1/2-\delta}C_{x}^{0}}(\|v_{q}\|_{C_{[t-1,t],x}^{0}}+\|{z_{q}}\|_{C_{[t-1,t],x}^{0}})$
$\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}(|\!|\!|v_{q}|\!|\!|_{C^{0},1}+|\!|\!|z_{q}|\!|\!|_{C^{0},1})+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{1/2-\delta}_{t}C^{0}_{x},2}(|\!|\!|v_{q}|\!|\!|_{C^{0},2}+|\!|\!|z_{q}|\!|\!|_{C^{0},2})$
$\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}|\!|\!|v_{q}|\!|\!|_{C^{0},1}+\ell\lambda_{q}^{\frac{3}{2}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},1}$
$\qquad\qquad+\ell^{1/2-\delta}|\!|\!|z|\!|\!|_{C^{1/2-\delta}_{t}H^{\frac{3}{2}+\sigma},2}|\!|\!|v_{q}|\!|\!|_{C^{0},2}+\ell^{1/2-\delta}|\!|\!|z|\!|\!|^{2}_{C^{1/2-\delta}_{t}H^{\frac{3}{2}+\sigma},2}$
$\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}\lambda_{q}^{\beta}+\ell^{1/2-\delta}\lambda_{q}^{\beta}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $14b^{2}\beta+2\beta<$ 1 to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-\frac{5}{6}+\beta}<\lambda_{q}^{-2b^{2}\beta}$, and we $a$ was chosen sufficiently large to absorb the constant. For the last term in (4.54), recalling the definition of $\eta_{k,j}$ and $a_{(\xi)}$ in (4.12), we know that for any fixed $t\in[k,k+1]$, the sum with respect to $j,j^{\prime}$ is actually the finite sum of at most two terms. Hence, applying (4.40) with $n=m$ and using the estimates (4.14), (4.15) and (C.7) we obtain
$\mathbf{E}$ $\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{j}\otimes W_{\xi^{\prime}}\circ\Phi_{j^{\prime}}){\mathord{{\rm d}}}x\right|$
$\lesssim\mathbf{E}\sup_{j.j^{\prime}}\sum_{\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}a_{(\xi)}a_{(\xi^{\prime})}B_{(\xi)}\otimes B_{(\xi^{\prime})}e^{i\lambda_{q+1}(\xi\cdot\Phi_{k,j}+\xi^{\prime}\cdot\Phi_{k,j^{\prime}})}{\mathord{{\rm d}}}x\right|$
$\lesssim\sup_{j,j^{\prime}}\sum_{\xi+\xi^{\prime}\neq 0}\frac{\|a_{(\xi)}a_{(\xi^{\prime})}\|_{C^{0}_{t}C^{m}_{x}}+\|a_{(\xi)}a_{(\xi^{\prime})}\|_{C^{0}_{t,x}}(\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C^{m}_{x}}+\|\nabla\Phi_{k,j^{\prime}}\|_{C^{0}_{t}C^{m}_{x}})}{\lambda_{q+1}^{m}}$
$\lesssim\frac{\ell^{-2m-1}\lambda_{q+3}^{(m+2)\gamma}}{\lambda_{q+1}^{m}}+\frac{\ell^{-m}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m}}$
$\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $b^{3}\gamma+2b^{2}\beta<1$ and $5m+5<bm$ to have $\lambda_{q}^{4m+2+(m+2)b^{3}\gamma}<\lambda_{q}^{bm-2b^{2}\beta}$, and we chose $a$ sufficiently large to absorb the constant. This completes the bound for (4.54). Going back to (4.53), it remains to control
$\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(c)}|^{2}+2w_{q+1}^{(p)}\cdot w_{q+1}^{(c)}{\mathord{{\rm d}}}x\right|+\mathbf{E}\left|\int_{\mathbb{T}^{3}}|v_{\ell}-v_{q}+z_{q+1}-z_{q}|^{2}{\mathord{{\rm d}}}x\right|$ (4.55)
$\qquad+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}-v_{q}+z_{q+1}-z_{q})(v_{q}+z_{q}){\mathord{{\rm d}}}x\right|+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}+z_{q+1})\cdot w_{q+1}{\mathord{{\rm d}}}x\right|.$
Using the estimates (4.20) and (4.23), we have
$\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(c)}|^{2}+2w_{q+1}^{(p)}\cdot w_{q+1}^{(c)}{\mathord{{\rm d}}}x\right|\lesssim\mathbf{E}\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}^{2}+2\mathbf{E}\|w_{q+1}^{(p)}\|_{C^{0}_{t,x}}\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}$
$\qquad\qquad\lesssim\frac{\lambda_{q}^{8}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-1}+\frac{\lambda_{q}^{4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-\frac{1}{2}}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $b\geqslant 7$, $14b^{3}\gamma+4b^{2}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{4+3b^{3}\gamma}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Next, we use (3.3), (3.7), (3.8b), (4.6) and (4.33) to obtain
$\mathbf{E}$ $\int_{\mathbb{T}^{3}}|v_{\ell}-v_{q}+z_{q+1}-z_{q}|^{2}{\mathord{{\rm d}}}x+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}-v_{q}+z_{q+1}-z_{q})(v_{q}+z_{q}){\mathord{{\rm d}}}x\right|$
$\lesssim\mathbf{E}\left(\|v_{\ell}-v_{q}+z_{q+1}-z_{q}\|_{C_{t,x}^{0}}^{2}\right)+\mathbf{E}\left(\|v_{\ell}-v_{q}+z_{q+1}-z_{q}\|_{C_{t,x}^{0}}\|z_{q}+v_{q}\|_{C_{t,x}^{0}}\right)$
$\lesssim\mathbf{E}\left(\|v_{\ell}-v_{q}\|^{2}_{C_{t,x}^{0}}\right)+|\!|\!|z_{q+1}-z_{q}|\!|\!|^{2}_{C^{0},2}$
$\qquad+\left(|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2}+|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2}\right)(|\!|\!|z_{q}|\!|\!|_{C^{0},2}+|\!|\!|v_{q}|\!|\!|_{C^{0},2})$
$\lesssim\ell^{2}\lambda_{q}^{3}+(\lambda_{q+1}^{-\frac{\gamma\sigma}{4}}+\lambda_{q+2}^{-2\gamma}\lambda_{q+1}^{\gamma})+(\ell\lambda_{q}^{\frac{3}{2}}+\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{4}})(|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2}+\lambda_{q}^{\beta})$
$\leqslant\lambda_{q}^{-\frac{1}{2}}\lambda_{q+1}^{\beta}+\lambda_{q+1}^{-\frac{\gamma\sigma}{8}+\beta}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\beta+\gamma}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $2b\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta+2b^{2}\beta}<\lambda_{q}^{\frac{\gamma\sigma}{8}b}$, and used $14b^{2}\beta+2b\beta<$ 1, $3b\beta<\gamma$ to have $\lambda_{q}^{b\beta-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-b^{2}\gamma+b\beta+b\gamma}<\lambda_{q}^{-2b^{2}\beta}$, and $a$ was chosen large enough to absorb the constant. For the last term of (4.55), we recall (cf.(4.19)) that $w_{q+1}$ may be written as the curl of a vector field as
$w_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j,\xi}\mathrm{curl}(a_{\xi}B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}}).$
Integrating by parts and using the estimates (3.3), (3.7), (3.8b) and (4.15), we have
$2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}+z_{q+1})\cdot w_{q+1}{\mathord{{\rm d}}}x\right|$ $\lesssim\frac{1}{\lambda_{q+1}}\mathbf{E}\left(\sup_{j}\sum_{\xi}\big{\|}a_{\xi}B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}}\big{\|}_{C^{0}_{t,x}}\|v_{\ell}+z_{q+1}\|_{C_{t,x}^{1}}\right)$
$\lesssim\frac{1}{\lambda_{q+1}}\sum_{\xi}\mathbf{E}(\|a_{\xi}\|_{C^{0}_{t,x}}\cdot\|v_{\ell}+z_{q+1}\|_{C^{0}_{t}C_{x}^{1}})$
$\lesssim\frac{\lambda_{q+3}^{\gamma}}{\lambda_{q+1}}(\|v_{q}\|_{C^{1}_{[t-1,t+1],x}}+|\!|\!|z_{q+1}|\!|\!|_{C^{1},1})$
$\lesssim\frac{\lambda_{q+3}^{\gamma}}{\lambda_{q+1}}(\lambda_{q}^{\frac{3}{2}}+\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},1})\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $b\geqslant 7$ and $12b^{2}\beta+b^{3}\gamma<$ 1 to have $\lambda_{q}^{b^{3}\gamma+\frac{3}{2}}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{b^{3}\gamma+b\gamma}<\lambda_{q}^{b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Summing over the above estimates, we obtain
$\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}\leqslant\frac{1}{4}\delta_{q+2}e(t).$
Hence, the proof of Proposition 3.2 is completed. ∎
5. Stationary solutions to the stochastic Euler equations
We recall that the trajectory space defined in Section 1.1 is $\mathcal{T}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$ and we take $\kappa=\frac{\vartheta}{2}$ in this section. The corresponding shifts $S_{t}$, $t\in{\mathbb{R}}$, on trajectories are given by
$S_{t}(u,B)(\cdot)=(u(\cdot+t),B(\cdot+t)-B(t)),\quad t\in{\mathbb{R}},\quad(u,B)\in{\mathcal{T}}.$
The notion of stationary solution was introduced in Definition 1.4. Our main result of this section is existence of stationary solutions as limits of ergodic averages of solutions constructed in Section 3. This in particular implies their non-uniqueness.
Theorem 5.1.
Let $u$ be a solution obtained in Theorem 3.3 with $e(t)=K$ for some constant $K\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}rL^{2}$ and satisfying (3.12), where $r,L,c_{R}$ are constants as in Theorem 3.3. Then there exists a sequence $T_{n}\to\infty$ and a stationary solution $((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}}),\tilde{u},\tilde{B})$ to (1.1) such that
$\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathcal{L}[S_{t}(u,B)]{\mathord{{\rm d}}}t\to\mathcal{L}[\tilde{u},\tilde{B}]$
weakly in the sense of probability measures on $\mathcal{T}$ as $n\to\infty$. Moreover, it holds true that
[…]
Proof.
First, we claim that there exist $\beta^{\prime},\beta^{\prime\prime}>0$ such that for any $N\in\mathbb{N}$
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
Indeed, reviewing the proof of Theorem 3.3, and by interpolation we deduce for some $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$
$\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}$ $\lesssim\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\vartheta}_{t,x},2r}$
$\lesssim\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{1-\vartheta}_{C^{0}_{t,x},2r}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{\vartheta}_{C^{1}_{t,x},2r}$
$\lesssim\sum_{q\geqslant 0}\delta_{q+1}^{\frac{1-\vartheta}{2}}\lambda_{q+1}^{\frac{3}{2}\vartheta}\delta_{q+1}^{\frac{1}{2}\vartheta}\leqslant\sqrt{3r}La^{\frac{3}{2}b\vartheta}+\lambda_{2}^{\beta}\sum_{q\geqslant 1}\lambda_{q+1}^{\frac{3}{2}\vartheta-\beta}<\infty.$
Hence, we conclude that $v=\lim_{q\rightarrow\infty}v_{q}$ exists and lies in $L^{2r}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$. Similarly, using (3.3), (3.7), (4.33) and interpolation we deduce for the same $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$ and any $p\geqslant 1$
$\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},p}$
$\lesssim$ $\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})}_{C^{0}_{t,x},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{(\frac{1}{4}-\beta^{\prime})\beta^{{}^{\prime\prime}}}_{C^{0}_{t}C^{1}_{x},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{\beta^{\prime}(1-\beta^{{}^{\prime\prime}})}_{C^{\frac{1}{4}}_{t}C^{0}_{x},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{\beta^{\prime}\beta^{{}^{\prime\prime}}}_{C^{\frac{1}{4}}_{t}C^{1}_{x},p}$
$\lesssim$ $\sum_{q\geqslant 0}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})^{(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})}\lambda_{q+2}^{\frac{\gamma}{8}[(\frac{1}{4}-\beta^{\prime})\beta^{\prime\prime}+\beta^{\prime}\beta^{\prime\prime}]}$
$\lesssim$ $\sum_{q\geqslant 0}\lambda_{q+1}^{-\frac{\gamma}{8}{\cdot}\frac{\sigma}{16}}\lambda_{q+2}^{\frac{\gamma}{32}\beta^{\prime\prime}}+\lambda_{q+2}^{-\frac{\gamma}{16}+\frac{\gamma}{32}\beta^{\prime\prime}}\lambda_{q+1}^{\frac{\gamma}{8}}\lesssim\sum_{q\geqslant 0}\lambda_{q+1}^{-\frac{\gamma}{8}{\cdot}\frac{\sigma}{16}+\frac{\gamma}{32}b\beta^{\prime\prime}}+\lambda_{q+1}^{\frac{\gamma}{8}+\frac{\gamma}{32}b\beta^{\prime\prime}-\frac{\gamma}{16}b}<\infty,$
where we used $(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})>\frac{1}{16}$, $b\geqslant 7$ and $\beta^{\prime\prime}<{\mathord{{\rm min}}}{\{\frac{\sigma}{4b},1\}}$ in the last two inequalities. Hence we obtain $\lim_{q\rightarrow\infty}z_{q}=z$ in $L^{p}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$ for any $p\geqslant 1$. Then letting $u=v+z$, we obtain an $(\mathcal{F}_{t})_{t\in\mathbb{R}}$-adapted analytically weak solution $u$ to (1.1) and
$|\!|\!|u|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}<\infty.$
Hence we obtain
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}$ $\leqslant\sup_{s\in{\mathbb{R}}}\sum_{i=-N}^{N-1}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([i,i+1],C^{\beta^{\prime\prime}}_{x})}$
$\leqslant 2N\sup_{t\in\mathbb{R}}\mathbf{E}\|u\|^{2r}_{C^{\beta^{\prime}}([t,t+1],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
Now, we define the ergodic averages of the solution $(u,B)$ as the probability measures on the space of trajectories ${\mathcal{T}}$
$\nu_{T}=\frac{1}{T}\int_{0}^{T}\mathcal{L}[S_{t}(u,B)]{\mathord{{\rm d}}}t,\qquad T\geqslant 0.$
For $R_{N}>0$, $N\in\mathbb{N}$, by Arzelà–Ascoli Theorem and the above assertion, we note that the set
$K_{M}:=\cap_{N=M}^{\infty}\bigg{\{}g;\,$ $\|g\|_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\leqslant R_{N}\bigg{\}}$
is relatively compact in $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$ since $\frac{\vartheta}{2}<\beta^{\prime\prime}$. As a consequence, we deduce that the time shifts $S_{t}u$, $t\in{\mathbb{R}}$, are tight on $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$. Since $S_{t}B$ is a Wiener process for every $t\in{\mathbb{R}}$, the law of $S_{t}B$ does not change with $t\in{\mathbb{R}}$ and is tight. Accordingly, for any $\varepsilon>0$ there is a compact set $\bar{K}_{\varepsilon}$ in ${\mathcal{T}}$ such that
$\sup_{t\in{\mathbb{R}}}{\mathbf{P}}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon})<\varepsilon.$
This implies
$\nu_{T}(\bar{K}_{\varepsilon}^{c})=$ $\frac{1}{T}\int_{0}^{T}\mathbf{P}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon}){\mathord{{\rm d}}}t<\varepsilon$
and therefore there is a weakly converging subsequence of the probability measures $\nu_{T}$, $T\geqslant 0$. That is, there is a subsequence $T_{n}\to\infty$ and $\nu\in{\mathcal{P}}({\mathcal{T}})$ such that $\nu_{T_{n}}\to\nu$ weakly in ${\mathcal{P}}({\mathcal{T}})$. Define the set
$A$ =$=\Big{\{}(u,B)\in{\mathcal{T}};\ \langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r$
$\qquad\qquad=\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle,\quad\forall\psi\in C^{\infty}(\mathbb{T}^{3};{\mathbb{R}}^{3}),\ \mathord{{\rm div}}\psi=0,\ t\geqslant s\Big{\}}.$
Since $(u,B)$ in the statement of the theorem satisfies the equation, we have for all $t\in{\mathbb{R}}$
${\mathcal{L}}[S_{t}(u,B)](A)=1.$
Hence, also $\nu_{T_{n}}(A)=1$ for all $n\in\mathbb{N}$. By Jakubowski–Skorokhod representation theorem, there is a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}})$ with a sequence of random variables $(\tilde{u}^{n},\tilde{B}^{n})$, $n\in\mathbb{N}$, such that ${\mathcal{L}}[\tilde{u}^{n},\tilde{B}^{n}]={\nu_{T_{n}}}$ and $(\tilde{u}^{n},\tilde{B}^{n})$ satisfy equation (1.1) on ${\mathbb{R}}$. By (3.12) we know
$\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2}=\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|{S_{s}u(t)}\|_{L^{2}}^{2}{\mathord{{\rm d}}}s=K.$ (5.2)
By (3.11)
$\sup_{n}\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2r}=\sup_{n}\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|S_{s}u(t)\|_{L^{2}}^{2r}{\mathord{{\rm d}}}s<\infty.$ (5.3)
Moreover, there is a random variable $(\tilde{u},\tilde{B})$ having the law ${\mathcal{L}}[\tilde{u},\tilde{B}]=\nu$ so that
$(\tilde{u}^{n},\tilde{B}^{n})\to(\tilde{u},\tilde{B})\qquad\tilde{\mathbf{P}}\text{-a.s. in }\mathcal{T}.$
Thus, we can pass to the limit in the equation to deduce that $\nu$ is a law of a solution on ${\mathbb{R}}$. The shift invariance follows from the same argument as in [BFH20e, Lemma 5.2]. Namely, it holds for $G\in C_{b}({\mathcal{T}})$ and $r\in{\mathbb{R}}$
$\int_{{\mathcal{T}}}G\circ S_{r}(u,B){\mathord{{\rm d}}}\nu(u,B)=\int_{{\mathcal{T}}}G(u,B){\mathord{{\rm d}}}\nu(u,B).$
Finally, (5.1) follows from (5.2) and (5.3). ∎
Using the above result and choosing different $K$, the second claim in Theorem 1.5 follows.
Appendix 1 A Beltrami waves
In this part we recall the Beltrami waves from [BV19, Section 5.4] whcih is adpated to the convex integration scheme in Proposition 3.2. We point out that the construction is entirely deterministic, that is, none of the functions below depends on $\omega$. Let us begin with the definition of Beltrami waves. Given $\xi\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$, let $A_{\xi}\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$ obey
$A_{\xi}\cdot\xi=0,\quad A_{-\xi}=A_{\xi}.$
We define the complex vector
$B_{\xi}=\frac{1}{\sqrt{2}}(A_{\xi}+i\xi\times A_{\xi}).$
By construction, the vector $B_{\xi}$ has the properties
$|B_{\xi}|=1,\quad B_{\xi}\cdot\xi=0,\quad i\xi\times B_{\xi}=B_{\xi},\quad B_{-\xi}=\overline{B_{\xi}}.$
This implies that for any $\lambda\in\mathbb{Z}$, such that $\lambda\xi\in\mathbb{Z}^{3}$, the function
$W_{(\xi)}:=W_{\xi,\lambda}(x):=B_{\xi}e^{i\lambda\xi\cdot x}$ (A.1)
is $\mathbb{T}^{3}$ periodic, divergence free, and is an eigenfunction of the $\mathrm{curl}$ operator with eigenvalue $\lambda$. That is, $W_{(\xi)}$ is a complex Beltrami plane wave. The following lemma states a useful property for linear combinations of complex Beltrami plane waves, which can be found in [BV19, Proposition 5.5].
Lemma A.1.
Let $\Lambda$ be a given finite subset of $\mathbb{S}^{2}\cap\mathbb{Q}^{3}$ such that $-\Lambda=\Lambda$, and let $\lambda\in\mathbb{Z}$ be such that $\lambda\Lambda\subset\mathbb{Z}$. Then for any choice of coefficients $a_{\xi}\in\mathbb{C}$ with $\overline{a_{\xi}}=a_{-\xi}$ the vector field
$W(x)=\sum_{\xi\in\Lambda}a_{\xi}B_{\xi}e^{i\lambda\xi\cdot x}$ (A.2)
is a real-valued, divergence-free Beltrami vector field $\mathrm{curl}W=\lambda W$, and thus it is a stationary solution of the Euler equations
$\mathord{{\rm div}}(W\otimes W)=\nabla\frac{|W|^{2}}{2}.$
Furthermore, since $B_{\xi}\otimes B_{-\xi}+B_{-\xi}\otimes B_{\xi}=2Re(B_{\xi}\otimes B_{-\xi})=\mathrm{Id}-\xi\otimes\xi$, we have
[…]
The key point of the construction is that the abundance of Beltrami waves allows to find several such flows $v$ with the property that
$\frac{1}{(2\pi)^{3}}\int_{\mathbb{T}^{3}}v\otimes v(t,x){\mathord{{\rm d}}}x$
equals a prescribed symmetric matrix $R$. Indeed we will need to select these flows so as to depend smoothly on the matrix $R$, at least when $R$ belongs to a neighborhood of the identity matrix. In view of (A.3), such selection is made possible by the following lemma.
Lemma A.2.
(Geometric lemma) There exists a sufficiently small $c_{*}>0$ with the following property. Let $B_{c_{*}(\mathrm{Id})}$ denote the closed ball of symmetric $3\times 3$ matrices, centered at $\mathrm{Id}$, of radius $c_{*}$. Then, there exist pairwise disjoint subsets
$\Lambda_{\alpha}\subset\mathbb{S}^{2}\cap\mathbb{Q}^{3}\qquad\alpha\in\{0,1\},$
and smooth positive functions
$\gamma_{\xi}^{(\alpha)}\in C^{\infty}(B_{c_{*}}(\mathrm{Id}))\qquad\alpha\in\{0,1\},\xi\in\Lambda_{\alpha},$
such that the following hold. For every $\xi\in\Lambda_{\alpha}$ we have $-\xi\in\Lambda_{\alpha}$ and $\gamma_{\xi}^{(\alpha)}=\gamma_{-\xi}^{(\alpha)}$. For each $R\in B_{c_{*}}(\mathrm{Id})$ we have the identity
$R=\frac{1}{2}\sum_{\xi\in\Lambda_{\alpha}}\left(\gamma_{\xi}^{(\alpha)}(R)\right)^{2}(\mathrm{Id}-\xi\otimes\xi).$ (A.4)
We label by $n_{*}$ the smallest natural number such that $n_{*}\Lambda_{\alpha}\subset\mathbb{Z}^{3}$ for all $\alpha\in\{0,1\}$.
[…]
$\sum_{\xi\in\Lambda_{\alpha}}\|\gamma^{(\alpha)}_{\xi}\|_{C^{n}(B_{c_{*}}(\mathrm{Id}))}\leqslant M$ (A.5)
holds for $n$ large enough, $\alpha\in\{0,1\}$ and $\xi\in\Lambda_{\alpha}$. This parameter is universal.
Appendix 2 B Estimates of $a_{(\xi)}$ in the proof of Proposition 3.2
For completeness, we include here the detailed proof of the estimates (4.14) and (4.15). We first aim at estimating the $C^{N}_{t,x}$-norm of $\rho=\sqrt{|\mathring{R}_{\ell}|^{2}+\ell^{2}}+c_{*}\zeta_{\ell}$ for $N\in\mathbb{N}$. To this end, we first apply the chain rule [BDLIS15, Proposition C.1] to the function $\Psi(z)=\sqrt{\ell^{2}+z^{2}}$, since $|D^{m}\Psi(z)|\lesssim\ell^{-m+1}$, for $m\geqslant 1$, we obtain for any $N\in\mathbb{N}$
$\|\rho\|_{C^{N}_{t,x}}$ $\lesssim\left\|\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}\right\|_{C^{0}_{t,x}}+\|D\Psi\|_{C^{0}}\|\mathring{R}_{\ell}\|_{C^{N}_{t,x}}+\|D\Psi\|_{C^{N-1}}\|\mathring{R}_{\ell}\|^{N}_{C^{1}_{t,x}}+\|\zeta_{\ell}\|_{C^{N}_{t}}$ (B.1)
$\lesssim\ell+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-N}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2N+1}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{N}+\ell^{-N}\delta_{q+1}\bar{e},$
where we have used definition of $\zeta_{\ell}$ (4.10) in the last inequality. Let us now estimate the $C^{N}_{t,x}$-norm of $a_{(\xi)}$. By Leibniz rule, we get
$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\sum_{m=0}^{N}\|\rho^{\frac{1}{2}}\|_{C^{m}_{t,x}}\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}},$ (B.2)
and estimate each term separately. First, by (4.10)
$\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\lesssim\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}},$
and by Lemma A.2
$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{0}_{t,x}}\lesssim 1.$
Second, applying [BDLIS15, Proposition C.1] to the function $\Psi(z)=z^{1/2},$ $|D^{m}\Psi(z)|\lesssim|z|^{1/2-m},$ for $m=1,\dots,N$, and using (B.1) and $\rho\geqslant\ell$ and $\ell^{-1}\geqslant\bar{e}$ we obtain for $m\geqslant 1$
$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim$ $\|\rho^{1/2}\|_{C^{0}_{t,x}}+\ell^{-1/2}\|\rho\|_{C^{m}_{t,x}}+\ell^{1/2-m}\|\rho\|_{C^{1}_{t,x}}^{m}$ (B.3)
$\lesssim$ $\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}}+\ell^{\frac{1}{2}-2m}\|\mathring{R}_{q}\|^{m}_{C^{0}_{[t-1,t+1],x}}$
$+\ell^{-\frac{1}{2}-m}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}-2m}\delta_{q+1}^{m}+\ell^{-\frac{1}{2}-m}\delta_{q+1}.$
For $m\geqslant 1$, using $\delta_{q+1}\leqslant 3rL^{2}$ the above is bounded by
$\ell^{\frac{1}{2}-2m}(3rL^{2}+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{m}.$
We proceed with a bound for $\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ for $m=0,\dots,N-1$. Keeping [BDLIS15, Proposition C.1] in mind, we need to estimate
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}+\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}+\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}.$ (B.4)
We use $\rho\geqslant\ell$ to have
$\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}},$ (B.5)
and in view of $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$
$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)},$
and by (B.1) and $\delta_{q+1}\leqslant 3rL^{2}$
$\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)},$
hence
$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$ (B.6)
To estimate the first term in (B.4), we write
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\sum_{k=0}^{N-m}\|\mathring{R}_{\ell}\|_{C^{k}_{t,x}}\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}},$ (B.7)
where for $N-m-k=0$ we have
$\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}\lesssim\ell^{-1},$
and for $k=0,\dots,N-m-1$ using (B.1)
$\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}}$ $\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}+\ell^{-2}\|\rho\|_{C^{N-m-k}_{t,x}}+\ell^{-(N-m-k){-1}}\|\rho\|_{C^{1}_{t,x}}^{N-m-k}$
$\lesssim\ell^{-2(N-m-k)-1}+\ell^{-(N-m-k)-2}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m-k)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{N-m-k}.$
Altogether, we therefore bound (B.7) as
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\ell^{-2(N-m)-1}$ $\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$ (B.8)
$+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}.$
Finally, plugging (B.5), (B.6) and (B.8) into (B.4) leads to
$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ $\lesssim\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$
$+\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$
For $N-m\geqslant 1$ the above is bounded by
$\ell^{-2(N-m)-1}(\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+1).$
Combining this with the bounds (B.3) above and plugging into (B.2) yields for $N\geqslant 2$
$\|a_{(\xi)}\|_{C^{N}_{t,x}}$ $\lesssim\ell^{-2N-\frac{1}{2}}(1+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})^{N+1}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (B.9)
where we used inductive assumption (3.8d) for $\mathring{R}_{q}$ in the last inequality. Moreover, for $N=1$, by (B.3) and chain rule, we obtain
$\|a_{(\xi)}\|_{C^{1}_{t,x}}$ $\lesssim\|\rho^{\frac{1}{2}}\|_{C^{1}_{t,x}}+\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\left\|\gamma_{\xi}\left(\mathrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{1}_{t,x}}$ (B.10)
$\lesssim\ell^{-\frac{3}{2}}(\delta_{q+1}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})+(\delta_{q+1}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{\frac{1}{2}})\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{1}_{t,x}}.$
Since $\rho\geqslant\ell$ and $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$, by (B.3) and using chain rule again, we obtain
$\left|\nabla_{t,x}\left(\frac{\mathring{R}_{\ell}}{\rho}\right)\right|\lesssim\frac{|\nabla_{t,x}\mathring{R}_{\ell}|}{\rho}+\frac{|\mathring{R}_{\ell}\cdot\nabla_{t,x}\rho|}{\rho^{2}}\lesssim\ell^{-2}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2}.$ (B.11)
Pluging (B.11) into (B.10), we obtain
$\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{\frac{3}{2}}\leqslant\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (B.12)
And by the definition of $a_{(\xi)}$, it is easy to get
$\|a_{(\xi)}\|_{C^{0}_{t,x}}$ $\lesssim\|\mathring{R}_{q}\|^{\frac{1}{2}}_{C_{[t-1,t+1],x}^{0}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\gamma}.$ (B.13)
Hence, (4.14) and (4.15) follow from the estimates (B.9), (B.12) and (B.13).
Appendix 3 C Estimates for transport equantions
We provide a detailed estimate of the solutions to the transport equations in Section 4.3.1. First, we consider the following transport equation on $[t_{0},T],t_{0}\geqslant 0$:
$(\partial_{t}+v\cdot\nabla)f$ $=g,$ (C.1)
$f(t_{0},x)$ $=f_{0}.$
We recall the following results from [BDLIS15, Proposition D.1, (133), (134)]: {equ} ∥f(t)∥_C_x^1⩽∥f_0∥_C_x^1e^(t-t_0)∥v∥_C^0_[t_0,t]C^1_x+∫_t_0^te^(t-τ)∥v∥_C_[t_0,t]^0C_x^1∥g(τ)∥_C_x^1 dτ, and more generally, for any $N\geqslant 2$, there exists a constant $C=C_{N}$ such that
$\|f(t)\|_{C_{x}^{N}}\leqslant$ $\left(\|f_{0}\|_{C_{x}^{N}}+C(t-t_{0})\|{v}\|_{C^{0}_{[t_{0},t]}C^{N}_{x}}\|f_{0}\|_{C_{x}^{1}}\right)e^{C(t-t_{0})\|{v}\|_{C^{0}_{[t_{0},t]}C^{1}_{x}}}$ (C.2)
$+\int_{t_{0}}^{t}e^{C(t-\tau)\|{v}\|_{C^{0}_{[t_{0},t]}C^{1}_{x}}}\left(\|g(\tau)\|_{C_{x}^{N}}+C(t-\tau)\|{v}\|_{C^{0}_{[t_{0},t]}C^{N}_{x}}\|g(\tau)\|_{C_{x}^{1}}\right){\mathord{{\rm d}}}\tau.$
Consider the following case:
$(\partial_{t}+v\cdot\nabla)\Phi$ $=0,$
$\Phi(t_{0},x)$ $=x.$
We aim at estimating the $C_{x}^{0}$-norm of $\nabla\Phi(t)-\textrm{Id}$. First, let $\Psi(s,x)=\Phi(s,x)-x$, then $\Psi$ statifies the following equation:
$(\partial_{t}+v\cdot\nabla)\Psi=-v,$
$\Psi(t_{0},x)=0.$
We then apply (C) to conclude
$\|\nabla\Phi(t)-\textrm{Id}\|_{C_{x}^{0}}=\|\Psi(t)\|_{C_{x}^{1}}\leqslant\int_{t_{0}}^{t}e^{(t-\tau)\|{v}\|_{C_{[t_{0},t]}^{0}C_{x}^{1}}}\|{v}\|_{C^{0}_{[t_{0},t]}C_{x}^{1}}{\mathord{{\rm d}}}\tau=e^{(t-t_{0})\|{v}\|_{C_{[t_{0},t]}^{0}C_{x}^{1}}}-1.$ (C.3)
Now we use the above result to estimate the solution $\Phi_{k,j}$ of (4.7), $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$. Using (3.8c) and the definition of $z_{q}$ in (3.5), we obtain for any $k\in\mathbb{Z}$
$\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}\leqslant\ell(\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}+\lambda_{q+3}^{\frac{\gamma}{4}})\lesssim\lambda_{q}^{-\frac{1}{2}}\ll 1,$ (C.4)
where we used $b^{3}\gamma<1$ in the second inequality. Since $e^{x}-1\leqslant 2x$ for $x\in[0,1]$ and combining (C.3) with (C.4), we obtain
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\textrm{Id}\|_{C_{x}^{0}}$ $\leqslant e^{2\ell\|v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}-1$ (C.5)
$\lesssim\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}\lesssim\lambda_{q}^{-\frac{1}{2}}\ll 1,$
hence (4.8a) holds ture. Then, (4.8b) follows from (C.5), we have
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\leqslant\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\mathrm{Id}\|_{C_{x}^{0}}+1\leqslant 2,$ (C.6)
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\geqslant 1-\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\mathrm{Id}\|_{C_{x}^{0}}\geqslant\frac{1}{2}.$
Similarly, we use (C.2) and (C.4) again, we obtain for any $n\geqslant 1$
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{n}}$ $\lesssim\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{n+1}}}e^{2\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}}$ (C.7)
$\lesssim\ell^{-(n-1)}\|v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}\lesssim\ell^{-n+\frac{1}{4}},$
where we used $b^{3}\gamma<1$ in the last inequality. Using (3.8a), (4.7), (4.31) and (C.6), we obtain
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\partial_{t}\Phi_{k,j}\|_{C_{x}^{0}}\leqslant\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|(v_{\ell}+z_{\ell})\cdot\nabla\Phi_{k,j}\|_{C_{x}^{0}}\leqslant\lambda_{q+3}^{\frac{\gamma}{2}}.$ (C.8)
Differentiating both sides of (4.7) and using (3.8a), (3.8c), (4.31), (C.6) and (C.7), we can also obtain
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\partial_{t}\nabla\Phi_{k,j}\|_{C_{x}^{0}}\lesssim$ $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla(v_{\ell}+z_{\ell})\cdot\nabla\Phi_{k,j}\|_{C_{x}^{0}}$ (C.9)
$+\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|(v_{\ell}+z_{\ell})\cdot\nabla^{2}\Phi_{k,j}\|_{C_{x}^{0}}$
$\lesssim$ $\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}+\lambda_{q+3}^{\frac{\gamma}{4}}+(\lambda_{q+3}^{\frac{\gamma}{4}}+\lambda_{q+2}^{\gamma})\ell^{-\frac{3}{4}}\leqslant\lambda_{q}^{\frac{3}{2}}\lambda_{q+3}^{\gamma}.$
The estimates (C.7)–(C.9) are used in Section 4.3 and Section 4.6 to estimate the $C^{1}_{t,x}$-norm of $w_{q+1}$ and the $C^{0}_{t,x}$-norm of $\mathring{R}_{q+1}$.

### v3

@@
 Abstract.
-We establish the existence of infinitely many global and stationary solutions in $C(\mathbb{R};C^{\vartheta})$ space for some $\vartheta>0$ to the three dimensional Euler equations driven by an additive noise.
+We establish the existence of infinitely many global and stationary solutions in $C(\mathbb{R},C^{\vartheta})$ space for some $\vartheta>0$ to the three-dimensional Euler equations driven by an additive stochastic forcing.
 The result is based on a new stochastic version of the convex integration method, incorporating the stochastic convex integration method developed in [HZZ22b] and pathwise estimates to derive uniform moment estimates independent of time.
@@
 1. Introduction
-In this paper, we are considered with stochastic Euler equations governing the time evolution of the velocity $u$ of an inviscid fluid on the three dimensional torus $\mathbb{T}^{3}=\mathbb{R}^{3}/(2\pi\mathbb{Z})^{3}$.
+In this paper, we consider the stochastic Euler equations governing the time evolution of the velocity $u$ of an inviscid fluid on the three-dimensional torus $\mathbb{T}^{3}=\mathbb{R}^{3}/(2\pi\mathbb{Z})^{3}$.
 The system reads as
@@
 The right hand side represents a random external force acting on the fluid, which is given by a $GG^{*}$-Wiener process $B$ defined on some probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $GG^{*}$ being trace class operator.
-In early works, stochastic Euler equations have been widely studied in both two and three dimensions, see e.g.[BFM16], [BP01], [CFH19], [GHV14], where local-in-time solutions for (1.1) have been constructed.
-Our goal is to investigate the existence of global and stationary solutions with Hölder continuity to the above system (1.1).
-To achieve this, we employ the method of convex integration, a technique that has gained prominence in fluid dynamics over the past decades.
+The notion of solution to (1.1) we use throughout this paper is that of analytically weak solutions, here recalled.
+Definition 1.1.
+We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is an analytically weak solution to the stochastic Euler system (1.1) provided
+  1. (1)
+    $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ is a stochastic basis with a complete right-continuous filtration;
+  2. (2)
+    $B$ is an $\mathbb{R}^{3}$-valued, mean zero and divergence-free, two-sided $GG^{*}$-Wiener process with respect to the filtration $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$;
+  3. (3)
+    the velocity $u\in C({\mathbb{R}}\times\mathbb{T}^{3})$ $\mathbf{P}$-a.s. and is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted;
+  4. (4)
+    for every $-\infty<s\leqslant t<\infty$ it holds $\mathbf{P}$-a.s.
+    $\langle u(t),\psi\rangle=\langle u(s),\psi\rangle+\int_{s}^{t}\langle u,u\cdot\nabla\psi\rangle{\mathord{{\rm d}}}r+\langle B(t)-B(s),\psi\rangle$
+    for all $\psi\in C^{\infty}(\mathbb{T}^{3},{\mathbb{R}}^{3})$, $\mathord{{\rm div}}\psi=0$.
+Stochastic Euler equations have been extensively studied in both two and three dimensions in previous works, including [BFM16, BP01, CFH19, GHV14], where local-in-time solutions for (1.1) have been constructed.
+Our objective is to investigate the existence of global and stationary solutions in the sense of Definition 1.1 with Hölder continuity to the above system (1.1).
+To achieve this goal, we employ the method of convex integration, a technique that has gained prominence in fluid dynamics over the past decades.
 This method was initially introduced following [DLS10, DLS12, DLS13] by De Lellis and Székelyhidi Jr.
-In these works, the method of convex integration was used in order to prove non-uniqueness of weak solutions with Hölder continuity to Euler equations.
-Subsequently, this method has led to several breakthroughs in the fluid dynamics, including proof of Onsager’s conjecture for the incompressible Euler equations [Ise18, BDLSV19].
+In these works, the convex integration method was used to prove the non-uniqueness of weak solutions with Hölder continuity to Euler equations.
+Subsequently, this method has led to several breakthroughs in fluid dynamics, including proof of Onsager’s conjecture for the incompressible Euler equations [Ise18, BDLSV19].
 We refer to the excellent review article by Buckmaster and Vicol [BV19] for a gentle introduction and further references.
-In recent years, convex integration has already been applied in a stochastic setting, namely, to the isentropic Euler system by Breit, Feireisl and Hofmanová [BFH20] and to the full Euler system by Chiodaroli, Feireisl and Flandoli [CFF19] with linear multiplicative noise.
-Futhermore, ill/well-posedness of dissipative martingale solutions to stochastic 3D Euler equations was studied by Hofmanová, Zhu and Zhu [HZZ22a], along with the existence and non-uniqueness of strong Markov solutions.
-Later, Hölder continuous, global-in-time probabilistically strong solutions to 3D Euler equations perturbed by Stratonovich transport noise was constructed in [HLP22].
-In the mostly recent work [HZZ22b], existence and non-uniqueness of stationary solutions in $L^{2}$ to the stochastic 3D Euler equations was established, using a new stochastic convex integration method.
+In recent years, convex integration has already been applied in a stochastic setting, namely, to the isentropic Euler system by Breit, Feireisl and Hofmanová [BFH20] and to the full Euler system by Chiodaroli, Feireisl and Flandoli [CFF21] with linear multiplicative noise.
+Furthermore, the ill/well-posedness of dissipative martingale solutions to stochastic 3D Euler equations was studied by Hofmanová, Zhu and Zhu [HZZ22a], along with the existence and non-uniqueness of strong Markov solutions.
+Later, Hölder continuous, global-in-time probabilistically strong solutions to 3D Euler equations perturbed by Stratonovich transport noise was constructed in [HLP24].
+In the most recent work [HZZ22b], the existence and non-uniqueness of stationary solutions in $H^{\vartheta}$ to the stochastic 3D Euler equations was established, using a new stochastic convex integration method.
 However, the regularity of the global and stationary solutions to the stochastic 3D Euler equations with additive noise, as constructed in the previous work, cannot achieve Hölder continuity.
@@
 Another motivation for our research is linked to the Onsager’s conjecture.
-It is constructed in [Ise18, BDLSV19] that solutions in $C^{\frac{1}{3}-}$ do not conserve energy to the deterministic Euler equations.
-It is natural to ask whether the regularity of the global, nonunique solutions in the stochastic setting can stay $C^{\frac{1}{3}-}$ a.s..
-To summarize the above discussion, it is natural to investigate the validity of the following claims.
+It is constructed in [Ise18, BDLSV19] that solutions in Hölder space $C^{\frac{1}{3}-}$ to the deterministic Euler equations do not conserve energy.
+It is natural to ask whether the global and nonunique solutions in the stochastic setting can stay $C^{\frac{1}{3}-}$ almost surely.
+To summarize the above discussion, we investigate the validity of the following claims.
   1. (i)
@@
     Existence and non-uniqueness of stationary solutions to the stochastic Euler equations (1.1).
+Our objective is to address the aforementioned issues (i) and (ii) within the context of the stochastic Euler equations on $\mathbb{T}^{3}$ driven by additive stochastic noise.
+We employ a stochastic version of the convex integration method as the primary tool in this study.
+Unlike previous works applying convex integration for the Navier–Stokes and Euler equations in the stochastic setting [HZZ22a, HZZ23a, HZZ23b, HZZ24], our approach is inspired by [CDZ24] and [HZZ22b].
+In these works, the limitations stemming from stopping times, which were previously used to control the noise terms in the iteration, are overcome.
+To be more specific, we have moved away from working with stopping times.
+Instead, we incorporate expectations into the inductive iteration and provide uniform moment estimates.
+However, the construction of Hölder continuous Euler flows in the deterministic setting, as demonstrated in [BV19], requires to introduce transport equations, and pathwise positive lower bounds for the solutions are used to improve regularity.
+As a result, we also introduce some pathwise estimates during iteration for convex integration and use the cutoff technique to control the growth of noise pathwisely.
+Compared to only using stochastic convex integration, the combination of pathwise estimates with it can provide better Hölder regularity of the solutions (see Remark 1.4 below).
+This approach also allows us not to control all the moment bounds during the iteration.
+We hope this cutoff method can be applied to enhance the regularity of the solutions in future work.
 1.1. Main results
-Our aim is to provide some answers to the above problems (i) and (ii) in the physically relevant context of the stochastic Euler equations on $\mathbb{T}^{3}$ driven by additive stochastic noise.
-As the main tool within this study, we apply a stochastic version of the convex integration method.
-Unlike the previous works using convex integration for the Navier–Stokes and Euler equations in the stochastic setting [HZZ19, HZZ22a, HZZ23a, HZZ23b], we are now inspired by [CDZ22] and [HZZ22b] where they overcome the limitation originating from stopping times, previously used to control the noise terms in the iteration.
-More precisely, we no longer work with stopping times and instead we include expectations in the inductive iteration and provide uniform moment estimates.
-However, the construction of Hölder continuous Euler flows in the deteministic setting, as demonstrated in [BV19], requires to introduce transport equations, and pathwise positive lower bounds for the solutions are used to improve regularity.
-As a result, we also introduce some pathwise estimates during iteration for convex integration.
-To this end, we use cutoff method to control the growth of noise pathwisely.
-Compared to only using stochastic convex integration, the combination of pathwise estimates with it can provide better Hölder regularity (see Remark 1.3 below).
-This approach also allows us not to control all the moments bounds during the iteration.
-We hope this cutoff method can be applied to enhance regularity of the solutions in the future work.
-As we are interested in the long time behavior of solutions and particularly in the construction of statistically stationary solutions, we work with entire solutions solving the equations for all times $t\in{\mathbb{R}}$.
-Accordingly, the norms in the convex integration scheme must be chosen appropriately in order to provide the desired global-in-time estimates.
-This is achieved through bounds of the form
-$\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{t\leqslant s\leqslant t+1}\|u(s)\|_{C^{\vartheta}}^{p}\right]<\infty,$
-with $p\geqslant 1$ and small $\vartheta>0$.
-Such bounds provide uniform moment estimates locally in $C(\mathbb{R};C^{\vartheta})$ and guarantee the convergence of the corresponding ergodic averages.
-This leads to the existence of stationary solutions.
-Within this study, we focus on analytically weak solutions which satisfy the equations in the following sense.
-Definition 1.1.
-We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is an analytically weak solution to the Euler system (1.1) provided
-  1. (1)
-    $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ is a stochastic basis with a complete right-continuous filtration;
-  2. (2)
-    $B$ is an $\mathbb{R}^{3}$-valued, spatial mean and divergence free, two-sided $GG^{*}$-Wiener process with respect to the filtration $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$;
-  3. (3)
-    the velocity $u\in C({\mathbb{R}}\times\mathbb{T}^{3})$ $\mathbf{P}$-a.s. and is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted;
-  4. (4)
-    for every $-\infty<s\leqslant t<\infty$ it holds $\mathbf{P}$-a.s.
-    $\langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r\ =\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle$
-    for all $\psi\in C^{\infty}(\mathbb{T}^{3})$, $\mathord{{\rm div}}\psi=0$.
-With this definition at hand, our first main result reads as follows, tackling the problem (i) above.
+Our first main result reads as follows and is proved in Theorem 3.4, tackling the problem (i) above.
 Theorem 1.2.
-Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exists an analytically weak solution to (1.1) in the sense of Definition 1.1 which belongs to $C(\mathbb{R},C^{\vartheta})$, for some $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}\right)$.
+Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exists an analytically weak solution to (1.1) in the sense of Definition 1.1 with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ being the normal filtration generated by the Wiener process $B$, which belongs to $C(\mathbb{R},C^{\vartheta})$ for some $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}\right)$.There are infinitely many such solutions $u$.
 Remark 1.3.
-(i) In Section 3, we prove this result by combining stochastic convex integration method with pathwise estimates to derive uniform moment estimates independent of time, and obtain Hölder continuous solutions, where $\vartheta$ can be chosen less than ${\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}$.
-While we also prove this result in a separable article (see [Lü24]), which applies stochastic convex integration method to the the construction of Hölder continuous Euler flows in [DLS13] to derive uniform moment estimates independent of time.
-The value of $\vartheta$ in this case decreases due to the higher moments required, and we may take $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{73{\cdot}62^{3}},\frac{1}{7{\cdot}62^{3}}\}}$.
-(ii) It is natural to ask whether we could construct the solutions with better regulartiy to (1.1) by combining the gluing method in [BDLSV19] with our cutoff method.
-This will be studied in our future work.
-The second result is the existence and non-uniqueness of stationary solutions in Hölder space.
-In the sequel, we understand stationarity in the sense of shift invariance of laws of solutions on the space of trajectories, see also [BFHM19, BFH20e, FFH21, HZZ22b].
-More precisely, we define the joint trajectory space for the solution and the driving Brownian motion as
-${\mathcal{T}}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$
-for some $\kappa>0$ and let $S_{t}$, $t\in{\mathbb{R}}$, be shifts on trajectories given by
+In [HZZ22b], the authors have constructed solutions to (1.1) with continuous paths in a suitable Sobolev space, namely $C({\mathbb{R}},H^{\vartheta})$ $\mathbf{P}$-almost surely.
+In contrast, we are able to construct solutions achieving the Hölder continuity.
+Additionally, the stochastic convex integration developed in [HZZ22b] requires estimating all the moments of the solutions, while our analysis only necessitates specifying an iteration procedure for a finite number of moments, along with additional pathwise bounds.
+Remark 1.4.
+We can also establish Theorem 1.2 in a separate article (see [Lü24]), which sorely employs the stochastic convex integration method to provide uniform moment bounds.
+In other words, we do not incorporate pathwise estimates (see (3.8a), (3.8c), (3.8d) below) during the iteration for convex integration.
+In that paper, the construction of building blocks of the convex integration method follows from [DLS13], and we do not require pathwise lower bounds of the solutions to the transport equation (4.3) below.
+Therefore, the stochastic convolution $z$ is not cut off.
+However, the value of $\vartheta$ in this case decreases due to the arbitrarily higher moments required to be estimated, and we may take $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{73{\cdot}62^{3}},\frac{1}{7{\cdot}62^{3}}\}}$.
+Our second result establishes the existence and non-uniqueness of stationary solutions in the Hölder space.
+Here, stationarity is interpreted in terms of the shift invariance of laws governing solutions on the space of trajectories, as discussed in [BFHM19, BFH24, FFH21, HZZ22b].
+To be more specific, we define the joint trajectory space for the solution and the driving Wiener process as follows:
+${\mathcal{T}}=C({\mathbb{R}},C^{\kappa})\times C({\mathbb{R}},C^{\kappa})$
+for some $\kappa>0$. And let $S_{t}$, $t\in{\mathbb{R}}$, be shifts on trajectories given by
 $S_{t}(u,B)(\cdot)=(u(\cdot+t),B(\cdot+t)-B(t)),\quad t\in{\mathbb{R}},\quad(u,B)\in{\mathcal{T}}.$
-We note that the shift in the second component acts differently in order to guarantee that for a Brownian motion $B$ the shift $S_{t}B$ is again a Brownian motion.
+We observe that the shift in the second component operates differently to ensure that for a Wiener process $B$, the shift $S_{t}B$ remains a Wiener process.
 Definition 1.5.
@@
 $\mathcal{L}[S_{t}(u,B)]=\mathcal{L}[u,B]\qquad\text{ for all }\quad t\in{\mathbb{R}}.$
-With these definitions at hand, we are able to state our second main result, tackling the problem (ii) above. The result then reads as follows and is proved in Section 5.
+With these definitions at hand, our second main result then reads as follows and is proved in Theorem 6.1, tackling the problem (ii) above.
 Theorem 1.6.
@@
 Organization of the paper
-In Section 2, we collect the basic notations used throughout the paper.
-Section 3 and Section 4 are the core of our proofs: the stochastic convex integration is developed and employed to construct entire non-unique analytically weak solutions.
+Section 2 compiles the fundamental notations used throughout the paper.
+Section 3, 4, 5 constitute the core of our proofs, where stochastic convex integration is developed and employed to construct entire analytically weak solutions.
 In Section 3, we present the main iteration Proposition 3.3 to prove our main Theorem 3.4.
-Section 4 is devoted to the construction of stochastic convex integration with pathwise estimates, namely, proof of iteration Proposition 3.2.
-Then the results in Section 3 are used in Section 5 together with a Krylov–Bogoliubov’s argument to obtain existence of non-unique stationary solutions to the stochastic Euler equations.
-In Appendix A, we recall the construction of Beltrami waves from [BV19] .
-In Appendix B, we give estimates on amplitude functions used in the convex integration construction.
+Section 4 and Section 5 are devoted to the construction of stochastic convex integration with pathwise estimates, namely, proof of iteration Proposition 3.3.
+Then the results from Section 3 are applied in Section 6, coupled with a Krylov–Bogoliubov’s argument, to establish the existence of non-unique stationary solutions to the stochastic Euler equations.
+In Appendix A, we recall the construction and property of Beltrami waves from [BV19] .
+In Appendix B, we estimate amplitude functions used in the convex integration construction.
 Finally, the estimates for transport equations are completed in Appendix C.
-2. Notations
+2. Preliminaries
+In the sequel, we use the notation $a\lesssim b$ if there exists a constant $c>0$ such that $a\leqslant cb$.
 2.1. Function spaces
-Throughout the paper, we employ the notation $a\lesssim b$ if there exists a constant $c>0$ such that $a\leqslant cb$.
-We let ${\mathbb{N}}_{0}:={\mathbb{N}}\cup\{0\}$.
-Given a Banach space $E$ with a norm $\|\cdot\|_{E}$ and $t\in{\mathbb{R}}$, we write $C_{t}E=C([t,t+1];E)$ for the space of continuous functions from $[t,t+1]$ to $E$, equipped with the supremum norm $\|f\|_{C_{t}E}=\sup_{s\in[t,t+1]}\|f(s)\|_{E}$.
-For $\alpha\in(0,1)$ we define $C^{\alpha}_{t}E$ as the space of $\alpha$-Hölder continuous functions from $[t,t+1]$ to $E$, endowed with the norm $\|f\|_{C^{\alpha}_{t}E}=\sup_{s,r\in[t,t+1],s\neq r}\frac{\|f(s)-f(r)\|_{E}}{|r-s|^{\alpha}}+\sup_{s\in[t,t+1]}\|f(s)\|_{E}$.
-For $\alpha\in(0,1)$, we also write $C^{\alpha}$ for functions in the space of $\alpha$-Hölder continuous functions from $\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$ endowed with the norm $\|f\|_{C^{\alpha}}=\sup_{x,y\in\mathbb{T}^{3},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in\mathbb{T}^{3}}|f(x)|.$ We use $L^{p}$ to denote the set of standard $L^{p}$-integrable functions from $\mathbb{T}^{3}$ to $\mathbb{R}^{3}$.
-For $s>0$, $p>1$ we set $W^{s,p}:=\{f\in L^{p};\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}<\infty\}$ with the norm $\|f\|_{W^{s,p}}=\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}$.
-Set $L^{2}_{\sigma}=\{f\in L^{2};\int_{\mathbb{T}^{3}}f\,{\mathord{{\rm d}}}x=0,\mathord{{\rm div}}f=0\}$.
-For $s>0$, we define $H^{s}:=W^{s,2}\cap L^{2}_{\sigma}$.
-For $s<0$ we define $H^{s}$ to be the dual space of $H^{-s}$.
-For $t\in\mathbb{R}$ and a domain $D\subset\mathbb{R}$ we denote by $C^{N}_{t,x}$ and $C^{N}_{D,x}$, respectively, the space of $C^{N}$-functions on $[t,t+1]\times\mathbb{T}^{3}$ and on $D\times\mathbb{T}^{3}$, respectively, $N\in\mathbb{N}_{0}$.
+Given a Banach space $X$ with a norm $\|\cdot\|_{X}$ and $t\in{\mathbb{R}}$, we denote $C_{t}X:=C([t,t+1],X)$ as the space of continuous functions from $[t,t+1]$ to $X$, equipped with the supremum norm
+$\|f\|_{C_{t}X}:=\sup_{s\in[t,t+1]}\|f(s)\|_{X}.$
+For $\alpha\in(0,1)$, we use $C^{\alpha}_{t}X$ to denote the space of $\alpha$-Hölder continuous functions from $[t,t+1]$ to $X$, endowed with the norm
+$\|f\|_{C^{\alpha}_{t}X}:=\sup_{s,r\in[t,t+1],s\neq r}\frac{\|f(s)-f(r)\|_{X}}{|r-s|^{\alpha}}+\|f\|_{C_{t}X}.$
+We also write $C^{0}:=C(\mathbb{T}^{3},{\mathbb{R}}^{3})$ equipped with the supremum norm $\|f\|_{C^{0}}:=\sup_{x\in\mathbb{T}^{3}}|f(x)|$.
+For $\alpha\in(0,1)$, we write $C^{\alpha}:=C^{\alpha}(\mathbb{T}^{3},{\mathbb{R}}^{3})$ as the space of $\alpha$-Hölder continuous functions from $\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$ endowed with the norm
+$\|f\|_{C^{\alpha}}:=\sup_{x,y\in\mathbb{T}^{3},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in\mathbb{T}^{3}}|f(x)|.$
+We denote $L^{p}$ as the set of standard $L^{p}$-integrable functions from $\mathbb{T}^{3}$ to $\mathbb{R}^{3}$.
+For $s>0$, $p>1$ the Sobolev space $W^{s,p}:=\{f\in L^{p};\|f\|_{W^{s,p}}:=\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}<\infty\}$.
+We set $L^{2}_{\sigma}:=\{f\in L^{2};\int_{\mathbb{T}^{3}}f\,{\mathord{{\rm d}}}x=0,\mathord{{\rm div}}f=0\}$.
+For $s>0$, we also denote $H^{s}:=W^{s,2}\cap L^{2}_{\sigma}$.
+For $t\in\mathbb{R}$ and a domain $D\subset\mathbb{R}$ and $N\in\mathbb{N}_{0}:={\mathbb{N}}\cup\{0\}$, we denote the space of $C^{N}$-functions from $[t,t+1]\times\mathbb{T}^{3}$ and $D\times\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$, respectively, by $C^{N}_{t,x}$ and $C^{N}_{D,x}$.
 The spaces are equipped with the norms
 $\|f\|_{C^{N}_{t,x}}=\sum_{\begin{subarray}{c}0\leqslant n+|\alpha|\leqslant N\\ n\in\mathbb{N}_{0},\alpha\in\mathbb{N}^{3}_{0}\end{subarray}}\|\partial_{t}^{n}D^{\alpha}f\|_{L^{\infty}_{[t,t+1]}L^{\infty}},\qquad\|f\|_{C^{N}_{D,x}}=\sum_{\begin{subarray}{c}0\leqslant n+|\alpha|\leqslant N\\ n\in\mathbb{N}_{0},\alpha\in\mathbb{N}^{3}_{0}\end{subarray}}\sup_{t\in D}\|\partial_{t}^{n}D^{\alpha}f\|_{L^{\infty}}.$
-For a Polish space $H$ we denote by $\mathcal{B}(H)$ the $\sigma$-algebra of Borel sets in $H$.
-We also use $\mathring{\otimes}$ to denote the trace-free part of the tensor product.
-We recall the inverse divergence operator $\mathcal{R}$ from [DLS13, Section 4.5], which acts on vector fields $v$ with $\int_{\mathbb{T}^{3}}v{\mathord{{\rm d}}}x=0$ as
+We also use $\mathring{\otimes}$ to denote the trace-free part of the tensor product. For a tensor $T$, we denote its traceless part by $\mathring{T}:=T-\frac{1}{3}\mathrm{Tr}(T)\rm{Id}$.
+We recall [DLS13, Definition 4.2] the inverse divergence operator $\mathcal{R}$ which acts on vector fields $v$ with $\int_{\mathbb{T}^{3}}v{\mathord{{\rm d}}}x=0$ as
 $(\mathcal{R}v)^{kl}=(\partial_{k}\Delta^{-1}v^{l}+\partial_{l}\Delta^{-1}v^{k})-\frac{1}{2}(\delta_{kl}+\partial_{k}\partial_{l}\Delta^{-1})\mathord{{\rm div}}\Delta^{-1}v,$
 for $k,l\in\{1,2,3\}$.
-Then $\mathcal{R}v(x)$ is a symmetric trace-free matrix for each $x\in\mathbb{T}^{3}$, and $\mathcal{R}$ is a right inverse of the div operator, i.e.
+The above inverse divergence operator has the property that $\mathcal{R}v(x)$ is a symmetric trace-free matrix for each $x\in\mathbb{T}^{3}$, and $\mathcal{R}$ is a right inverse of the div operator, i.e.
 $\mathord{{\rm div}}(\mathcal{R}v)=v$.
@@
 2.2. Probabilistic elements
-For a given probability measure $\mathbf{P}$ we denote by $\mathbf{E}$ the expectation under $\mathbf{P}$.
-Regarding the driving noise, we assume that $B$ is $\mathbb{R}^{3}$-valued two-sided $GG^{*}$-Wiener process with zero spatial mean and zero divergence, defined on some probability space $(\Omega,\mathcal{F},\mathbf{P})$ and $G$ is a Hilbert–Schmidt operator from $U$ to $L_{\sigma}^{2}$ for some Hilbert space $U$.
-For $p\in[1,\infty)$, $\delta\in(0,\frac{1}{2})$ and some $\kappa>0$, we denote
-$|\!|\!|u|\!|\!|^{p}_{C^{0},p}:=\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{s\in[t,t+1]}\|u(s)\|^{p}_{C^{0}}\right],\quad|\!|\!|u|\!|\!|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa},p}:=\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|u\|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa}}\right].$
-These norms define function spaces of random variables on $\Omega$ taking values in $C({\mathbb{R}}\times\mathbb{T}^{3})$ and $C^{\frac{1}{2}-\delta}({\mathbb{R}},C^{\kappa})$, respectively, with bounds in $L^{p}(\Omega;{C^{0}(I\times\mathbb{T}^{3}))}$ and $L^{p}(\Omega;{C^{\frac{1}{2}-\delta}(I,C^{\kappa})})$ for bounded interval $I\subset{\mathbb{R}}$.
-Furthermore, the bounds only depend on the length of the interval $I$, not on its location within ${\mathbb{R}}$.
-In the sequel, we simply say that $u$ has a uniform moment of order $p$ locally in $C^{0}({\mathbb{R}}\times\mathbb{T}^{3})$ provided $|\!|\!|u|\!|\!|^{p}_{C^{0},p}<\infty$.
-Similarly, we define the corresponding norms with $C^{0}$ replaced by $C^{\kappa}$, $C^{1}$, $L^{1}$ and $H^{\frac{3}{2}+\kappa}$.
+For a given probability measure $\mathbf{P}$, we use $\mathbf{E}$ to denote the expectation under $\mathbf{P}$.
+Concerning the driving noise, we assume that $B$ is $\mathbb{R}^{3}$-valued two-sided $GG^{*}$-Wiener process with mean zero and divergence-free.
+This process is defined on some probability space $(\Omega,\mathcal{F},\mathbf{P})$, where $G$ is a Hilbert–Schmidt operator from $U$ to $L_{\sigma}^{2}$ for some Hilbert space $U$.
+Given a Banach space $X=C(\mathbb{T}^{3},{\mathbb{R}}^{3})$ or $X=C^{\kappa}(\mathbb{T}^{3},{\mathbb{R}}^{3})$ for some $\kappa>0$. For $p\in[1,\infty)$ and $\delta\in(0,\frac{1}{2})$, we denote
+$|\!|\!|u|\!|\!|^{p}_{X,p}:=\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{s\in[t,t+1]}\|u(s)\|^{p}_{X}\right],\quad|\!|\!|u|\!|\!|^{p}_{C_{t}^{\frac{1}{2}-\delta}X,p}:=\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|u\|^{p}_{C_{t}^{\frac{1}{2}-\delta}X}\right].$
+The above norms denote function spaces of random variables on $\Omega$ taking values in $C({\mathbb{R}},X)$ and $C^{\frac{1}{2}-\delta}({\mathbb{R}},X)$, respectively, with bounds in $L^{p}(\Omega;{C(I,X))}$ and $L^{p}(\Omega;{C^{\frac{1}{2}-\delta}(I,X)})$ for bounded interval $I\subset{\mathbb{R}}$.
+Importantly, the bounds solely depend on the length of the interval $I$ and are independent of its location within ${\mathbb{R}}$.
+In addition, we denote the corresponding norms with $X$ replaced by $L^{1}$ and $H^{\frac{3}{2}+\kappa}$ for some $\kappa>0$.
 3. Stochastic convex integration with pathwise estimates and results
-In [CDZ22] and [HZZ22b], they presented a stochastic convex integration method constructing directly solutions on the whole time line $\mathbb{R}$ by introducing expectations to the iterative estimates in convex integration.
+In [CDZ24] and [HZZ22b], the authors presented a stochastic convex integration method to construct solutions on the entire timeline $\mathbb{R}$.
+They achieved this by introducing expectations to the iterative estimates in convex integration.
 In this section, we integrate their methods with pathwise estimates in the context of the stochastic Euler equations.
-Our goal is to construct a Hölder continuous Euler flow, mirroring the deterministic case as presented in [BV19], where the transport equations (see (4.7) below) are used to improve regularity.
-This method necessitates pathwise positive lower bounds for the solutions of the transport equations.
-As a result, moment bounds during the inductive iteration become insufficient.
-Hence, we introduce additional pathwise estimates in convex integration schemes.
-More precisely, our idea employs the cutoff method to control the growth of the noise, enabling the introduction of pathwise estimates during the inductive iteration.
-This is combined with moment bounds to derive desired uniform estimates.
-Compared to only using moment bounds in the inductive iteration in [CDZ22] and [HZZ22b], the advantage of this approach is that it only requires lower moments of the solutions, and avoids the requirements for higher moments as in [HZZ22b].
+Our objective is to construct Hölder continuous Euler flows, which are similar to the deterministic case as presented in [BV19].
+In that case, the transport equations (see (4.3) below) are used to improve regularity.
+However, this method necessitates pathwise positive lower bounds for the solutions of the transport equations.
+Moment bounds during the inductive iteration become insufficient.
+Therefore, we introduce additional pathwise estimates in convex integration schemes.
+To control the growth of the noise, we employ the cut-off technique to the stochastic convolution.
+This enables the introduction of pathwise estimates during the inductive iteration.
+We combine this with moment bounds to derive desired uniform estimates.
+Compared to only using moment bounds in the inductive iteration in [CDZ24] and [HZZ22b], the advantage of our approach is that it only requires lower moments of the solutions, and avoids the requirements for higher moments as in [HZZ22b].
 This can simplify the computations and enhance the regularity of the solutions.
-At the first step, we decompose stochastic Euler system (1.1) into two parts, one is linear and involves the noise, whereas the second one is a random PDE.
+We intend to develop an iteration procedure leading to the proof of Theorem 1.2.
+To this end, we decompose stochastic Euler system (1.1) into two parts, one is linear and involves the noise, whereas the second one is a random PDE.
 More precisely, we consider the stochastic linear equation
@@
 $\mathord{{\rm div}}z$ $=0,$
-where $B$ is $\mathbb{R}^{3}$-valued two-sided trace-class $GG^{*}$-Wiener process with spatial zero mean and divergence free (see e.g. [LR15, Section 2.1]), and $v$ solves the nonlinear equation
+where $B$ is $\mathbb{R}^{3}$-valued two-sided trace-class $GG^{*}$-Wiener process with divergence-free and mean zero, and $v$ solves the nonlinear equation
 $\partial_{t}v-z+\mathord{{\rm div}}((v+z)\otimes(v+z))+\nabla P$ $=0,$ (3.2)
 $\mathord{{\rm div}}v$ $=0.$
-Here, $z$ is divergence free by the assumptions on the noise $B$ and by $P$ we denote the pressure term associated to $v$.
-We include the linear zero order term $z$ on the left hand side of (3.1) in order to obtain a stabilization of the equation needed for the necessary global in time estimates.
-It will be seen in the course of the construction that the corresponding counter term in (3.2) will not cause any difficulties.
-Using the factorization method, it is standard to derive regularity of the stochastic convolution $z$ on a given stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ canonical filtration given in [LR15, Section 2.1].
-In particular, the following result follows from [DPZ92, Theorem 5.16] together with the Kolmogorov continuity criterion.
+Here, $z$ is divergence-free by the assumptions on the noise $B$ and we denote the pressure term associated with $v$ by $P$
+Furthermore, using the factorization method, it is standard to obtain the regularity of $z$ on a given stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ being normal filtration given in [LR15, Section 2.1].
+Specifically, the following result follows from [DPZ92, Theorem 5.16] together with the Kolmogorov continuity criterion.
 Proposition 3.1.
@@
 Proof.
-We recall that the unique stationary solution to (3.1) has the explicit form $z(t)=\int_{-\infty}^{t}e^{s-t}{\mathord{{\rm d}}}B_{s}$.
-By [LR15, Section 2.1], the Wiener process $B$ can be written as $B=\sum_{k\in{\mathbb{N}}}\sqrt{c_{k}}\beta_{k}e_{k}$ for an orthonormal basis $\{e_{k}\}_{k\in{\mathbb{N}}}$ of $L^{2}_{\sigma}$ consisting of eigenvectors of $GG^{*}$ with corresponding eigenvalues $c_{k}$, and the coefficients satisfy $\sum_{k\in{\mathbb{N}}}c_{k}<\infty$, where $\{\beta_{k}\}_{k\in{\mathbb{N}}}$ are a sequence of mutually independent standard two-sided real-valued Brownian motions.
+We recall that the unique stationary solution to (3.1) has an explicit form $z(t)=\int_{-\infty}^{t}e^{s-t}{\mathord{{\rm d}}}B_{s}$.
+According to [LR15, Section 2.1], the Wiener process $B$ can be written as $B=\sum_{k\in{\mathbb{N}}}\sqrt{c_{k}}\beta_{k}e_{k}$ for an orthonormal basis $\{e_{k}\}_{k\in{\mathbb{N}}}$ of $L^{2}_{\sigma}$ consisting of eigenvectors of $GG^{*}$ with corresponding eigenvalues $c_{k}$, and the coefficients satisfy $\sum_{k\in{\mathbb{N}}}c_{k}<\infty$.
+Here, $\{\beta_{k}\}_{k\in{\mathbb{N}}}$ denotes a sequence of mutually independent standard two-sided real-valued Brownian motions.
 Then it holds for $t\geqslant s$,
@@
 $\mathbf{E}\|z(t)-z(s)\|^{p}_{H^{\frac{3}{2}+\kappa}}\leqslant(p-1)^{\frac{p}{2}}\Big{(}\mathbf{E}\|z(t)-z(s)\|_{H^{\frac{3}{2}+\kappa}}^{2}\Big{)}^{\frac{p}{2}}.$
-The result then follows from Kolmogorov’s continuity criterion and fundamental Sobolev embedding $H^{\frac{3}{2}+\kappa}\hookrightarrow C^{\kappa}$, i.e. for any $\delta\in(0,\frac{1}{2})$
+By Kolmogorov’s continuity criterion and fundamental Sobolev embedding $H^{\frac{3}{2}+\kappa}\hookrightarrow C^{\kappa}$, we obtain for any $\delta\in(0,\frac{1}{2})$
 $\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}{C^{\kappa}}}\right]\lesssim\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}H^{\frac{3}{2}+\kappa}}\right]\leqslant(p-1)^{p/2}L^{p},$
@@
 3.1. Stochastic convex integration set-up and iterative proposition
-Let us now explain how the convex integration iteration is set up.
+Now, we apply the convex integration method to the nonlinear equation (3.2).
+The convex integration iteration is indexed by a parameter $q\in\mathbb{N}_{0}$.
 We consider an increasing sequence $\{\lambda_{q}\}_{q\in\mathbb{N}_{0}}$ which diverges to $\infty$, and a bounded sequence $\{\delta_{q}\}_{q\in\mathbb{N}}$ which decreases to 00.
-We choose $a,b\in\mathbb{N}$ and $\beta\in(0,1)$ and let
+We choose $a,b\in\mathbb{N}$ sufficiently large and $\beta\in(0,1)$ sufficiently small and let
 $\lambda_{q}=a^{b^{q}},\quad\delta_{1}=3rL^{2},\quad\delta_{q}=\frac{1}{2}\lambda_{2}^{2\beta}\lambda_{q}^{-2\beta},\ q\geqslant 2.$
-Here $\beta$ will be chosen sufficiently small, $a$ will be chosen sufficiently large, and more details on the choice of these parameters will be given below in the course of the construction.
-The iteration is indexed by a parameter $q\in\mathbb{N}_{0}$.
+More details on the choice of these parameters will be given in Section 4.1 below.
 At each step $q$, a pair $(v_{q},\mathring{R}_{q})$ is constructed to solve the following system
@@
 $z_{q}(t,x)=\chi_{q}\left(\|\tilde{z}_{q}(t)\|_{C^{1}_{x}}\right)\tilde{z}_{q}(t,x),$ (3.5)
-where $\chi_{q}$ is a non-increasing smooth function such that
-$\chi_{q}(x)=\begin{cases}1,&x\in[0,\lambda_{q+2}^{\gamma}],\\ 0,&x\in(\lambda_{q+3}^{\frac{1}{4}\gamma},\infty),\end{cases}$
-with derivative bounded by $1$, which requires
-$\frac{1}{\lambda_{q+3}^{\gamma/4}-\lambda_{q+2}^{\gamma}}\leqslant 1.$ (3.6)
-And $\tilde{z}_{q}=\mathbb{P}_{\leqslant f(q)}z$ with $f(q)=\lambda_{q+1}^{\gamma/8}$.
+where $\chi_{q}$ is a non-increasing smooth function satisfying
+$\chi_{q}(x)=\begin{cases}1,&x\in[0,\lambda_{q+2}^{\gamma}],\\ 0,&x\in(\lambda_{q+3}^{\frac{1}{4}\gamma},\infty).\end{cases}$
+Moreover, we suppose $\chi_{q}$ with derivative bounded by $1$. This is achieved by choosing $a$ sufficiently large such that
+$\frac{1}{\lambda_{q+3}^{\gamma/4}-\lambda_{q+2}^{\gamma}}\ll 1.$ (3.6)
+In addition, the $\tilde{z}_{q}$ in (3.5) is given by $\tilde{z}_{q}=\mathbb{P}_{\leqslant f(q)}z$ with $f(q)=\lambda_{q+1}^{\gamma/8}$.
 Here $\mathbb{P}_{\leqslant f(q)}$ is the Fourier multiplier operator, which projects a function onto its Fourier frequencies $\leqslant f(q)$ in absolute value, and $\gamma$ is another small parameter which will be fixed in Section 4.1.
@@
 $|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{1}_{x},2p}|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2p}\lesssim\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}H^{\frac{3}{2}+\kappa},2p}^{2}\lesssim(2p-1)L^{2}\lambda_{q+1}^{\frac{\gamma}{8}}.$
-The above estimates (3.7) will be frequently used in the sequel. Under the above assumptions, our main iteration reads as follows, and the proof of this result is presented in Section 4 below.
+The above estimates (3.7) will be frequently used in the sequel.
+Remark 3.2 (On the choice of cut-off).
+In order to apply the stationary phase lemma (see Lemma 5.3 below) to derive the desired bound of $\mathring{R}_{q+1}$, it is necessary to have a $pathwise$ positive lower bound for the solution $\Phi_{k,j}$ of equation (4.3) below (as shown in (4.4b)).
+This requires a pathwise bound of $C^{1}_{x}$-norm of $z_{q}$ part from the noise.
+Consequently, we truncate the $C^{1}_{x}$-norm of $\tilde{z}_{q}$ by a cut-off function depending on $q$.
+Although the pathwise $C^{1}_{x}$-norm for $z_{q}$ may blow up, we choose the parameter $\lambda_{q+3}^{\gamma/4}$ in the definition of $\chi_{q}$ to ensure it does not increase too fast.
+On the other hand, our technique requires absorbing the difference of $z_{q+1}$ and $z_{q}$, i.e.
+$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},r}$ into the small parameter $\delta_{q+3}$ to derive the inductive estimates for $\mathring{R}_{q+1}$ (see (3.8e) and (3.8f) below).
+Therefore, the choice of parameter $\lambda_{q+2}^{\gamma}$ is to bound $|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},r}$ by $\lambda_{q+2}^{-\gamma}$, which can be absorbed into $\delta_{q+3}$ by choosing suitable parameters as stated in Section 4.1 below.
+Under the above assumptions, the key result is the following iterative proposition, and the proof of this result is presented in Section 4 and Section 5 below.
 Proposition 3.3.
 Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$ and let $r>1$ be fixed.
-Given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$, where $L$ is from Proposition 3.2 and $c_{R}>0$ is a small universal constant (see Section 4.3).
+Given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$, where $L$ is from Proposition 3.1 and $c_{R}>0$ is a small universal constant (see (4.8) below).
 There exists a choice of parameters $a,b,\beta$ and $\gamma$ such that the following holds true: Let $(v_{q},\mathring{R}_{q})$ for some $q\in\mathbb{N}_{0}$ be an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted solution to (3.4) satisfying
@@
 $|\!|\!|\mathring{R}_{q}|\!|\!|_{C^{0},r}$ $\leqslant\delta_{q+1},$ (3.8e)
-$|\!|\!|\mathring{R}_{q}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{48}c_{R}\delta_{q+2}\underline{e},$ (3.8f)
+$|\!|\!|\mathring{R}_{q}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{48}c_{R}\delta_{q+2}\underline{e}.$ (3.8f)
 Moreover, for any $t\in{\mathbb{R}}$
-$\frac{3}{4}\delta_{q+1}e(t)\leqslant e(t)-\mathbf{E}\|(v_{q}+z_{q})(t)\|_{L^{2}}^{2}\leqslant\frac{5}{4}\delta_{q+1}e(t),\quad$ $q\geqslant 1,$ (3.9)
-$\mathbf{E}\|v_{0}(t)+z_{0}(t)\|^{2}_{L^{2}}\leqslant\frac{1}{2}e(t),\quad$ $q=0.$
+$\frac{3}{4}\delta_{q+1}e(t)\leqslant e(t)-\mathbf{E}\|(v_{q}+z_{q})(t)\|_{L^{2}}^{2}\leqslant\frac{5}{4}\delta_{q+1}e(t),$ $q\geqslant 1,$ (3.9)
+$\mathbf{E}\|v_{0}(t)+z_{0}(t)\|^{2}_{L^{2}}\leqslant\frac{1}{2}e(t),$ $q=0.$
 Then there exists an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted process $(v_{q+1},\mathring{R}_{q+1})$ which solves (3.4), obeys (3.8a)-(3.8f) and (3.9) at the level $q+1$ and satisfies
 $|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\bar{M}\delta_{q+1}^{\frac{1}{2}},$ (3.10)
-where $\bar{M}$ is a universal constant which will be fixed throughout the iteration.
+where $\bar{M}$ is a universal constant that will be fixed throughout the iteration.
 We intend to start the iteration from $v_{0}\equiv 0$ on ${\mathbb{R}}$.
@@
 $\|\mathring{R}_{0}\|_{C^{0}_{t,x}}\leqslant\|z_{0}\|^{2}_{C^{0}_{t,x}}+\|z_{0}\|_{C^{0}_{t,x}}\leqslant\lambda_{3}^{\frac{1}{2}\gamma}+\lambda_{3}^{\frac{1}{4}\gamma}\leqslant\lambda_{3}^{\gamma},$
-where we choose $a$ sufficiently large to absorb the constant in the last inequality. Taking expectation and using (3.7), we obtain
+where we chose $a$ sufficiently large to absorb the constant in the last inequality. Taking expectation and using (3.7), we obtain
 $|\!|\!|\mathring{R}_{0}|\!|\!|_{C^{0},r}\leqslant|\!|\!|z_{0}|\!|\!|^{2}_{C^{0},2r}+|\!|\!|z_{0}|\!|\!|_{C^{0},r}\leqslant|\!|\!|z|\!|\!|^{2}_{H^{\frac{3}{2}+\sigma},2r}+|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},r}\leqslant 2rL^{2}+rL\leqslant\delta_{1},$
@@
 where we used $6{\cdot}48{\cdot}(2\pi)^{3}c_{R}^{-1}L^{2}\leqslant\underline{e}$, $\delta_{1}=3rL^{2}$ and $\delta_{2}=\frac{1}{2}$.
-Hence, (3.8d), (3.8e) and (3.8f) are satisfied on the level $q=0$.
+Hence, (3.8d), (3.8e) and (3.8f) are satisfied at the level $q=0$.
 Similarly, (3.9) also holds at the level $q=0$, since we have
 $\mathbf{E}\|z_{0}(t)\|^{2}_{L^{2}}\leqslant(2\pi)^{3}\mathbf{E}\|z_{0}(t)\|^{2}_{C^{0}}\leqslant(2\pi)^{3}\mathbf{E}\|z(t)\|^{2}_{H^{\frac{3}{2}+\sigma}}\leqslant(2\pi)^{3}L^{2}\leqslant\frac{1}{2}e(t).$
-We will see in the following context, as long as the $q=0$ step in (3.9) holds ture, $\zeta_{0}(t)$ in (4.11) below is well-defined, thus the iteration can proceed.
+We will see in the following context, as long as the $q=0$ step in (3.9) holds, $\zeta_{0}(t)$ in (4.7) below is well-defined, thus the iteration can proceed.
 3.2. Proof of main result
@@
 $\mathbf{E}\|u(t)\|_{L^{2}}^{2}=e(t).$ (3.12)
-There are infinitely many such solutions by choosing different energies $e$.
+There are infinitely many such solutions by choosing different energy functions $e$.
 Proof.
-We may thus start the iteration Proposition 3.2 with the pair $(v_{0},\mathring{R}_{0})$ and obtain a sequence of solutions $(v_{q},\mathring{R}_{q})$.
+We start the iteration Proposition 3.3 with the pair $(v_{0},\mathring{R}_{0})$ and obtain a sequence of solutions $(v_{q},\mathring{R}_{q})$.
 By (3.8c), (3.10) and interpolation we deduce for any $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}b^{5}},\frac{1}{21{\cdot}b^{4}}\}}\right)\in(0,\frac{2}{3}\beta)$, the following series is summable
@@
 $\lesssim p\sum_{q\geqslant 1}\left(\lambda_{q+1}^{\frac{\gamma}{8}(b\vartheta+\sigma\vartheta-\sigma)}+\lambda_{q+1}^{\gamma(\frac{9}{8}b\vartheta+\frac{1}{8}-b)}\right)<\infty,$
-since $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{b+\sigma},\frac{8b-1}{9b}\}}$.
-Then we obtain $\lim_{q\to\infty}z_{q}=z$ in $L^{p}(\Omega;C({\mathbb{R}},C^{\vartheta}))$ for any $p\geqslant 1$.
+where we used $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{b+\sigma},\frac{8b-1}{9b}\}}$ in the last inequality.
+Then, for any $p\geqslant 1$, we have $\lim_{q\to\infty}z_{q}=z$ in $L^{p}(\Omega;C({\mathbb{R}},C^{\vartheta}))$.
 Furthermore, it follows from (3.8e) that $\lim_{q\to\infty}\mathring{R}_{q}=0$ in $L^{1}(\Omega;C({\mathbb{R}},C^{0}))$.
-Thus $v$ is an analytically weak solution to (3.2).
-Hence letting $u=v+z$ we obtain an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted analytically weak solution to (1.1).
-Moreover, the estimate (3.11) for $u$ holds.
-Finally, (3.12) follows from (3.9), and choosing different energy functions, we can obtain infinitely many such solutions.
+Thus, $v$ is an analytically weak solution to (3.2).
+Letting $u=v+z$ we obtain an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted analytically weak solution to (1.1) and the estimate (3.11) for $u$ holds.
+Finally, (3.12) follows from (3.9).
+By choosing different energy functions, we can obtain infinitely many such solutions.
 This concludes the proof of Theorem 3.4.
 ∎
-4. Proof of Proposition 3.2
-The proof proceeds in several main steps which are similar as in many convex integration schemes.
-First of all, we start the construction by fixing the parameters in Section 4.1 and proceed with a mollification step in Section 4.2.
-Section 4.3 is devoted to introduce the new iteration $v_{q+1}$.
-This is the main part of the construction which differs in general convex integration schemes since we aim to construct global Hölder continuous flows.
-Here, we combine the construction methods of [BV19] and [HZZ22b] to obtain a new perturbation $w_{q+1}$.
-More precisely, we construct new amplitude functions $a_{(\xi)}$ similarly to [HZZ22b], meanwhile, we incorporate the solutions for transport equations into Beltrami waves to obtain an acceptable transport error similar to [BV19].
-As pathwise positive lower bounds of the solutions to the transport equations are required, we use cutoff method to control the growth of noise and add pathwise estimtes of the velocity field $v_{q}$ in inductive iteration.
-Section 4.4 contains the inductive estimates of $v_{q+1}$, especially the pathwise and moment estimates.
+Remark 3.5.
+It appears challenging to specify the energy pathwise, as done in [HZZ23b].
+Indeed, to obtain the pathwise estimates for energy, our technique requires absorbing the pathwise $C^{0}_{x}$-norms of $\mathring{R}_{q}$ and $z_{q+1}-z_{q}$ into the small parameter $\delta_{q+2}$, as demonstrated in the proof of Proposition 5.4 below.
+Instead of using stopping time as in [HZZ23b] to control the noise, we utilize a cut-off function $\chi_{q}$ here, the support of which also increases with $q$.
+This makes it not easy to absorb the above two terms into $\delta_{q+2}$.
+4. Convex integration scheme
+In this section, we aim to construct a pair $(v_{q+1},\mathring{R}_{q+1})$ to solve the system (3.4), given $(v_{q},\mathring{R}_{q})$ as in the statement of Proposition 3.3.
+First of all, we fix the parameters used during the construction in Section 4.1 and continue with a mollification step in Section 4.2.
+Section 4.3 is devoted to introducing the new perturbation $w_{q+1}$.
+We follow the construction method in [HZZ22b] and [BV19] to achieve this.
+More precisely, we construct new amplitude functions $a_{(\xi)}$ similarly to [HZZ22b], meanwhile, we incorporate the solutions for transport equations into Beltrami waves to obtain an acceptable transport error as in [BV19].
+As pathwise positive lower bounds of the solutions to the transport equations are required, we use the cut-off technique to control the growth of noise and add pathwise estimates of the velocity field $v_{q}$ in inductive iteration.
 In Section 4.4, we show how to construct the new stress $\mathring{R}_{q+1}$.
-We establish the inductive pathwise and moment estimates on $\mathring{R}_{q+1}$ in Section 4.6.
-Finally, in Section 4.7 we show how the energy is controlled.
 4.1. Choice of parameters
-In the sequel, additional parameters will be indispensable and their value has to be carefully chosen in order to respect all the compatibility conditions appearing in the estimations below.
+In the sequel, additional parameters will be carefully chosen to verify all the conditions appearing in the estimates below.
 First, for a fixed integer $b\geqslant 7$, we let $\beta,\gamma\in(0,1)$ be small parameters satisfying
@@
 $2b^{2}\beta+\beta<\frac{\gamma\sigma}{8},$
-which can be obtained by choosing $\gamma=\frac{2}{9b^{3}}$ and choosing $\beta$ small such that $\beta<{\mathord{{\rm min}}}{\{\frac{\sigma}{78{\cdot}b^{5}},\frac{2}{27{\cdot}b^{4}}\}}$.
+which can be obtained by choosing $\gamma=\frac{2}{9b^{3}}$ and $\beta<{\mathord{{\rm min}}}{\{\frac{\sigma}{78{\cdot}b^{5}},\frac{2}{27{\cdot}b^{4}}\}}$.
 The last free parameter $a\in 2^{36b^{2}\mathbb{N}}$ is chosen sufficiently large such that $a^{\frac{b\gamma}{8}}\in\mathbb{N}$ and $2<a^{(\frac{b}{4}-1)b^{2}\gamma}$.
-The choice of $a$ ensures $f(q)\in\mathbb{N}$ and (3.6) holds.
-The above requirements for parameters will be used in the following sections.
-In the sequel, we increase $a$ in order to absorb various implicit and universal constants.
+This ensures $f(q)\in\mathbb{N}$ and (3.6) holds.
+In the following sections, we increase $a$ to absorb various implicit and universal constants.
 4.2. Mollification
-We replace $v_{q}$ by a mollified velocity field $v_{\ell}$, we choose a small parameter $\ell\in(0,1)$ as $\ell=\lambda_{q}^{-2}$.
+In order to guarantee smoothness throughout the construction, we replace $v_{q}$ with a mollified velocity field $v_{\ell}$.
+For this purpose, we choose a small parameter $\ell=\lambda_{q}^{-2}$.
 Let $\{\phi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers on $\mathbb{R}^{3}$, and $\{\varphi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers with support in $(0,1)$.
-The one-sided mollifiers here is used in order to preserve adaptedness.
-We define a mollification of $v_{q}$ and $\mathring{R}_{q}$ in space and time by
-$v_{\ell}=(v_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad\mathring{R}_{\ell}=(\mathring{R}_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad z_{\ell}=({z_{q}}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$
+Here we use the one-sided mollifiers to reserve adaptedness.
+We define a mollification of $v_{q}$, $z_{q}$ and $\mathring{R}_{q}$ in space and time by
+$v_{\ell}=(v_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad z_{\ell}=({z_{q}}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad\mathring{R}_{\ell}=(\mathring{R}_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$
 where $\phi_{\ell}=\frac{1}{\ell^{3}}\phi(\frac{\cdot}{\ell})$ and $\varphi_{\ell}=\frac{1}{\ell}\varphi(\frac{\cdot}{\ell})$.
-By definition it is easy to know that $z_{\ell}$ is adapted.
-Then using (3.4) we obtain that $(v_{\ell},\mathring{R}_{\ell})$ satisfies
-$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{com}}),$ (4.1)
+By definition, it is easy to see that $z_{\ell}$, $v_{\ell}$ and $\mathring{R}_{\ell}$ are $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ adapted.
+Then using (3.4) we obtain that $(v_{\ell},\mathring{R}_{\ell})$ obey
+$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{commutator}}),$ (4.1)
 $\mathord{{\rm div}}v_{\ell}$ $=0,$
-where
-$R_{\textrm{com}}=(v_{\ell}+z_{\ell})\mathring{\otimes}(v_{\ell}+z_{\ell})-(((v_{q}+z_{q})\mathring{\otimes}(v_{q}+z_{q}))*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$ (4.2)
+where the commutator error $R_{\textrm{commutator}}$ and pressure term $p_{\ell}$ are given by
+$R_{\textrm{commutator}}=(v_{\ell}+z_{\ell})\mathring{\otimes}(v_{\ell}+z_{\ell})-(((v_{q}+z_{q})\mathring{\otimes}(v_{q}+z_{q}))*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$ (4.2)
 $p_{\ell}=(p_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell}-\frac{1}{3}\big{(}|v_{\ell}+z_{\ell}|^{2}-(|v_{q}+z_{q}|^{2}*_{x}\phi_{\ell})*_{t}\varphi_{\ell}).$
-Using the definition of $z_{q}$ (3.5) and the pathwise inductive assumption (3.8a), we can obtain the following pathwise estimate for $R_{\textrm{com}}$
-$\|R_{\textrm{com}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q}+z_{q}\|^{2}_{C^{0}_{[t-1,t+1],x}}\leqslant 4(\lambda_{q+2}^{2\gamma}+\lambda_{q+3}^{\frac{\gamma}{2}})\leqslant\frac{1}{10}\lambda_{q+4}^{\gamma},$ (4.3)
-where we choose $a$ sufficiently large to absorb the constant in the last inequality.
-Using a mollification estimate and inductive assumption (3.8a) and (3.8c), for any $t\in\mathbb{R}$ and $\delta\in(0,\frac{1}{12})$, we have
-$\|R_{\textrm{com}}(t)\|_{L^{\infty}}$
-$\lesssim\left(\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}+\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C_{x}^{1}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{\frac{1}{2}-\delta}_{[t-1,t]}C_{x}^{0}}\right)\left(\|z_{q}\|_{C^{0}_{[t-1,t],x}}+\|v_{q}\|_{C^{0}_{[t-1,t],x}}\right).$
-Taking expectation and using (3.7), (3.8a) and (3.8b), we obtain
-$|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{C^{0},r}$ $\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell|\!|\!|z_{q}|\!|\!|_{C^{1},2r}+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2r}\right)\left(|\!|\!|z_{q}|\!|\!|_{C^{0},2r}+|\!|\!|v_{q}|\!|\!|_{C^{0},2r}\right)$ (4.4)
-$\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell\lambda_{q+1}^{\frac{\gamma}{8}}2rL+\ell^{\frac{1}{2}-\delta}\lambda_{q+1}^{\frac{\gamma}{8}}2pL^{2}\right)\left(\lambda_{q}^{\beta}+2rL\right)$
-$\lesssim\lambda_{q}^{-\frac{1}{2}+\beta}+\lambda_{q}^{-\frac{3}{4}+\beta}\lambda_{q+1}^{\frac{\gamma}{8}}$
-$\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
-where we used $b\gamma+4b^{3}\beta<1$ to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{-\frac{3}{4}+\beta+\frac{1}{8}b\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $a$ was chosen sufficiently large to absorb the constant.
-By (4.4), we also obtain
-$\qquad|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$ (4.5)
-Similarly, by inductive assumption (3.8c) and the same requirements for parameters as above, we also obtain for any $t\in\mathbb{R}$
-$\|(v_{q}-v_{\ell})(t)\|_{L^{\infty}}\lesssim\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}\leqslant\ell\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (4.6)
 4.3. Perturbation of the velocity
-As usual in convex integration schemes, in order to guarantee smoothness throughout the construction, the new velocity field $v_{q+1}$ will be constructed as a perturbation of $v_{\ell}$:
-$v_{q+1}=:v_{\ell}+w_{q+1},$
-where $w_{q+1}:=w_{q+1}^{(p)}+w_{q+1}^{(c)}$ is constructed following [BV19]. We proceed construction with introducing transport equations and time cutoffs.
+As usual in convex integration schemes, the new velocity field $v_{q+1}$ is a perturbation of $v_{\ell}$:
+$v_{q+1}:=v_{\ell}+w_{q+1},$
+where $w_{q+1}:=w_{q+1}^{(p)}+w_{q+1}^{(c)}$ is constructed following [BV19]. We proceed with construction by introducing flow maps that satisfy transport equations and time cutoffs.
 4.3.1. Flow maps and cutoffs
@@
 Similar to [BV19], a natural way to achieve this, is to replace the linear phase $\xi\cdot x$ in the definition of the Beltrami wave $W_{\xi,\lambda}$ by the nonlinear phase $\xi\cdot\Phi(x,t)$, where $\Phi$ is transported by the aforementioned vector field.
-A slight difference from [BV19] is that we extend the definition of $\Phi$ to the whole line w.r.t time variable.
+A slight difference from [BV19] is that we extend the definition of $\Phi$ to the whole time line ${\mathbb{R}}$.
 For any integer $k\in\mathbb{Z}$, we subdivide $[k,k+1]$ into time intervals of size $\ell$ and solve transport equations on these intervals.
-For $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$ we define the adapted map $\Phi_{k,j}:\Omega\times{{\mathbb{R}}^{3}}\times[k+(j-1)\ell,k+(j+1)\ell]\to{{\mathbb{R}}^{3}}$ as the ${\mathbb{T}}^{3}$ periodic solution of
+For $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$, we define the adapted map $\Phi_{k,j}:\Omega\times{{\mathbb{R}}^{3}}\times[k+(j-1)\ell,k+(j+1)\ell]\to{{\mathbb{R}}^{3}}$ as the ${\mathbb{T}}^{3}$ periodic solution of
 $(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\Phi_{k,j}$ $=0,$ (4.3)
 $\Phi_{k,j}(k+(j-1)\ell,x)$ $=x.$
-We put further details on the $C_{x}^{n}$-estimates of $\Phi_{k,j}$ in Appendix C. Specifically, the following two estimates which follows from (C.5) and (C.6)
+We put further details on the $C_{x}^{n}$-estimates of $\Phi_{k,j}$ in Appendix C. Specifically, the following two estimates which follow from (C.5) and (C.6)
 $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\textrm{Id}\|_{C_{x}^{0}}\leqslant\lambda^{-\frac{1}{2}}_{q}\ll 1,$ (4.4a)
 $\frac{1}{2}\leqslant$ $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\leqslant 2.$ (4.4b)
-As indicated by the proof in Appendix C, the validity of (4.8a) requires that both $v_{q}$ and $z_{q}$ have pointwise bounds, and this is why we need to cut off $\tilde{z}_{q}$ and introduce (3.8c) in the inductive iteration, instead of solely considering the moment bounds in the inductive iteration as in [HZZ22b].
-The above estimates play a crucial role in the cancellation of the estimates for the new Reynolds stress $\mathring{R}_{q+1}$.
+As indicated by the proof in Appendix C, the validity of (4.4a) requires that both $v_{q}$ and $z_{q}$ have pathwise bounds.
+Therefore, we cut off $\tilde{z}_{q}$ and introduce (3.8c) in the inductive iteration, instead of solely considering the moment bounds in the inductive iteration as in [HZZ22b].
+The above estimates play a crucial role in canceling the oscillation error in the new Reynolds stress $\mathring{R}_{q+1}$.
 In particular, (4.4b) ensures the validity of stationary phase Lemma 5.3 below in Section 5.2, which requires the positive lower bounds of $\nabla\Phi_{k,j}$.
-We could easily extend $\Phi_{k,j}$ from $\Omega\times\mathbb{R}^{3}\times[k+(j-1)\ell,k+(j+1)\ell]$ to $\Omega\times\mathbb{R}^{3}\times[k,k+1]$ in an adapted way, which is still denoted as $\Phi_{k,j}$.
-The extension is not unique but in the following we only use the value of $\Phi_{k,j}$ on $[k+(j-1)\ell,k+(j+1)\ell]$.
-We also let $\eta$ be a non-negative function supported in $(-1,1)$, which is 1 on $(-\frac{1}{4},\frac{1}{4})$ and such that the square of the shifted bump functions
+Moreover, we could easily extend $\Phi_{k,j}$ from $\Omega\times\mathbb{R}^{3}\times[k+(j-1)\ell,k+(j+1)\ell]$ to $\Omega\times\mathbb{R}^{3}\times[k,k+1]$ in an adapted way, which is still denoted as $\Phi_{k,j}$.
+The extension is not unique but in the following, we only use the value of $\Phi_{k,j}$ on $[k+(j-1)\ell,k+(j+1)\ell]$.
+We also let $\eta$ be a non-negative function supported in $(-1,1)$, which equals to 1 on $(-\frac{1}{4},\frac{1}{4})$ and such that the square of the shifted bump functions
 $\eta_{j}(t)=\eta(\ell^{-1}t-j)$
-form a partition of unity
-$\sum_{j}\eta^{2}_{j}(t)=1,$
+form a partition of unity, namely, for all $t\in[0,1]$
+$\sum_{j}\eta^{2}_{j}(t)=1.$
 […]
@@
 $\zeta_{q}(t):=\frac{1}{3{\cdot}(2\pi)^{3}}\Big{[}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q}(t)+{z_{q}}(t)\|_{L^{2}}^{2}\Big{]},$ (4.7)
-and
-$\zeta_{\ell}:=\zeta_{q}*_{t}\varphi_{\ell},$
-where $c_{*}>0$ is a constant introduced in Lemma A.2 and we take the constant $c_{R}$ in (3.8f) satisfying $c_{R}<c_{*}$. Now, we define the amplitude functions
+and $\zeta_{\ell}:=\zeta_{q}*_{t}\varphi_{\ell}$. The constant $c_{*}>0$ is introduced in Lemma A.2 and we take the constant $c_{R}$ in Proposition 3.3 satisfying
+$c_{R}<c_{*}.$ (4.8)
+Now, we define the amplitude functions
 $a_{(\xi)}(t,x)=a_{q+1,j,\xi}(t,x)=c_{*}^{-\frac{1}{2}}\rho^{\frac{1}{2}}\eta_{k,j}(t)\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right),\ t\in[k,k+1],$ (4.9)
 where $\gamma_{\xi}^{(j)}$ is introduced in Appendix A. By the definition of $\rho$ we know
-$\left\|\textrm{Id}-\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}(t,x)}{\rho}\right)\right\|_{C^{0}}\leqslant c_{*}.$
+$\left\|\textrm{Id}-\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{0}_{t,x}}\leqslant c_{*}.$
 […]
@@
 […]
-$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (4.14)
-and
-$\|a_{(\xi)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma},\qquad\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (4.15)
-With these preparations in hand, we define the principal part $w_{q+1}^{(p)}$ of the perturbation $w_{q+1}$ as in [BV19], more precisely, for $t\in[k,k+1]$, we define
+With these preparations in hand, we define the principal part $w_{q+1}^{(p)}$ of the perturbation $w_{q+1}$ as in [BV19]. More precisely, for $t\in[k,k+1]$, we define
 $w_{(\xi)}(t,x):=a_{q+1,j,\xi}(t,x)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)},$
@@
 $=\sum_{j}\sum_{\xi\in\Lambda_{j}}c_{*}^{-\frac{1}{2}}\rho^{\frac{1}{2}}\eta_{k,j}(t)\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)},$
-where $B_{\xi}$ is suitable vector defined in Appendix A.
-Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(p)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted.
+where $B_{\xi}$ is a suitable vector defined in Appendix A.
+Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}$ are $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted, we deduce that $w_{q+1}^{(p)}$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted.
 Next, we define the incompressibility correction $w_{q+1}^{(c)}$.
-We aim to add a corrector to $w_{q+1}^{(p)}$ such that the resulting function is perfect $\mathrm{curl}$, making it thus divergence free.
+We aim to add a corrector to $w_{q+1}^{(p)}$ such that the resulting function is perfect $\mathrm{curl}$, making it thus divergence-free.
 To this end, it is useful to introduce the following scalar $phase$ function
@@
 Using the same notation as in Appendix A, we define
-$W_{(\xi)}=B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}.$
-Since from Appendix A $\mathrm{curl}$ $W_{(\xi)}=\lambda_{q+1}W_{(\xi)}$, and $\phi_{(\xi)}$ and $a_{(\xi)}$ are scalar functions, we have
-$a_{(\xi)}\phi_{(\xi)}W_{(\xi)}=\frac{1}{\lambda_{q+1}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right)-\frac{1}{\lambda_{q+1}}\nabla\left(a_{(\xi)}\phi_{(\xi)}\right)\times W_{(\xi)}.$
+$W_{(\xi)}(x)=B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}.$
+Since $\phi_{(\xi)}$ and $a_{(\xi)}$ are scalar functions, it follows from Appendix A that $\mathrm{curl}$ $W_{(\xi)}=\lambda_{q+1}W_{(\xi)}$. By a direct computation, we deduce
+$a_{(\xi)}\phi_{(\xi)}W_{(\xi)}=\frac{1}{\lambda_{q+1}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right)-\frac{1}{\lambda_{q+1}}\nabla\left(a_{(\xi)}\phi_{(\xi)}\right)\times W_{(\xi)}.$ (4.13)
 We therefore define
 $w_{(\xi)}^{(c)}(t,x):$ $=\frac{1}{\lambda_{q+1}}\nabla(a_{(\xi)}\phi_{(\xi)})\times B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}$ (4.14)
-$=\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}(t,x)-\textrm{Id})\xi\right)\times W_{(\xi)}(\Phi_{k,j}(t,x)),\ t\in[k.k+1].$
+$=\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}(t,x)-\textrm{Id})\xi\right)\times W_{(\xi)}(\Phi_{k,j}(t,x)),\ t\in[k,k+1].$
 The incompressibility correction $w^{(c)}_{q+1}$ is defined as
-$w^{(c)}_{q+1}(t,x)=\sum_{j}\sum_{\xi\in\Lambda_{j}}w_{(\xi)}^{(c)}(t,x)$
-for all $t\in[k,k+1]$, $k\in\mathbb{Z}$.
-Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(c)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted, and the new perturbation is defined as
-$w_{q+1}=w^{(p)}_{q+1}+w^{(c)}_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j}\sum_{\xi\in\Lambda_{j}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right),$ (4.19)
-and so clearly $w_{q+1}$ is mean zero, divergence free and $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. We define the velocity field at level $q+1$ as
+$w^{(c)}_{q+1}(t,x)=\sum_{j}\sum_{\xi\in\Lambda_{j}}w_{(\xi)}^{(c)}(t,x).$
+Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}$ are $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted, we deduce that $w_{q+1}^{(c)}$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted.
+Moreover, by (4.13) we deduce that
+$w^{(p)}_{q+1}+w^{(c)}_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j}\sum_{\xi\in\Lambda_{j}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right).$
+Finally, the new perturbation is defined as
+$w_{q+1}:=w^{(p)}_{q+1}+w^{(c)}_{q+1},$ (4.15)
+and so clearly $w_{q+1}$ is mean zero, divergence-free and $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted. The new velocity at the level $q+1$ is defined as
 $v_{q+1}:=v_{\ell}+w_{q+1}.$
-Thus, it is also $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted.
+Thus, by the previous discussion, it is also divergence-free and $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted.
+4.4. Definition of the Reynolds stress $\mathring{R}_{q+1}$
+Recalling the system (4.1)
+$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{commutator}}),$
+[…]
+$\mathord{{\rm div}}\mathring{R}_{q+1}-\nabla p_{q+1}$ (4.16)
+$=\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}}_{\mathord{{\rm div}}(R_{\textrm{transport}})}+\underbrace{\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}}$
+$+\underbrace{w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{Nash}})}+\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}+\mathord{{\rm div}}(w_{q+1}^{(c)}\otimes w_{q+1}+w_{q+1}^{(p)}\otimes w_{q+1}^{(c)})}_{\mathord{{\rm div}}(R_{\textrm{corrector}})+\nabla p_{\textrm{corrector}}}$
+$+\underbrace{\mathord{{\rm div}}(v_{q+1}\otimes(z_{q+1}-z_{\ell})+(z_{q+1}-z_{\ell})\otimes v_{q+1}+z_{q+1}\otimes z_{q+1}-z_{\ell}\otimes z_{\ell})-z_{q+1}+z_{\ell}}_{\mathord{{\rm div}}(R_{\textrm{commutator1}})+\nabla p_{\textrm{commutator1}}}$
+$+\mathord{{\rm div}}(R_{\textrm{commutator}})-\nabla p_{\ell}.$
+[…]
+$R_{\textrm{transport}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}\right),$
+$R_{\textrm{Nash}}$ $:=\mathcal{R}(w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})),$
+$R_{\textrm{corrector}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right)+\left(w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}\right),$
+$R_{\textrm{commutator1}}$ $:=v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})+(z_{q+1}-z_{\ell})\mathring{\otimes}v_{q+1}+z_{q+1}\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}-\mathcal{R}(z_{q+1}-z_{\ell}),$
+$\nabla p_{\textrm{corrector}}$ $:=\frac{1}{3}(2w_{q+1}^{(c)}\cdot w_{q+1}^{(p)}+|w_{q+1}^{(c)}|^{2}),$
+$\nabla p_{\textrm{commutator1}}$ $:=\frac{1}{3}(v_{q+1}\cdot z_{q+1}-v_{q+1}\cdot z_{\ell}+z_{q+1}\cdot v_{q+1}-z_{\ell}\cdot v_{q+1}+z_{q+1}\cdot z_{q+1}-z_{\ell}\cdot z_{\ell}).$
+In order to define the remaining oscillation error from the second line in (4.16), we first note that for $j$ and $j^{\prime}$ such that $|j-j^{\prime}|\geqslant 2$, we have $\eta_{k,j}(t)\eta_{k,j^{\prime}}(t)=0$ Second, we have $\Lambda_{j}\cap\Lambda_{j^{\prime}}=\emptyset$ for $|j-j^{\prime}|=1$, and by Lemma A.1 we have $\mathord{{\rm div}}(W_{\xi}\otimes W_{\xi^{\prime}}+W_{\xi^{\prime}}\otimes W_{\xi})=\nabla(W_{\xi}\cdot W_{\xi^{\prime}})$.
+Therefore, we may use Lemma A.2 and (4.10) to obtain
+$\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})$ $=\mathord{{\rm div}}\left(\sum_{j,j^{{}^{\prime}},\xi,\xi^{{}^{\prime}}}w_{(\xi)}\otimes w_{(\xi^{{}^{\prime}})}+\mathring{R}_{\ell}\right)$ (4.17)
+$=\mathord{{\rm div}}\left(\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell}\right)^{2}B_{\xi}\otimes B_{-\xi}+\mathring{R}_{\ell}\right)$
+$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
+$=\mathord{{\rm div}}\left(c^{-1}_{*}\rho(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell})+\mathring{R}_{\ell}\right)$
+$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
+=$\nabla\left(c_{*}^{-1}\rho\right)+\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})})\nabla\left(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\right)$
+$\qquad\qquad\qquad+\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\nabla\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)$
+$=\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}.$
+Note that compared to [BV19], a slight difference is the appearance of $\mathord{{\rm div}}(\rho\mathrm{Id})$.
+This term can be put into the pressure term $p_{\textrm{oscillation}}$.
+Above, we have denoted
+$p_{\textrm{oscillation}}:=\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)+c_{*}^{-1}\rho,$
+and
+$R_{\textrm{oscillation}}:=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}\left(\left(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})}-\frac{W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}}{2}\mathrm{Id}\right)\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})})\right).$
+Finally we define the Reynolds stress at the level $q+1$ by
+$\mathring{R}_{q+1}=R_{\textrm{transport}}+R_{\textrm{oscillation}}+R_{\textrm{Nash}}+R_{\textrm{corrector}}+R_{\textrm{commutator}}+R_{\textrm{commutator1}},$ (4.18)
+and
+$p_{q+1}=p_{\ell}-p_{\textrm{oscillation}}-p_{\textrm{corrector}}-p_{\textrm{commutator1}}.$
+5. Inductive estimates and proof of Proposition 3.3
+In this section, we collect all the necessary estimates of $v_{q+1}$, $\mathring{R}_{q+1}$ and the energy to complete the proof of the Proposition 3.3.
+As a preliminary step, we need to give estimates on the $C_{t,x}^{N}$-norm of the amplitude functions $a_{(\xi)}$ defined in Section 4.3.
+This is done in forthcoming Proposition 5.1, whose proof is given in Appendix B.
+Proposition 5.1.
+Let $a_{(\xi)}$ be defined by (4.9). Then we have for any $N\geqslant 2$
+$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (5.1)
+and
+$\|a_{(\xi)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma},\qquad\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}\lambda_{q+3}^{2\gamma},$ (5.2)
+where the implicit constants are independent of $q$.
+This section is organized as follows: Section 5.1 contains the inductive estimates of $v_{q+1}$, especially the pathwise and moment estimates.
+We establish the inductive pathwise and moment estimates on $\mathring{R}_{q+1}$ in Section 5.2.
+Finally, in Section 5.3, we show how the energy is controlled.
+Note that we first give pathwise estimates by using (5.1) and (5.2), then taking expectation and using inductive assumptions can derive uniform moment estimates independent of time.
+In addition, we also need the estimates of transport equations, as provided in the appendix C.
+A useful stationary phase lemma on estimating $\mathring{R}_{q+1}$ previously used in the deterministic setting is recalled in Section 5.2.
 5.1. Inductive estimates for $v_{q+1}$
-In this section, we verify the inductive estimates (3.8a), (3.8b), (3.8c) and (3.10). By the definition of $w^{(p)}_{q+1}$ (4.16), it follows from (3.8d), (4.10) and (A.5) that
+In this section, we verify the inductive estimates (3.8a), (3.8b), (3.8c) and (3.10). By the definition of $w^{(p)}_{q+1}$ in (4.11), it follows from (3.8d), (4.6) and (A.5) that
 $\|w^{(p)}_{q+1}\|_{C^{0}_{t,x}}\lesssim 2|\Lambda_{j}|c_{*}^{-\frac{1}{2}}M(\|\mathring{R}_{q}\|^{\frac{1}{2}}_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}})+2|\Lambda_{j}|M\bar{e}\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\frac{3}{4}\gamma},$ (5.3)
-where $M$ is the universal constant in (A.5). Taking expectation on the both side of (4.20), we then use (3.8e) to obtain
+where $M$ is the universal constant in (A.5). Taking expectation on both sides of (5.3), we use (3.8e) to obtain
 $|\!|\!|w^{(p)}_{q+1}|\!|\!|_{C^{0},2r}\lesssim 2|\Lambda_{j}|c_{*}^{-\frac{1}{2}}M(|\!|\!|\mathring{R}_{q}|\!|\!|_{C^{0},r}^{\frac{1}{2}}+\ell^{\frac{1}{2}})+2|\Lambda_{j}|M\bar{e}\delta_{q+1}^{\frac{1}{2}}\leqslant\frac{1}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$ (5.4)
 where $\bar{M}$ is a universal constant satisfying $9c_{*}^{-\frac{1}{2}}M|\Lambda_{j}|+8M\bar{e}|\Lambda_{j}|<\bar{M}$ and we choose $a$ sufficiently large to absorb the constant in the above two inequalities.
-Applying (4.8b), (4.15) and (C.8), we also obtain
+Using (4.4b), (5.2) and (C.8), we obtain
 $\|w^{(p)}_{q+1}\|_{C^{1}_{t,x}}$ $\lesssim\sup_{j}\sum_{\xi\in\Lambda_{j}}\|a_{(\xi)}\|_{C^{1}_{t,x}}+\lambda_{q+1}\|a_{(\xi)}\|_{C^{0}_{t,x}}\left(\|\partial_{t}\Phi_{k,j}\|_{C^{0}}+\|\nabla\Phi_{k,j}\|_{C^{0}}\right)$ (5.5)
 $\lesssim\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}+\lambda_{q+3}^{2\gamma}\lambda_{q+1}\leqslant\frac{1}{4}\lambda_{q+1}^{\frac{3}{2}}\delta_{q+1}^{\frac{1}{2}},$
-where we used $b\geqslant 7$ and $12b^{3}\gamma+b\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma+4}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}b-b\beta{\tiny}}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant.
-We proceed to estimate $w^{(c)}_{q+1}$, using a standard mollification estimate, (4.8a) and (4.15) implies
+where we used $b\geqslant 7$ and $12b^{3}\gamma+b\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma+4}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}b-b\beta{\tiny}}$ in the last inequality.
+We also chose $a$ sufficiently large to absorb the constant.
+We next estimate $w^{(c)}_{q+1}$. Using a standard mollification estimate, (4.4a) and (5.2) we obtain
 $\|w^{(c)}_{q+1}\|_{C^{0}_{t,x}}$ $\lesssim\sup_{j}\sum_{\xi\in\Lambda_{j}}\left(\frac{\|\nabla a_{(\xi)}\|_{C^{0}_{t,x}}}{\lambda_{q+1}}+\|a_{(\xi)}\|_{C^{0}_{t,x}}\cdot\|\nabla\Phi_{k,j}-\textrm{Id}\|_{C^{0}}\right)$ (5.6)
@@
 where we used $b\geqslant 7$ and $12b\gamma^{3}+b\beta<$ 1 to have $\lambda_{q}^{b^{3}\gamma}<\lambda_{q}^{\frac{b}{2}}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{9+3b^{3}\gamma}<\lambda_{q}^{\frac{5}{2}b-b\beta}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant.
-By (4.21) and (4.24), we have
-$|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\frac{3}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$
-[…]
-$|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$
-Combining the above two estimates, we obtain
+We use a standard mollification estimate and inductive assumption (3.8c) to obtain for any $t\in\mathbb{R}$
+$\|(v_{q}-v_{\ell})(t)\|_{L^{\infty}}\lesssim\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}\leqslant\ell\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}},$ (5.9)
+where we used $b\beta<1$ and chose $a$ sufficiently large to absorb the constant. Taking expectation we obtain
+$|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (5.10)
+Combining (5.4), (5.7) and (5.10) we obtain
 $|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{0},2r}\leqslant|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}+|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\bar{M}\delta_{q+1}^{\frac{1}{2}}.$
@@
 Therefore, we have verified that (3.8a), (3.8b) and (3.8c) hold at the level $q+1$.
-4.5. Definition of the Reynolds stress $\mathring{R}_{q+1}$
-[…]
-$\mathord{{\rm div}}\mathring{R}_{q+1}-\nabla p_{q+1}$ (4.28)
-$=\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}}_{\mathord{{\rm div}}(R_{\textrm{transport}})}+\underbrace{\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}}$
-$+\underbrace{w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{Nash}})}+\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}+\mathord{{\rm div}}(w_{q+1}^{(c)}\otimes w_{q+1}+w_{q+1}^{(p)}\otimes w_{q+1}^{(c)})}_{\mathord{{\rm div}}(R_{\textrm{corrector}})+\nabla p_{\textrm{corrector}}}$
-$+\underbrace{\mathord{{\rm div}}(v_{q+1}\otimes z_{q+1}-v_{q+1}\otimes z_{\ell}+z_{q+1}\otimes v_{q+1}-z_{\ell}\otimes v_{q+1}+z_{q+1}\otimes z_{q+1}-z_{\ell}\otimes z_{\ell})-z_{q+1}+z_{\ell}}_{\mathord{{\rm div}}(R_{\textrm{commutator1}})+\nabla p_{\textrm{commutator1}}}$
-$+\mathord{{\rm div}}(R_{\textrm{commutator}})-\nabla p_{\ell}.$
-[…]
-$R_{\textrm{transport}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}\right),$
-$R_{\textrm{Nash}}$ $:=\mathcal{R}(w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})),$
-$R_{\textrm{corrector}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right)+\left(w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}\right),$
-$R_{\textrm{commutator1}}$ $:=v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})+(z_{q+1}-z_{\ell})\mathring{\otimes}v_{q+1}+z_{q+1}\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}-\mathcal{R}(z_{q+1}-z_{\ell}),$
-$\nabla p_{\textrm{corrector}}$ $:=\frac{1}{3}(2w_{q+1}^{(c)}\cdot w_{q+1}^{(p)}+|w_{q+1}^{(c)}|^{2}),$
-$\nabla p_{\textrm{commutator1}}$ $:=\frac{1}{3}(v_{q+1}\cdot z_{q+1}-v_{q+1}\cdot z_{\ell}+z_{q+1}\cdot v_{q+1}-z_{\ell}\cdot v_{q+1}+z_{q+1}\cdot z_{q+1}-z_{\ell}\cdot z_{\ell}).$
-In order to define the remaining oscillation error from the second line in (4.28), first note that for $j,j^{\prime}$ such that $|j-j^{\prime}|\geqslant 2$, we have $\eta_{k,j}(t)\eta_{k,j^{\prime}}(t)=0$ Second, we have $\Lambda_{j}\cap\Lambda_{j^{\prime}}=\emptyset$ for $|j-j^{\prime}|=1$, and by Lemma A.1 we have $\mathord{{\rm div}}(W_{\xi}\otimes W_{\xi^{\prime}}+W_{\xi^{\prime}}\otimes W_{\xi})=\nabla(W_{\xi}\cdot W_{\xi^{\prime}})$.
-Therefore, we may use (4.13) and Lemma A.2 to obtain
-$\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})$ $=\mathord{{\rm div}}\left(\sum_{j,j^{{}^{\prime}},\xi,\xi^{{}^{\prime}}}w_{(\xi)}\otimes w_{(\xi^{{}^{\prime}})}+\mathring{R}_{\ell}\right)$ (4.29)
-$=\mathord{{\rm div}}\left(\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell}\right)^{2}B_{\xi}\otimes B_{-\xi}+\mathring{R}_{\ell}\right)$
-$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
-$=\mathord{{\rm div}}\left(c^{-1}_{*}\rho(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell})+\mathring{R}_{\ell}\right)$
-$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
-=$\nabla\left(c_{*}^{-1}\rho\right)+\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})})\nabla\left(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\right)$
-$\qquad\qquad\qquad+\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\nabla\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)$
-$=\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}.$
-Note that compared to [BV19], a slight difference in this term is the appearance of $\mathord{{\rm div}}(\rho\mathrm{Id})$, which can be put into the pressure term $p_{\textrm{oscillation}}$.
-Thus, we have denoted
-$p_{\textrm{oscillation}}:=\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)+c_{*}^{-1}\rho,$
-and
-$R_{\textrm{oscillation}}:=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}\left(\left(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})}-\frac{W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}}{2}\mathrm{Id}\right)\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})})\right).$
-Finally we define the Reynolds stress at the level $q+1$ by
-$\mathring{R}_{q+1}=R_{\textrm{transport}}+R_{\textrm{oscillation}}+R_{\textrm{Nash}}+R_{\textrm{corrector}}+R_{\textrm{commutator}}+R_{\textrm{commutator1}},$ (4.30)
-and
-$p_{q+1}=p_{\ell}-p_{\textrm{oscillation}}-p_{\textrm{corrector}}-p_{\textrm{commutator1}}.$
 5.2. Inductive estimates for $\mathring{R}_{q+1}$
-In this section, we shall verify (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$.
+In this section, we shall verify (3.8d), (3.8e) and (3.8f) hold at the level $q+1$.
 We first proceed with estimates of $R_{\mathrm{commutator1}}$ in subsection 5.2.1.
-Then we recall a useful sationary phase lemma in subsection 4.6.2.
-Finally, the remianing errors are controlled in subsection 4.6.34.6.6
+Then we recall a stationary phase lemma in subsection 5.2.2.
+Finally, the remaining errors are controlled in subsection 5.2.35.2.6.
 5.2.1. Estimate on $R_{\mathrm{commutator1}}$
-First, by the definition of $z_{q}$ we have
+First, by the definition of $z_{q}$ in (3.5), we have
 $\|z_{q}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+3}^{\frac{\gamma}{4}}.$ (5.13)
@@
 $\|\mathcal{R}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}$ $\leqslant\|z_{q+1}-z_{\ell}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\frac{\gamma}{4}}+\lambda_{q+3}^{\frac{\gamma}{4}}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
-where we chose $a$ sufficiently large to absorb the constant in the above. Summing over the estimates above, the following pathwise bound holds true for any $t\in\mathbb{R}$
+where we chose $a$ sufficiently large to absorb the constant in the above. Summing over the estimates above, the following pathwise bound holds for any $t\in\mathbb{R}$
 $\|R_{\textrm{commutator1}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}+\|z_{q+1}$$\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}\|_{C^{0}_{t,x}}$ (5.14)
 $+\|\mathcal{R}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}\leqslant\frac{4}{5}\lambda_{q+4}^{\gamma}.$
-Next we estimate the $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of $R_{\textrm{com1}}$. Before proceeding, we give an estimate of $z_{q+1}-z_{q}$.
+Next we estimate the $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of $R_{\textrm{commutator1}}$, and we first give an estimate of $z_{q+1}-z_{q}$.
 Lemma 5.2.
@@
 Proof.
-By the definition of $z_{q}$, we deduce for any $p\geqslant 1$
+By the definition of $z_{q}$ in (3.5), we deduce for any $p\geqslant 1$
 $|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},p}$ $=\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|z_{q+1}(s)-z_{q}(s)\|_{C_{x}^{0}}^{p}\right)^{\frac{1}{p}}$
@@
 […]
+where we used $9b\gamma\leqslant 8b^{2}\gamma+\gamma$ to have $\lambda_{q+3}^{-\gamma}\lambda_{q+2}^{\gamma/8}\leqslant\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\gamma/8}$.
 Finally, combining (5.17) with (5.18) implies
 […]
-In particular, taking $p=2r$ impies (4.33). Hence, the proof of Lemma 4.1 is completed. ∎
-With Lemma 4.1 in hand, we proceed to estimate $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$, using estimates (4.27), (4.34) and Hölder’s inequality, we have
+In particular, taking $p=2r$ impies (5.15). Hence, the proof of Lemma 5.2 is complete. ∎
+Next, we estimate $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$. Using (5.12), (5.16) and Hölder’s inequality, we have
 $|\!|\!|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}\leqslant|\!|\!|v_{q+1}|\!|\!|_{C^{0},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\beta}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}).$
@@
 $\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2r}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}.$
-Using a standard mollification estimate, we obtain for any $\delta\in(0,\frac{1}{12})$
+We use a standard mollification estimate to have for any $\delta\in(0,\frac{1}{12})$
 $\|z_{\ell}(t)-z_{q}(t)\|_{L^{\infty}}\lesssim\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C^{1}_{x}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{1/2-\delta}_{[t-1,t]}C^{0}_{x}}.$
-Taking expectation on the both side and using (3.7), we deduce
+Taking expectation on both sides and using (3.7), we deduce
 $|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}$ $\lesssim\ell\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z_{q}\|^{2r}_{C_{[t-1,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2r}}+\ell^{\frac{1}{2}-\delta}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z_{q}\|^{2r}_{C_{[t-1,t+1]}^{1/2-\delta}C_{x}^{0}}\right)^{\frac{1}{2r}}$ (5.19)
 $\lesssim\ell^{\frac{5}{12}}\lambda_{q+1}^{\frac{\gamma}{8}}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z\|^{2r}_{C_{[t-1,t+1]}^{1/2-\delta}H^{\frac{3}{2}+\sigma}}\right)^{\frac{1}{2r}}\lesssim\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}\sqrt{2r}L.$
-Hence, using (4.27), (4.37) and Hölder’s inequality, we deduce
+Hence, using (5.12), (5.19) and Hölder’s inequality, we have
 $|\!|\!|v_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}\leqslant|\!|\!|v_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\beta+\gamma}\lambda_{q}^{-\frac{3}{4}}.$
-Similarly, using (3.7) and (4.37) again, we deduce
+Similarly, we use (3.7) and (5.19) again to deduce
 $|\!|\!|z_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|z_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}$
 $\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}.$
-Summarizing over the estimates above, we obtain
+Summarizing the estimates above, we obtain
 $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|v_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},r}$ (5.20)
@@
 and
-$|\!|\!|\mathring{R}_{\textrm{commutator1}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e},$
+$|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e},$
 where we used $2b^{2}\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta-\frac{b}{8}\gamma\sigma}<\lambda_{q}^{-2b^{3}\beta}$.
-We also used $3b\beta<\gamma$, $14b^{3}\beta+2b\gamma<$ 1 to have $\lambda_{q}^{b\beta+b\gamma-b^{2}\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{b\gamma+b\beta-\frac{3}{4}}<\lambda_{q}^{-2b^{3}\beta}$ in the above two inequalities and we chose $a$ sufficiently large to absorb the constant.
+We also used $3b\beta<\gamma$, $14b^{3}\beta+2b\gamma<$ 1 to have $\lambda_{q}^{b\beta+b\gamma-b^{2}\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{b\gamma+b\beta-\frac{3}{4}}<\lambda_{q}^{-2b^{3}\beta}$ in the above two inequalities and chose $a$ sufficiently large to absorb the constant.
 5.2.2. Stationary phase Lemma
-Before we estimate the remaining terms in (4.30) separately to show that the stress $\mathring{R}_{q+1}$ obeys the bound (3.8d), (3.8e) and (3.8f) at level $q+1$, we first introduce the following lemma which makes rigorous the fact that $\mathcal{R}$ obeys the same elliptic regularity estimates as $|\nabla|^{-1}$.
+Before we estimate the remaining terms in (4.18), we first introduce the following lemma which makes rigorous the fact that $\mathcal{R}$ obeys the same elliptic regularity estimates as $|\nabla|^{-1}$.
 We recall the following stationary phase lemma (see for example [BV19, Lemma 5.7] and [DS17, Lemma 2.2]), adapted to our setting.
@@
 where the implicit constant depends on $C,\alpha$ and m (in particular, not on the frequency $\lambda$).
-The above lemma is used when estimating the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of the remaining terms in (4.30).
+The above lemma is used to estimate the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of the remaining terms in (4.18).
 Indeed, for a fixed $t\in[k+(j-1)\ell,k+(j+1)\ell]$, in view of the bound (4.4b), we have $\frac{1}{2}\leqslant\|\Phi_{k,j}(\cdot,t)\|_{C^{0}_{x}}\leqslant 2$ on $\mathbb{T}^{3}$.
@@
 The implicit constant depends only on $\alpha$ and $n$.
-In the following subsection, we aim to use the above discussion to control the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$ norms, respectively, of the transport error, oscillation error, Nash error and corrector error.
+In the following subsections, we aim to use the above discussion to control the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$ norms, respectively, of the transport error, oscillation error, Nash error and corrector error.
 We first fix an integer $b\geqslant 7$, then choose $\beta$, $\alpha$ sufficiently small such that $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 and choose $m\in\mathbb{N}$ such that $m>\frac{2b+5}{b-5}$, at last we choose $a$ sufficiently large to absorb the constant.
 5.2.3. Estimate on the transport error.
-Inspecting the definition of $w_{q+1}^{(p)}$ as (4.16), we notice that the material derivative $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$ cannot lead to higher frequency term, since the term $W_{(\xi)}\circ\Phi_{k,j}$ is perfectly transported by $v_{\ell}+z_{\ell}$.
+Inspecting the definition of $w_{q+1}^{(p)}$ as (4.11), we note that the material derivative $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$ cannot lead to higher frequency term, since the term $W_{(\xi)}\circ\Phi_{k,j}$ is perfectly transported by $v_{\ell}+z_{\ell}$.
 Therefore, for $t\in[k,k+1]$, $k\in\mathbb{Z}$ we have
 $(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}=\sum_{j,\xi}(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}W_{(\xi)}\circ\Phi_{k,j}.$
-Returning to the $C^{N}_{t,x}$-norm estimate of $a_{(\xi)}$ in (4.14) and (4.15) and the definition of $z_{q}$ (3.5), we use chain rule and (3.8a) to obtain that for any $n\in\mathbb{N}_{0}$
+Recalling the $C^{N}_{t,x}$-norm estimate of $a_{(\xi)}$ in (5.1) and (5.2) and the definition of $z_{q}$ in (3.5), we use chain rule and (3.8a) to obtain that for any $n\in\mathbb{N}_{0}$
 $\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}\|_{C^{n}_{t,x}}$ $\lesssim\|a_{(\xi)}\|_{C^{n+1}_{t,x}}+\sum_{k=0}^{n}\|v_{\ell}+z_{\ell}\|_{C^{k}_{t,x}}\cdot\|\nabla a_{(\xi)}\|_{C^{n-k}_{t,x}}$ (5.24)
@@
 $\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
-where we used $b\geqslant 7$ and $13b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant.
+where we used $b\geqslant 7$ and $13b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$.
+We also used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant.
 Then taking expectation we obtain
-$|\!|\!|R_{\textrm{transport}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
-and
-$|\!|\!|R_{\textrm{transport}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
+$|\!|\!|R_{\textrm{transport}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{transport}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
 5.2.4. Estimate on the oscillation error.
-Based on previous computations (4.29) in Section 4.5, we can rewrite oscillation error as follows, for $t\in[k,k+1]$, $k\in\mathbb{Z}$
+Based on the previous computations in (4.17), we can rewrite oscillation error as follows, for $t\in[k,k+1]$, $k\in\mathbb{Z}$
 $R_{\textrm{oscillation}}=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}$ $\bigg{[}\left(W_{(\xi)}\circ\Phi_{k,j}\otimes W_{(\xi^{{}^{\prime}})}\circ\Phi_{k,j^{{}^{\prime}}}-\frac{1}{2}(W_{(\xi)}\circ\Phi_{k,j}\cdot W_{(\xi^{{}^{\prime}})}\circ\Phi_{k,j^{{}^{\prime}}})\textrm{Id}\right)$
 ×$\times\left(\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\right)\bigg{]}.$
-Then using (4.14), (4.15) and chain rule, we have that for $n\in\mathbb{N}_{0}$
+Then using (5.1), (5.2) and chain rule, we have for any $n\in\mathbb{N}_{0}$
 $\|\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})\|_{C^{n}_{t,x}}$ $\lesssim\sum_{k=0}^{n+1}\|a_{(\xi)}\|_{C^{k}_{t,x}}\|a_{(\xi)}\|_{C^{n-k+1}_{t,x}}$ (5.25)
@@
 $\lesssim$ $\lambda_{q+1}\sum_{l=0}^{n}\ell^{-2l-1}\lambda_{q+3}^{(l+2)\gamma}\ell^{-(n-l)}\lesssim\lambda_{q+1}\ell^{-2n-1}\lambda_{q+3}^{(n+2)\gamma},$
-where we used $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) in the second inequality, namely $\|\nabla\Phi_{k,j}(t)\|_{C^{n}_{x}}\leqslant\ell^{-n}$ for $n\geqslant 1$.
-Combining (4.43), (4.44) with (4.45) and using $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) again, applying (4.41) with $n=m$ and $a$ replaced by $\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})$ implies
+where we used $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) in the second inequality, namely, $\|\nabla\Phi_{k,j}(t)\|_{C^{n}_{x}}\leqslant\ell^{-n}$ for $n\geqslant 1$.
+Combining (5.25), (5.26), (5.27) with (C.7) again, applying (5.23) with $n=m$ and $a$ replaced by $\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})$, we obtain
 $\|R_{\textrm{oscillation}}\|_{C^{0}_{t,x}}\lesssim$ $\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
@@
 $\leqslant$ $\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
-where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\alpha-2b^{3}\beta}$, and used $5m+2b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-2)b}$, and we chose $a$ sufficiently large to absorb the constant.
+where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\alpha-2b^{3}\beta}$.
+We also used $5m+2b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-2)b}$ and we chose $a$ sufficiently large to absorb the constant.
 Then taking expectation we obtain
-$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
-and
-$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
+$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
 5.2.5. Estimate on the Nash error.
@@
 $w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell})=\sum_{j,\xi}a_{(\xi)}W_{(\xi)}\circ\Phi_{k,j}\cdot\nabla(v_{\ell}+z_{\ell}).$
-Similar to (4.42), using (3.8a), (4.14), (4.15), (4.31) and chain rule, we have for $n\in\mathbb{N}_{0}$
+Using (3.8a), (5.1), (5.2), (5.13) and chain rule, we have for $n\in\mathbb{N}_{0}$
 $\|a_{(\xi)}\nabla(v_{\ell}+z_{\ell})\|_{C^{n}_{t,x}}\lesssim\ell^{-2n-2}\lambda_{q+3}^{(n+2)\gamma}.$ (5.29)
@@
 $\lesssim\left(\frac{\ell^{-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}\right)\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
-where we used $b\geqslant 7$ and $13b^{3}\gamma+3b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{4+2b^{3}\gamma+b\alpha}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+3b^{3}\gamma+b\alpha}<\lambda_{q}^{2b-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant.
+where we used $b\geqslant 7$ and $13b^{3}\gamma+3b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{4+2b^{3}\gamma+b\alpha}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+3b^{3}\gamma+b\alpha}<\lambda_{q}^{2b-2b^{3}\beta}$.
+We also used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$ and chose $a$ sufficiently large to absorb the constant.
 Then taking expectation, we obtain
-$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
-and
-$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
+$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{Nash}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
 5.2.6. Estimate on the corrector error.
-The corrector error has two pieces, the transport derivative of $w^{(c)}_{q+1}$ by the flow of $v_{\ell}+z_{\ell}$, and the residual contribution from the nonlinear term, which are easier to estimate due to (4.20) and (4.23).
-Thus we obtain
+The corrector error has two pieces, one is the transport derivative of $w^{(c)}_{q+1}$ by the flow of $v_{\ell}+z_{\ell}$.
+The orther is a residual contribution from the nonlinear terms, which is easier to estimate due to (5.3) and (5.6).
+Therefore, we have
 $\|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\leqslant\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}^{2}\leqslant\frac{\lambda_{q}^{8}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-1}\leqslant\frac{1}{3{\cdot}6{\cdot}96}c_{R}\delta_{q+3},$
@@
 $\|w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma}\left(\frac{\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{\gamma}\lambda_{q}^{-\frac{1}{2}}\right)\leqslant\frac{1}{3{\cdot}6{\cdot}96}c_{R}\delta_{q+3},$
-where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-2b^{3}\beta}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{3}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
-Combining the above estimates and taking expectation, we obtain
+where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-2b^{3}\beta}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{3}\beta}$.
+Also, $a$ was chosen sufficiently large to absorb the constant.
+Combining the above estimates and taking expectations we obtain
 $|\!|\!|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}|\!|\!|_{C^{0},r}$ $\leqslant\frac{1}{12}\delta_{q+2},$ (5.32)
 $|\!|\!|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{6{\cdot}96}c_{R}\delta_{q+3}\underline{e}.$
-The remaining term in corrector error is
+The remaining term in the corrector error is
 $\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right).$
@@
 $\leqslant$ $\frac{1}{6{\cdot}96}c_{R}\delta_{q+3},$
-where we used $b\geqslant 7$ and $14b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{10+4b^{3}\gamma}<\lambda_{q}^{2b-\alpha b-2b^{3}\beta}$ and $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-\alpha b-2b^{3}\beta}$, and used $5m+5+b<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant.
+where we used $b\geqslant 7$ and $14b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{10+4b^{3}\gamma}<\lambda_{q}^{2b-\alpha b-2b^{3}\beta}$ and $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-\alpha b-2b^{3}\beta}$.
+We also used $5m+5+b<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$ and chose $a$ sufficiently large to absorb the constant.
 Taking expetation and combining (5.32), we obatin
-$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
-and
-$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6{\cdot}48}c_{R}\delta_{q+3}\underline{e}.$
-Summarizing over all the estimates we have established above and by (4.3), (4.4) and (4.32), we obtain
+$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{corrector}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6{\cdot}48}c_{R}\delta_{q+3}\underline{e}.$
+5.2.7. Estimate on the commutator error
+Using the definition of $z_{q}$ in (3.5) and the pathwise inductive assumption (3.8a), we obtain
+$\|R_{\textrm{commutator}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q}+z_{q}\|^{2}_{C^{0}_{[t-1,t+1],x}}\leqslant 4(\lambda_{q+2}^{2\gamma}+\lambda_{q+3}^{\frac{\gamma}{2}})\leqslant\frac{1}{10}\lambda_{q+4}^{\gamma},$ (5.33)
+where we chose $a$ sufficiently large to absorb the constant in the last inequality.
+Using a mollification estimate and inductive assumption (3.8a) and (3.8c), we have for any $t\in\mathbb{R}$ and $\delta\in(0,\frac{1}{12})$
+$\|R_{\textrm{commutator}}(t)\|_{L^{\infty}}$
+$\lesssim\left(\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}+\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C_{x}^{1}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{\frac{1}{2}-\delta}_{[t-1,t]}C_{x}^{0}}\right)\left(\|z_{q}\|_{C^{0}_{[t-1,t],x}}+\|v_{q}\|_{C^{0}_{[t-1,t],x}}\right).$
+Taking expectation and using (3.7), (3.8a) and (3.8b), we obtain
+$|\!|\!|R_{\textrm{commutator}}|\!|\!|_{C^{0},r}$ $\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell|\!|\!|z_{q}|\!|\!|_{C^{1},2r}+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2r}\right)\left(|\!|\!|z_{q}|\!|\!|_{C^{0},2r}+|\!|\!|v_{q}|\!|\!|_{C^{0},2r}\right)$ (5.34)
+$\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell\lambda_{q+1}^{\frac{\gamma}{8}}2rL+\ell^{\frac{1}{2}-\delta}\lambda_{q+1}^{\frac{\gamma}{8}}2pL^{2}\right)\left(\lambda_{q}^{\beta}+2rL\right)$
+$\lesssim\lambda_{q}^{-\frac{1}{2}+\beta}+\lambda_{q}^{-\frac{3}{4}+\beta}\lambda_{q+1}^{\frac{\gamma}{8}}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
+where we used $b\gamma+4b^{3}\beta<1$ to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{-\frac{3}{4}+\beta+\frac{1}{8}b\gamma}<\lambda_{q}^{-2b^{3}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
+By (5.34), we also obtain
+$\qquad|\!|\!|R_{\textrm{commutator}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$ (5.35)
+Summarizing over all the estimates we have established above and by (5.14) and (5.33), we obtain
 $\|\mathring{R}_{q+1}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\gamma},\quad|\!|\!|\mathring{R}_{q+1}|\!|\!|_{C^{0},r}\leqslant\delta_{q+2},\quad|\!|\!|\mathring{R}_{q+1}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e}.$ (5.36)
-Hence, (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$.
+Hence, (3.8d), (3.8e) and (3.8f) hold at the level $q+1$.
 5.3. Inductive estimates for the energy
-To conclude the proof of Proposition 3.2, we shall verify (3.9) holds at level $q+1$. To this end, we define
-$\delta E(t):=\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}.$
+To conclude the proof of Proposition 3.3, we shall verify (3.9) holds at the level $q+1$.
 Proposition 5.4.
 It holds for $t\in{\mathbb{R}}$
-$\delta E(t)\leqslant$ $\frac{1}{4}\delta_{q+2}e(t).$ (4.52)
+$\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}\leqslant$ $\frac{1}{4}\delta_{q+2}e(t).$ (5.37)
 Proof.
-By the definition of $\zeta_{q}$ in Section 4.3, we find
+By the definition of $\zeta_{q}$ in (4.7), we have
 […]
-which shall be estimated.
-Let us begin with the bound of the first term on the right hand side of (4.53).
-Similar as the computation of oscillation error (4.29) in Section 4.5, by (4.13) we have for any $t\in[k,k+1]$, $k\in\mathbb{Z}$
-$w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}$
+Let us begin with the estimates for the first term on the right hand side of (5.38).
+Using (4.10) and the same computation of oscillation error in (4.17), we have for any $t\in[k,k+1]$, $k\in\mathbb{Z}$
+$w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}$ (5.39)
 $=\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)^{2}B_{\xi}\otimes B_{-\xi}+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}$
@@
 $=c^{-1}_{*}\rho\textrm{Id}-\mathring{R}_{\ell}+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}.$
-Taking the trace on both sides of the equation and using the fact that $\mathring{R}_{\ell}$ is traceless, we deduce for $t\in[k,k+1]$, $k\in\mathbb{Z}$
+Taking the trace on both sides of (5.39) and using the fact that $\mathring{R}_{\ell}$ is traceless, we deduce for $t\in[k,k+1]$, $k\in\mathbb{Z}$
 $|w_{q+1}^{(p)}|^{2}-3\zeta_{q}=3c_{*}^{-1}\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}+3(\zeta_{\ell}-\zeta_{q})+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}).$
-Hence, we have
+As a result, we have
 $\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(p)}|^{2}-3\zeta_{q}{\mathord{{\rm d}}}x\right|\leqslant$ $3c_{*}^{-1}(2\pi)^{3}\ell+3c_{*}^{-1}\mathbf{E}\|\mathring{R}_{\ell}\|_{L^{1}}+3\cdot(2\pi)^{3}|\zeta_{\ell}-\zeta_{q}|$ (5.40)
 $+$ $\mathbf{E}\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}){\mathord{{\rm d}}}x\right|.$
-In the following, we estimate each term separately, by the choice of parameters, we find
+In the following, we estimate each term separately. By the choice of parameters, we find
 $3c_{*}^{-1}(2\pi)^{3}\ell=3c_{*}^{-1}(2\pi)^{3}\lambda_{q}^{-2}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
 which requires $b^{2}\beta<1$ and $a$ sufficiently large to absorb the constant.
-Using (3.8f) on $\mathring{R}_{q}$, $\operatorname{supp}\varphi_{\ell}\subset[0,\ell]$ and $c_{R}<c_{*}$ we deduce for $t\in{\mathbb{R}}$
+We use (3.8f) on $\mathring{R}_{q}$, $\operatorname{supp}\varphi_{\ell}\subset[0,\ell]$ and $c_{R}<c_{*}$ to obtain for $t\in{\mathbb{R}}$
 $3c_{*}^{-1}\mathbf{E}\|\mathring{R}_{\ell}(t)\|_{L^{1}}\leqslant\frac{1}{16}\delta_{q+2}e(t).$
-For the third term in (4.54), we use (3.3), (3.7), (3.8b) and (3.8c) to obtain
+For the third term in (5.40), using (3.3), (3.7), (3.8b) and (3.8c) we obtain
 $3\cdot(2\pi)^{3}|\zeta_{\ell}-\zeta_{q}|$ $\lesssim\ell\|e^{\prime}\|_{C^{0}_{[t-1,t]}}+\ell\mathbf{E}\|v_{q}\|_{C^{1}_{[t-1,t],x}}(\|v_{q}\|_{C_{[t-1,t],x}^{0}}+{\|z_{q}\|_{C_{[t-1,t],x}^{0}}})$
@@
 $\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}\lambda_{q}^{\beta}+\ell^{1/2-\delta}\lambda_{q}^{\beta}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
-where we used $14b^{2}\beta+2\beta<$ 1 to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-\frac{5}{6}+\beta}<\lambda_{q}^{-2b^{2}\beta}$, and we $a$ was chosen sufficiently large to absorb the constant.
-For the last term in (4.54), recalling the definition of $\eta_{k,j}$ and $a_{(\xi)}$ in (4.12), we know that for any fixed $t\in[k,k+1]$, the sum with respect to $j,j^{\prime}$ is actually the finite sum of at most two terms.
+where we used $14b^{2}\beta+2\beta<$ 1 to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-\frac{5}{6}+\beta}<\lambda_{q}^{-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
+Recalling the definition of $\eta_{k,j}$ and $a_{(\xi)}$ in (4.9), we can simplify the last term in (5.40).
+For any fixed $t$ within the interval $[k,k+1]$, the sum concerning $j$ and $j^{\prime}$ is the finite sum of at most two terms.
 Hence, applying (5.22) with $n=m$ and using the estimates (5.1), (5.2) and (C.7) we obtain
@@
 This completes the bound for (5.40).
-Going back to (4.53), it remains to control
+Going back to (5.38), we have to control
 $\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(c)}|^{2}+2w_{q+1}^{(p)}\cdot w_{q+1}^{(c)}{\mathord{{\rm d}}}x\right|+\mathbf{E}\left|\int_{\mathbb{T}^{3}}|v_{\ell}-v_{q}+z_{q+1}-z_{q}|^{2}{\mathord{{\rm d}}}x\right|$ (5.41)
 $\qquad+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}-v_{q}+z_{q+1}-z_{q})(v_{q}+z_{q}){\mathord{{\rm d}}}x\right|+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}+z_{q+1})\cdot w_{q+1}{\mathord{{\rm d}}}x\right|.$
-Using the estimates (4.20) and (4.23), we have
+We use the estimates (5.3) and (5.6) to obtain
 $\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(c)}|^{2}+2w_{q+1}^{(p)}\cdot w_{q+1}^{(c)}{\mathord{{\rm d}}}x\right|\lesssim\mathbf{E}\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}^{2}+2\mathbf{E}\|w_{q+1}^{(p)}\|_{C^{0}_{t,x}}\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}$
@@
 where we used $b\geqslant 7$, $14b^{3}\gamma+4b^{2}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{4+3b^{3}\gamma}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
-Next, we use (3.3), (3.7), (3.8b), (4.6) and (4.33) to obtain
+Using (3.3), (3.7), (3.8b), (5.9) and (5.15) we obtain
 $\mathbf{E}$ $\int_{\mathbb{T}^{3}}|v_{\ell}-v_{q}+z_{q+1}-z_{q}|^{2}{\mathord{{\rm d}}}x+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}-v_{q}+z_{q+1}-z_{q})(v_{q}+z_{q}){\mathord{{\rm d}}}x\right|$
@@
 $\leqslant\lambda_{q}^{-\frac{1}{2}}\lambda_{q+1}^{\beta}+\lambda_{q+1}^{-\frac{\gamma\sigma}{8}+\beta}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\beta+\gamma}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
-where we used $2b\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta+2b^{2}\beta}<\lambda_{q}^{\frac{\gamma\sigma}{8}b}$, and used $14b^{2}\beta+2b\beta<$ 1, $3b\beta<\gamma$ to have $\lambda_{q}^{b\beta-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-b^{2}\gamma+b\beta+b\gamma}<\lambda_{q}^{-2b^{2}\beta}$, and $a$ was chosen large enough to absorb the constant.
-For the last term of (4.55), we recall (cf.(4.19)) that $w_{q+1}$ may be written as the curl of a vector field as
+where we used $2b\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta+2b^{2}\beta}<\lambda_{q}^{\frac{\gamma\sigma}{8}b}$.
+We also used $14b^{2}\beta+2b\beta<$ 1, $3b\beta<\gamma$ to have $\lambda_{q}^{b\beta-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{-b^{2}\gamma+b\beta+b\gamma}<\lambda_{q}^{-2b^{2}\beta}$ and chose $a$ large enough to absorb the constant.
+For the last term in (5.41), we recall (cf.(4.15)) that $w_{q+1}$ may be written as the curl of a vector field as
 $w_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j,\xi}\mathrm{curl}(a_{\xi}B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}}).$
@@
 $\lesssim\frac{\lambda_{q+3}^{\gamma}}{\lambda_{q+1}}(\lambda_{q}^{\frac{3}{2}}+\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},1})\leqslant\frac{1}{48}\delta_{q+2}e(t),$
-where we used $b\geqslant 7$ and $12b^{2}\beta+b^{3}\gamma<$ 1 to have $\lambda_{q}^{b^{3}\gamma+\frac{3}{2}}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{b^{3}\gamma+b\gamma}<\lambda_{q}^{b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
+where we used $b\geqslant 7$, $12b^{2}\beta+b^{3}\gamma<$ 1 to have $\lambda_{q}^{b^{3}\gamma+\frac{3}{2}}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{b^{3}\gamma+b\gamma}<\lambda_{q}^{b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant.
 Summing over the above estimates, we obtain
 $\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}\leqslant\frac{1}{4}\delta_{q+2}e(t).$
-Hence, the proof of Proposition 3.2 is completed. ∎
+Hence, the proof of Proposition 3.3 is complete. ∎
 6. Stationary solutions to the stochastic Euler equations
-We recall that the trajectory space defined in Section 1.1 is $\mathcal{T}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$ and we take $\kappa=\frac{\vartheta}{2}$ in this section.
+We consider that the trajectory space defined in Section 1.1 is $\mathcal{T}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$ and we take $\kappa=\frac{\vartheta}{2}$ in this section.
 The corresponding shifts $S_{t}$, $t\in{\mathbb{R}}$, on trajectories are given by
 $S_{t}(u,B)(\cdot)=(u(\cdot+t),B(\cdot+t)-B(t)),\quad t\in{\mathbb{R}},\quad(u,B)\in{\mathcal{T}}.$
-The notion of stationary solution was introduced in Definition 1.4.
-Our main result of this section is existence of stationary solutions as limits of ergodic averages of solutions constructed in Section 3.
-This in particular implies their non-uniqueness.
+Recall the notion of stationary solution introduced in Definition 1.5.
+Our main result of this section is the existence of stationary solutions which are constructed as limits of ergodic averages of solutions from Theorem 3.4.
+This also implies their non-uniqueness.
 Theorem 6.1.
@@
 First, we claim that there exist $\beta^{\prime},\beta^{\prime\prime}>0$ such that for any $N\in\mathbb{N}$
-$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
-Indeed, reviewing the proof of Theorem 3.3, and by interpolation we deduce for some $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$
+$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$ (6.2)
+Indeed, reviewing the proof of Theorem 3.4 and using interpolation, we deduce for some $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$
 $\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}$ $\lesssim\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\vartheta}_{t,x},2r}$
@@
 $\lesssim$ $\sum_{q\geqslant 0}\lambda_{q+1}^{-\frac{\gamma}{8}{\cdot}\frac{\sigma}{16}}\lambda_{q+2}^{\frac{\gamma}{32}\beta^{\prime\prime}}+\lambda_{q+2}^{-\frac{\gamma}{16}+\frac{\gamma}{32}\beta^{\prime\prime}}\lambda_{q+1}^{\frac{\gamma}{8}}\lesssim\sum_{q\geqslant 0}\lambda_{q+1}^{-\frac{\gamma}{8}{\cdot}\frac{\sigma}{16}+\frac{\gamma}{32}b\beta^{\prime\prime}}+\lambda_{q+1}^{\frac{\gamma}{8}+\frac{\gamma}{32}b\beta^{\prime\prime}-\frac{\gamma}{16}b}<\infty,$
-where we used $(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})>\frac{1}{16}$, $b\geqslant 7$ and $\beta^{\prime\prime}<{\mathord{{\rm min}}}{\{\frac{\sigma}{4b},1\}}$ in the last two inequalities.
+where we used $b\geqslant 7$, $(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})>\frac{1}{16}$ and $\beta^{\prime\prime}<{\mathord{{\rm min}}}{\{\frac{\sigma}{4b},1\}}$ in the last two inequalities.
 Hence we obtain $\lim_{q\rightarrow\infty}z_{q}=z$ in $L^{p}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$ for any $p\geqslant 1$.
@@
 $|\!|\!|u|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}<\infty.$
-Hence we obtain
-$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}$ $\leqslant\sup_{s\in{\mathbb{R}}}\sum_{i=-N}^{N-1}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([i,i+1],C^{\beta^{\prime\prime}}_{x})}$
-$\leqslant 2N\sup_{t\in\mathbb{R}}\mathbf{E}\|u\|^{2r}_{C^{\beta^{\prime}}([t,t+1],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
-Now, we define the ergodic averages of the solution $(u,B)$ as the probability measures on the space of trajectories ${\mathcal{T}}$
+Furthermore, by $\|u\|_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\leqslant\sum_{i=-N}^{N-1}\|u\|_{C^{\beta^{\prime}}_{[i,i+1]}C^{\beta^{\prime\prime}}_{x}}$ for any $N\in\mathbb{N}$ and Minkowski’s inequality, we obtain
+$\sup_{s\in{\mathbb{R}}}\left[\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\right]^{\frac{1}{2r}}$ $\leqslant\sup_{s\in{\mathbb{R}}}\sum_{i=-N}^{N-1}\left[\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}_{[i,i+1]}C^{\beta^{\prime\prime}}_{x}}\right]^{\frac{1}{2r}}$
+$\leqslant 2N\sup_{t\in\mathbb{R}}\left[\mathbf{E}\|u\|^{2r}_{C^{\beta^{\prime}}_{[t,t+1]}C^{\beta^{\prime\prime}}_{x}}\right]^{\frac{1}{2r}}=2N|\!|\!|u|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}\lesssim N,$
+where the implicit constants are independent of $N$. Hence, assertion (6.2) holds.
+Now, we define the ergodic averages of the solution $(u,B)$ as the probability measures on the trajectory space ${\mathcal{T}}$
 $\nu_{T}=\frac{1}{T}\int_{0}^{T}\mathcal{L}[S_{t}(u,B)]{\mathord{{\rm d}}}t,\qquad T\geqslant 0.$
-For $R_{N}>0$, $N\in\mathbb{N}$, by Arzelà–Ascoli Theorem and the above assertion, we note that the set
+For $R_{N}>0$, $N\in\mathbb{N}$, by Arzelà–Ascoli Theorem and the above claim, we have that the set
 $K_{M}:=\cap_{N=M}^{\infty}\bigg{\{}g;\,$ $\|g\|_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\leqslant R_{N}\bigg{\}}$
 is relatively compact in $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$ since $\frac{\vartheta}{2}<\beta^{\prime\prime}$.
-As a consequence, we deduce that the time shifts $S_{t}u$, $t\in{\mathbb{R}}$, are tight on $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$.
-Since $S_{t}B$ is a Wiener process for every $t\in{\mathbb{R}}$, the law of $S_{t}B$ does not change with $t\in{\mathbb{R}}$ and is tight.
-Accordingly, for any $\varepsilon>0$ there is a compact set $\bar{K}_{\varepsilon}$ in ${\mathcal{T}}$ such that
-$\sup_{t\in{\mathbb{R}}}{\mathbf{P}}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon})<\varepsilon.$
+We then deduce that the time shifts $S_{t}u$, $t\in{\mathbb{R}}$, are tight on $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$.
+Since $S_{t}B$ is a Wiener process for every $t\in{\mathbb{R}}$, the law of $S_{t}B$ is tight.
+Accordingly, for any $\varepsilon>0$ there is a compact set $F_{\varepsilon}$ in ${\mathcal{T}}$ such that
+$\sup_{t\in{\mathbb{R}}}{\mathbf{P}}(S_{t}(u,B)\in F^{c}_{\varepsilon})<\varepsilon.$
 This implies
-$\nu_{T}(\bar{K}_{\varepsilon}^{c})=$ $\frac{1}{T}\int_{0}^{T}\mathbf{P}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon}){\mathord{{\rm d}}}t<\varepsilon$
-and therefore there is a weakly converging subsequence of the probability measures $\nu_{T}$, $T\geqslant 0$.
-That is, there is a subsequence $T_{n}\to\infty$ and $\nu\in{\mathcal{P}}({\mathcal{T}})$ such that $\nu_{T_{n}}\to\nu$ weakly in ${\mathcal{P}}({\mathcal{T}})$.
-Define the set
-$A$ =$=\Big{\{}(u,B)\in{\mathcal{T}};\ \langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r$
-$\qquad\qquad=\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle,\quad\forall\psi\in C^{\infty}(\mathbb{T}^{3};{\mathbb{R}}^{3}),\ \mathord{{\rm div}}\psi=0,\ t\geqslant s\Big{\}}.$
-Since $(u,B)$ in the statement of the theorem satisfies the equation, we have for all $t\in{\mathbb{R}}$
-${\mathcal{L}}[S_{t}(u,B)](A)=1.$
-Hence, also $\nu_{T_{n}}(A)=1$ for all $n\in\mathbb{N}$.
-By Jakubowski–Skorokhod representation theorem, there is a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}})$ with a sequence of random variables $(\tilde{u}^{n},\tilde{B}^{n})$, $n\in\mathbb{N}$, such that ${\mathcal{L}}[\tilde{u}^{n},\tilde{B}^{n}]={\nu_{T_{n}}}$ and $(\tilde{u}^{n},\tilde{B}^{n})$ satisfy equation (1.1) on ${\mathbb{R}}$.
-By (3.12) we know
-$\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2}=\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|{S_{s}u(t)}\|_{L^{2}}^{2}{\mathord{{\rm d}}}s=K.$ (5.2)
-By (3.11)
-$\sup_{n}\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2r}=\sup_{n}\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|S_{s}u(t)\|_{L^{2}}^{2r}{\mathord{{\rm d}}}s<\infty.$ (5.3)
-Moreover, there is a random variable $(\tilde{u},\tilde{B})$ having the law ${\mathcal{L}}[\tilde{u},\tilde{B}]=\nu$ so that
-$(\tilde{u}^{n},\tilde{B}^{n})\to(\tilde{u},\tilde{B})\qquad\tilde{\mathbf{P}}\text{-a.s. in }\mathcal{T}.$
-Thus, we can pass to the limit in the equation to deduce that $\nu$ is a law of a solution on ${\mathbb{R}}$.
-The shift invariance follows from the same argument as in [BFH20e, Lemma 5.2].
-Namely, it holds for $G\in C_{b}({\mathcal{T}})$ and $r\in{\mathbb{R}}$
-$\int_{{\mathcal{T}}}G\circ S_{r}(u,B){\mathord{{\rm d}}}\nu(u,B)=\int_{{\mathcal{T}}}G(u,B){\mathord{{\rm d}}}\nu(u,B).$
-Finally, (5.1) follows from (5.2) and (5.3). ∎
-Using the above result and choosing different $K$, the second claim in Theorem 1.5 follows.
+$\nu_{T}(F_{\varepsilon}^{c})=$ $\frac{1}{T}\int_{0}^{T}\mathbf{P}(S_{t}(u,B)\in F^{c}_{\varepsilon}){\mathord{{\rm d}}}t<\varepsilon.$
+Therefore, there is a weakly converging subsequence of the probability measures $\nu_{T}$, $T\geqslant 0$.
+The rest of the proof follows by Jakubowski–Skorokhod representation theorem and the same argument as in [HZZ22b, Theorem 4.1].
+∎
+Using the above result and choosing different $K$, we conclude Theorem 1.6.
 Appendix 1 A Beltrami waves
-In this part we recall the Beltrami waves from [BV19, Section 5.4] whcih is adpated to the convex integration scheme in Proposition 3.2.
-We point out that the construction is entirely deterministic, that is, none of the functions below depends on $\omega$.
+In this part, we recall the Beltrami waves from [BV19, Section 5.4] which is adapted to the convex integration scheme in Proposition 3.3.
+We point out that the construction is entirely deterministic, meaning that none of the functions below depends on $\omega$.
 Let us begin with the definition of Beltrami waves.
@@
 $W_{(\xi)}:=W_{\xi,\lambda}(x):=B_{\xi}e^{i\lambda\xi\cdot x}$ (A.1)
-is $\mathbb{T}^{3}$ periodic, divergence free, and is an eigenfunction of the $\mathrm{curl}$ operator with eigenvalue $\lambda$.
+is $\mathbb{T}^{3}$ periodic, divergence-free, and is an eigenfunction of the $\mathrm{curl}$ operator with eigenvalue $\lambda$.
 That is, $W_{(\xi)}$ is a complex Beltrami plane wave.
@@
 […]
-The key point of the construction is that the abundance of Beltrami waves allows to find several such flows $v$ with the property that
+The key point of the construction is that the abundance of Beltrami waves allows us to find several such flows $v$ with the property that
 $\frac{1}{(2\pi)^{3}}\int_{\mathbb{T}^{3}}v\otimes v(t,x){\mathord{{\rm d}}}x$
@@
 Indeed we will need to select these flows so as to depend smoothly on the matrix $R$, at least when $R$ belongs to a neighborhood of the identity matrix.
-In view of (A.3), such selection is made possible by the following lemma.
+In view of (A.3), such selection is made possible by the following lemma from [BV19, Proposition 5.6].
 Lemma A.2.
@@
 Then, there exist pairwise disjoint subsets
-$\Lambda_{\alpha}\subset\mathbb{S}^{2}\cap\mathbb{Q}^{3}\qquad\alpha\in\{0,1\},$
+$\Lambda_{\alpha}\subset\mathbb{S}^{2}\cap\mathbb{Q}^{3},\qquad\alpha\in\{0,1\},$
 and smooth positive functions
-$\gamma_{\xi}^{(\alpha)}\in C^{\infty}(B_{c_{*}}(\mathrm{Id}))\qquad\alpha\in\{0,1\},\xi\in\Lambda_{\alpha},$
+$\gamma_{\xi}^{(\alpha)}\in C^{\infty}(B_{c_{*}}(\mathrm{Id})),\qquad\alpha\in\{0,1\},\ \xi\in\Lambda_{\alpha},$
 such that the following hold.
@@
 holds for $n$ large enough, $\alpha\in\{0,1\}$ and $\xi\in\Lambda_{\alpha}$. This parameter is universal.
-Appendix 2 B Estimates of $a_{(\xi)}$ in the proof of Proposition 3.2
-For completeness, we include here the detailed proof of the estimates (4.14) and (4.15).
-We first aim at estimating the $C^{N}_{t,x}$-norm of $\rho=\sqrt{|\mathring{R}_{\ell}|^{2}+\ell^{2}}+c_{*}\zeta_{\ell}$ for $N\in\mathbb{N}$.
-To this end, we first apply the chain rule [BDLIS15, Proposition C.1] to the function $\Psi(z)=\sqrt{\ell^{2}+z^{2}}$, since $|D^{m}\Psi(z)|\lesssim\ell^{-m+1}$, for $m\geqslant 1$, we obtain for any $N\in\mathbb{N}$
-$\|\rho\|_{C^{N}_{t,x}}$ $\lesssim\left\|\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}\right\|_{C^{0}_{t,x}}+\|D\Psi\|_{C^{0}}\|\mathring{R}_{\ell}\|_{C^{N}_{t,x}}+\|D\Psi\|_{C^{N-1}}\|\mathring{R}_{\ell}\|^{N}_{C^{1}_{t,x}}+\|\zeta_{\ell}\|_{C^{N}_{t}}$ (B.1)
-$\lesssim\ell+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-N}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2N+1}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{N}+\ell^{-N}\delta_{q+1}\bar{e},$
-where we have used definition of $\zeta_{\ell}$ (4.10) in the last inequality. Let us now estimate the $C^{N}_{t,x}$-norm of $a_{(\xi)}$. By Leibniz rule, we get
-$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\sum_{m=0}^{N}\|\rho^{\frac{1}{2}}\|_{C^{m}_{t,x}}\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}},$ (B.2)
-and estimate each term separately. First, by (4.10)
-$\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\lesssim\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}},$
-and by Lemma A.2
-$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{0}_{t,x}}\lesssim 1.$
-Second, applying [BDLIS15, Proposition C.1] to the function $\Psi(z)=z^{1/2},$ $|D^{m}\Psi(z)|\lesssim|z|^{1/2-m},$ for $m=1,\dots,N$, and using (B.1) and $\rho\geqslant\ell$ and $\ell^{-1}\geqslant\bar{e}$ we obtain for $m\geqslant 1$
-$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim$ $\|\rho^{1/2}\|_{C^{0}_{t,x}}+\ell^{-1/2}\|\rho\|_{C^{m}_{t,x}}+\ell^{1/2-m}\|\rho\|_{C^{1}_{t,x}}^{m}$ (B.3)
-$\lesssim$ $\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}}+\ell^{\frac{1}{2}-2m}\|\mathring{R}_{q}\|^{m}_{C^{0}_{[t-1,t+1],x}}$
-$+\ell^{-\frac{1}{2}-m}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}-2m}\delta_{q+1}^{m}+\ell^{-\frac{1}{2}-m}\delta_{q+1}.$
-For $m\geqslant 1$, using $\delta_{q+1}\leqslant 3rL^{2}$ the above is bounded by
-$\ell^{\frac{1}{2}-2m}(3rL^{2}+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{m}.$
-We proceed with a bound for $\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ for $m=0,\dots,N-1$.
-Keeping [BDLIS15, Proposition C.1] in mind, we need to estimate
-$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}+\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}+\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}.$ (B.4)
-We use $\rho\geqslant\ell$ to have
-$\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}},$ (B.5)
-and in view of $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$
-$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)},$
-and by (B.1) and $\delta_{q+1}\leqslant 3rL^{2}$
-$\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)},$
-hence
-$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$ (B.6)
-To estimate the first term in (B.4), we write
-$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\sum_{k=0}^{N-m}\|\mathring{R}_{\ell}\|_{C^{k}_{t,x}}\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}},$ (B.7)
-where for $N-m-k=0$ we have
-$\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}\lesssim\ell^{-1},$
-and for $k=0,\dots,N-m-1$ using (B.1)
-$\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}}$ $\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}+\ell^{-2}\|\rho\|_{C^{N-m-k}_{t,x}}+\ell^{-(N-m-k){-1}}\|\rho\|_{C^{1}_{t,x}}^{N-m-k}$
-$\lesssim\ell^{-2(N-m-k)-1}+\ell^{-(N-m-k)-2}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m-k)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{N-m-k}.$
-Altogether, we therefore bound (B.7) as
-$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\ell^{-2(N-m)-1}$ $\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$ (B.8)
-$+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}.$
-Finally, plugging (B.5), (B.6) and (B.8) into (B.4) leads to
-$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ $\lesssim\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$
-$+\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$
-For $N-m\geqslant 1$ the above is bounded by
-$\ell^{-2(N-m)-1}(\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+1).$
-Combining this with the bounds (B.3) above and plugging into (B.2) yields for $N\geqslant 2$
+Appendix 2 B Proof of Proposition 5.1
+We concisely present the proof of Proposition 5.1, following closely the calculation of [HZZ22b, Appendix B].
+We only write the necessary changes compared to [HZZ22b, Appendix B].
+By the chain rule, we have
+$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\sum_{m=0}^{N}\|\rho^{\frac{1}{2}}\|_{C^{m}_{t,x}}\big{\|}\gamma_{\xi}\big{(}\textrm{Id}-\rho^{-1}\mathring{R}_{\ell}\big{)}\big{\|}_{C^{N-m}_{t,x}}.$ (B.1)
+We estimate the two terms on the right hand side in (B.1) separately. First, using Lemma A.2 and (4.6) we have
+$\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\lesssim\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}},\qquad\|\gamma_{\xi}(\textrm{Id}-\rho^{-1}\mathring{R}_{\ell})\|_{C^{0}_{t,x}}\lesssim 1.$
+Second, we use [HZZ22b, (B.4)] to derive for $m\geqslant 1$
+$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim\|\rho^{1/2}\|_{C^{0}_{t,x}}+\ell^{-1/2}\|\rho\|_{C^{m}_{t,x}}+\ell^{1/2-m}\|\rho\|_{C^{1}_{t,x}}^{m}.$ (B.2)
+By the definition of $\rho$ in (4.6), it follows from [HZZ22b, (B.1)] that for any $m\geqslant 1$
+$\|\rho\|_{C^{m}_{t,x}}\lesssim\ell+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-m}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2m+1}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{m}+\ell^{-m}\delta_{q+1}\bar{e}.$ (B.3)
+Hence, combining (B.2) with (B.3), we have for any $m\geqslant 1$
+$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim\ell^{\frac{1}{2}-2m}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{m}.$ (B.4)
+We now consider the second term on the right hand side of (B.1). By [HZZ22b, (B.5)], we have for $m=0,\dots,N-1$
+$\left\|\gamma_{\xi}\left(\textrm{Id}-\rho^{-1}\mathring{R}_{\ell}\right)\right\|_{C^{N-m}_{t,x}}\lesssim\left\|\rho^{-1}\nabla_{t,x}\mathring{R}_{\ell}\right\|_{C^{0}_{t,x}}^{N-m}+\left\|\rho^{-2}\mathring{R}_{\ell}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}+\left\|\rho^{-1}\mathring{R}_{\ell}\right\|_{C^{N-m}_{t,x}}.$ (B.5)
+Using $\rho\geqslant\ell$, $|\rho^{-1}\mathring{R}_{\ell}|\leqslant 1$, (B.3) and [HZZ22b, (B.6)], we deduce that for any $N-m\geqslant 1$
+$\left\|\gamma_{\xi}\left(\textrm{Id}-\rho^{-1}\mathring{R}_{\ell}\right)\right\|_{C^{N-m}_{t,x}}\lesssim\ell^{-2(N-m)-1}(\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+1)^{N-m+1}.$ (B.6)
+Then combining (B.6) with the bound (B.4) and plugging into (B.1) yields for $N\geqslant 2$
 $\|a_{(\xi)}\|_{C^{N}_{t,x}}$ $\lesssim\ell^{-2N-\frac{1}{2}}(1+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})^{N+1}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (B.7)
-where we used inductive assumption (3.8d) for $\mathring{R}_{q}$ in the last inequality. Moreover, for $N=1$, by (B.3) and chain rule, we obtain
-$\|a_{(\xi)}\|_{C^{1}_{t,x}}$ $\lesssim\|\rho^{\frac{1}{2}}\|_{C^{1}_{t,x}}+\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\left\|\gamma_{\xi}\left(\mathrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{1}_{t,x}}$ (B.10)
-$\lesssim\ell^{-\frac{3}{2}}(\delta_{q+1}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})+(\delta_{q+1}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{\frac{1}{2}})\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{1}_{t,x}}.$
-Since $\rho\geqslant\ell$ and $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$, by (B.3) and using chain rule again, we obtain
-$\left|\nabla_{t,x}\left(\frac{\mathring{R}_{\ell}}{\rho}\right)\right|\lesssim\frac{|\nabla_{t,x}\mathring{R}_{\ell}|}{\rho}+\frac{|\mathring{R}_{\ell}\cdot\nabla_{t,x}\rho|}{\rho^{2}}\lesssim\ell^{-2}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2}.$ (B.11)
-Pluging (B.11) into (B.10), we obtain
-$\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{\frac{3}{2}}\leqslant\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (B.12)
-And by the definition of $a_{(\xi)}$, it is easy to get
+where we used inductive assumption (3.8d) for $\mathring{R}_{q}$ in the last inequality.
+Moreover, by a similar calculation as above, we also have
+$\|a_{(\xi)}\|_{C^{1}_{t,x}}$ $\lesssim\ell^{-2}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{\frac{3}{2}}\leqslant\ell^{-2}\lambda_{q+3}^{2\gamma},$ (B.8)
 $\|a_{(\xi)}\|_{C^{0}_{t,x}}$ $\lesssim\|\mathring{R}_{q}\|^{\frac{1}{2}}_{C_{[t-1,t+1],x}^{0}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\gamma}.$ (B.9)
 Hence, (5.1) and (5.2) follow from the estimates (B.7), (B.8) and (B.9).
-Appendix 3 C Estimates for transport equantions
+Appendix 3 C Estimates for transport equations
 We provide a detailed estimate of the solutions to the transport equations in Section 4.3.1. First, we consider the following transport equation on $[t_{0},T],t_{0}\geqslant 0$:
@@
 $f(t_{0},x)$ $=f_{0}.$
-We recall the following results from [BDLIS15, Proposition D.1, (133), (134)]: {equ} ∥f(t)∥_C_x^1⩽∥f_0∥_C_x^1e^(t-t_0)∥v∥_C^0_[t_0,t]C^1_x+∫_t_0^te^(t-τ)∥v∥_C_[t_0,t]^0C_x^1∥g(τ)∥_C_x^1 dτ, and more generally, for any $N\geqslant 2$, there exists a constant $C=C_{N}$ such that
+We recall the following results from [BDLIS15, Proposition D.1, (133), (134)]: {equ} ∥f(t)∥_C_x^1⩽∥f_0∥_C_x^1e^(t-t_0)∥v∥_C^0_[t_0,t]C^1_x+∫_t_0^te^(t-τ)∥v∥_C_[t_0,t]^0C_x^1∥g(τ)∥_C_x^1 dτ.
+More generally, for any $N\geqslant 2$, there exists a constant $C=C_{N}$ such that
 $\|f(t)\|_{C_{x}^{N}}\leqslant$ $\left(\|f_{0}\|_{C_{x}^{N}}+C(t-t_{0})\|{v}\|_{C^{0}_{[t_{0},t]}C^{N}_{x}}\|f_{0}\|_{C_{x}^{1}}\right)e^{C(t-t_{0})\|{v}\|_{C^{0}_{[t_{0},t]}C^{1}_{x}}}$ (C.2)
@@
 $\Phi(t_{0},x)$ $=x.$
-We aim at estimating the $C_{x}^{0}$-norm of $\nabla\Phi(t)-\textrm{Id}$. First, let $\Psi(s,x)=\Phi(s,x)-x$, then $\Psi$ statifies the following equation:
+We aim to estimate the $C_{x}^{0}$-norm of $\nabla\Phi(t)-\textrm{Id}$. First, let $\Psi(s,x)=\Phi(s,x)-x$, then $\Psi$ satisfies the following equation:
 $(\partial_{t}+v\cdot\nabla)\Psi=-v,$
@@
 $\lesssim\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}\lesssim\lambda_{q}^{-\frac{1}{2}}\ll 1,$
-hence (4.8a) holds ture. Then, (4.8b) follows from (C.5), we have
+hence (4.4a) holds. Then, (4.4b) follows from (C.5), we have
 $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\leqslant\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\mathrm{Id}\|_{C_{x}^{0}}+1\leqslant 2,$ (C.6)
 $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\geqslant 1-\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\mathrm{Id}\|_{C_{x}^{0}}\geqslant\frac{1}{2}.$
-Similarly, we use (C.2) and (C.4) again, we obtain for any $n\geqslant 1$
+Similarly, we use (C.2) and (C.4) again to obtain for any $n\geqslant 1$
 $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{n}}$ $\lesssim\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{n+1}}}e^{2\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}}$ (C.7)
@@
 $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\partial_{t}\Phi_{k,j}\|_{C_{x}^{0}}\leqslant\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|(v_{\ell}+z_{\ell})\cdot\nabla\Phi_{k,j}\|_{C_{x}^{0}}\leqslant\lambda_{q+3}^{\frac{\gamma}{2}}.$ (C.8)
-Differentiating both sides of (4.7) and using (3.8a), (3.8c), (4.31), (C.6) and (C.7), we can also obtain
+Differentiating both sides of (4.3) and using (3.8a), (3.8c), (5.13), (C.6) and (C.7), we also obtain
 $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\partial_{t}\nabla\Phi_{k,j}\|_{C_{x}^{0}}\lesssim$ $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla(v_{\ell}+z_{\ell})\cdot\nabla\Phi_{k,j}\|_{C_{x}^{0}}$ (C.9)